### v1

# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

Abstract.
The healthcare landscape is evolving, with patients seeking more reliable information about their health conditions, treatment options, and potential risks. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust doctors and hospital staff, highlighting the need for expert-endorsed health information. However, the pressure on experts has led to reduced communication time, impacting information sharing. To address this gap, we propose CataractBot, an experts-in-the-loop chatbot powered by large language models (LLMs). Developed in collaboration with a tertiary eye hospital in India, CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously. CataractBot features multimodal support and multilingual capabilities. In an in-the-wild deployment study with 49 participants, CataractBot proved valuable, providing anytime accessibility, saving time, and accommodating diverse literacy levels. Trust was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
1. Introduction
The evolving landscape of healthcare is witnessing a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007). It signifies a departure from the traditional passive patient role toward a more informed, empowered, and engaged approach to healthcare, resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their health conditions, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery or radiotherapy, this demand for information becomes more crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Conversely, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007). This poses a challenge in discerning trustworthy sources from misleading ones, leading to potential misconceptions. Additionally, patients frequently encounter difficulties finding information online at an appropriate level–ranging from oversimplified to overly technical–and harbor skepticism toward its reliability (Xiao et al., 2023). As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although locating them can be challenging (Civan-Hartzler et al., 2010). Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a). Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998). Study participants noted that being better informed increased their understanding of treatment plans, motivation to adhere to them, and satisfaction with hospital visits.
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Even providing all potentially relevant information can pose challenges as it can overwhelm patients with cognitive overload (Christalle et al., 2019). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, subsequently alleviating the burden on doctors and medical staff.
To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with the medical team at a tertiary eye hospital in Bangalore, India, we exemplified our approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized RAG (Retrieval-Augmented Generation) using a knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and updates to the knowledge base based on expert-provided edits to minimize future expert intervention. In particular, for high-stakes medical settings, access to reliable and trustworthy information is crucial, hence, each response that patients (or their attendants) receive undergoes scrutiny and potential editing by experts (doctors or patient coordinators). CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantaneously and subsequently providing verification at the experts’ convenience.
With that, the paper aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and their attendants) undergoing surgical treatment? (RQ2) How does the inclusion of an expert-in-the-loop influence the perceived trustworthiness of the information provided by the chatbot? (RQ3) How are such chatbots integrated into the workflow of doctors, patient coordinators, patients, and their attendants?
To answer them, we conducted an in-the-wild deployment study involving 49 information seekers (19 patients, 30 attendants) using CataractBot for at least two weeks—a week before and after the surgery. Additionally, four doctors and two patient coordinators served as experts in this study. Our findings revealed patients and attendants highly appreciated CataractBot for its anytime accessibility, alleviating hesitations associated with power differences in asking questions, saving time by minimizing hospital visits/calls, and accommodating individuals with low literacy and tech proficiency through multilingual and multimodal support. Trust in the system was established through rigorous verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous knowledge base updates. Our main contributions are: (1) The conception and development of a specialized chatbot111Our code is available at https://github.com/microsoft/BYOeB., designed in close collaboration with doctors and hospital staff, to assist patients undergoing Cataract surgery with their information needs. (2) The deployment of this chatbot CataractBot in a real-world setting, among 31 Cataract surgery patients and their attendants at a tertiary eye hospital in India. Four doctors and two patient coordinators served as experts, actively verifying and refining responses generated by the chatbot. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, our study not only presents pivotal insights specific to CataractBot but also offers broader implications of LLM-powered expert-in-the-loop chatbots.
2. Related Work
Our work is primarily informed by two areas of relevant research: solutions to address patients’ information needs and conversational agents used in healthcare settings.
2.1. Addressing Patients’ Information Needs
Prior work has focused on understanding the diverse informational requirements of patients, particularly in the context of healthcare decision-making (Tseng et al., 2023; Xiao et al., 2023; Beisecker and Beisecker, 1990). Patients and their attendants exhibit a pattern of seeking information throughout the various stages of a patient’s journey, such as to prepare for doctor consultations (Radionova et al., 2023; Leong et al., 2018), to verify diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009), to assess the need for clinical interventions (Yang et al., 2023; Tseng et al., 2023), and to aid in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018). Studies emphasize that satisfying these diverse information needs enhances patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998).
To fulfill these diverse informational needs, patients, while trusting physicians for their clinical expertise, often turn to the Internet for the ease of access to information (Xiao et al., 2023). However, the challenge lies in identifying relevant and appropriate information, with instances of individuals consuming inaccurate information (Clarke et al., 2016b). Information seekers also resort to consulting experienced patients within their social network who can share valuable knowledge and insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010).
Recent work in the field of Computer-Human Interaction has explored novel avenues to address the medical information gaps. For instance, Leong et al. (2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, while Wilcox et al. (2010) and Bickmore et al. (2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details based on their records. Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005). Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
In summary, patients require reliable, relevant, and timely information. However, challenges arise in accessing such information from busy or hesitant doctors and hospital staff. This dichotomy results in patients either receiving generic, unverified information from the internet and social circles or obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts. Our solution aims to strike a balance by providing a free, generic, high-quality answer instantaneously using Large Language Models (LLMs), which is later verified (and corrected if needed) by a medical professional with minimal increase in their workload.
2.2. Chatbots in Healthcare
Recently, researchers have been exploring chatbots as a means to provide easy and direct access to information, due to their ubiquity, minimal learning curve, and ability to offer a more personalized experience through natural language conversation (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020). Chatbots have been found to enhance the user experience by enabling users to engage in longer conversations and build complex queries message by message while retaining conversation context (Xiao et al., 2023). Stamenova et al. (2020) reveal that patients prefer messaging over video consultations, emphasizing that chatbots can help promote continuity in the treatment. Chatbots have been applied in various healthcare settings by the research community, including mental health support (Kumar et al., 2023; Jo et al., 2023), appointment scheduling, seeking information for chronic disease management (Montagna et al., 2023), assistance for new mothers (Yadav et al., 2019), and public health interventions to support socially isolated individuals (Jo et al., 2023). Commercial offerings, such as Babylon Health111Babylon Health: https://www.emed.com/uk, Ada222Ada: https://ada.com/, and Florence333Florence: https://florence.chat/, allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs, and the involvement of experts is limited to the data curation step (Xiao et al., 2023). Hence, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
Beyond standalone chatbots, prior work has explored chat as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both the patient and the doctor to be simultaneously active, resembling an in-person consultation, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp, WeChat (Wang et al., 2020), or email enable communication between the patient and doctor based on their convenience, providing flexibility. Previous research indicates that both doctors and patients prefer asynchronous texts over instant calls (audio or video) (Stamenova et al., 2020; Johansson et al., 2020a).
In our work, we combine synchronous and asynchronous modes of communication, leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode. This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.
3. CataractBot System Design
3.1. Design Requirements
Drawing from our discussions with doctors and patient coordinators, literature review, and understanding of the Indian healthcare ecosystem, we have identified several key requirements for the chatbot.
Accuracy: The chatbot’s capability to answer both medical and logistical questions with accuracy is a fundamental requirement. Providing precise, contextually relevant, and verified responses is paramount given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it might not be tailored to India-specific or hospital-specific nuances. For instance, while the phacoemulsification method is prevalent globally for cataract surgeries, India often favors manual small incision-based procedures due to their cost-effectiveness (Singh et al., 2017). Similarly, cultural considerations significantly impact advice on pre- and post-surgery choices, e.g., questions related to diet necessitates inclusion and exclusion of Indian food options for accurate guidance. To ensure accurate and contextual response, our chatbot generates responses from a knowledge base curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
Trustworthiness: In medical surgical treatment, access to reliable trustworthy information is crucial. Previous research (Xiao et al., 2023) highlights how patients often encounter misinformation from Internet-based (medical) sources, leading to skepticism towards them. Consequently, they tend to rely on the experiences of friends and family members who have undergone similar treatments, although such resources can sometimes be challenging to locate (Civan-Hartzler et al., 2010). Typically, patients place their utmost trust in the operating doctor and the hospital staff responsible for their care during the procedure (Larasati et al., 2023). Hence, each response the patient (or their attendant) receives is vetted and, if necessary, edited by doctors and patient coordinators to ensure trustworthiness.
Timeliness: Prompt responses are important as delays often lead patients to seek information elsewhere. Patient coordinators revealed that patients frequently reach out to them (over a phone call) for relaying medical queries to their consulting doctors. However, due to busy schedules of doctors and coordinators, immediate responses are not always feasible. That often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation. To counter this, we decided the chatbot should provide instantaneous responses to every patient query using a custom knowledge base curated by hospital staff. Subsequently, these responses were reviewed and refined by doctors/coordinators later.
Usability: The developed system must be usable by patients, their attendants, doctors and patient coordinators. Thus, we chose WhatsApp as our platform to develop CataractBot, as it is the most widely used messaging platform in India (Statista, 2022). Despite Telegram’s richer feature set or a custom chatbot offering more development flexibility, we opted for WhatsApp to minimize the learning curve and ensure seamless onboarding without requiring additional app installations. Furthermore, due to the frequent use of WhatsApp among Indians (Statista, 2022), we anticipated higher adoption and usage of a WhatsApp-based chatbot. To cater to the diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, we made the chatbot multilingual—supporting five languages (English, Hindi, Kannada, Tamil, and Telugu)—and multimodal (offering both text and speech inputs and outputs), similar to (Jain et al., 2018; Findlater et al., 2009).
3.2. Components
From our discussions with doctors and hospital staff, we have identified four key stakeholders in the cataract surgery ecosystem: the patient, their attendants, the doctor, and the patient coordinator. Once onboarded on CataractBot, a patient (and their attendant) scheduled for cataract surgery within the next 7 days can ask any question related to the surgical treatment and receive a response verified by the hospital staff. In addition, we have a few more roles: an escalation doctor, an escalation patient coordinator, and a KB (Knowledge Base) update expert. Below, we describe the key components of our CataractBot system.
3.2.1. Input Language and Modality
The Sankara Eye Hospital situated in Bangalore (in the state of Karnataka), known as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language via WhatsApp. Among these messages is an option allowing users to modify their preferred language using a drop-down menu at any time in the future. Note: To converse in a specific language, users must ensure their preferred language is selected before interacting. For example, if a user has chosen English as their preferred language, they cannot begin conversing with CataractBot in Kannada. However, they can navigate to the welcome messages, switch their preferred language to Kannada, and then engage with the bot in Kannada.
To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, both text and speech inputs are enabled (Jain et al., 2018; Findlater et al., 2009). Users can ask questions by holding the mic button on WhatsApp, speaking in their preferred language, and sending a voice message. For every voice message, CataractBot not only provides a text response but also sends an audio message (Figure 2D). Given that the doctor and patient coordinator may not be proficient in all five languages, their interactions with the bot for verifying and editing messages are exclusively conducted in English.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot addresses this by providing a set of three frequently asked questions in a drop-down menu along with the welcome message. Additionally, after every medical and logistical question from the patient/attendant, the bot’s response includes a “What to do next?” prompt, offering three related questions in a drop-down menu based on the preceding query (Figure 2F). This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time. Users are free to select one of the suggested questions or pose a different query to continue the conversation.
3.2.2. Response Generation, Verification and Icons
Upon receiving a message from a patient/attendant, CataractBot classifies it as a medical question, logistical question, small talk, or other queries.
For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response.
In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response.
For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
For miscellaneous questions unrelated to medical or logistical matters, e.g., about sports or weather, the bot responds with a template “I don’t know” message.
The patient/attendant receives these responses in real-time from CataractBot.
Specifically, for medical and logistical questions, the response includes a question mark icon
as a reaction indicating the unverified status (Figure 2E).
For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, patient demographics, and citations of the documents used to generate the response.
The doctor is then prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
Tapping ‘Yes’ replaces the question mark icon
in the user’s received answer with a green tick icon
, confirming verification (Figure 1F). Additionally, the bot notifies the user that the answer has been verified, tagging that particular response (Figure 1G).
Tapping ‘No’ replaces the question mark
with a red cross icon
, indicating an incorrect answer (Figure 1C), and asks the user to await a corrected response from the doctor (Figure 1D).
The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they need to offer a correction in free form.
The CataractBot system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
(Figure 1E).
If a question is misclassified, i.e., a logistical question is sent to the doctor, they have the option to ‘Send to Patient Coordinator’.
Similar to the doctor’s workflow, for logistical questions, the patient coordinators verify and provide corrections to the bot’s response.
In deciding how to deliver chatbot responses to patients, we deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction. Through our discussions with hospital staff, it became clear that providing real-time responses was crucial. Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
3.2.3. Escalation Mechanism
The CataractBot system employs an intricate escalation mechanism to ensure swift verification of every query posed by the patient/attendant.
In the doctor’s WhatsApp interface, unanswered questions are marked with a question mark icon
(Figure 2E), enabling them to easily identify pending queries.
If an answer is not verified by the operating doctor within three hours, it automatically escalates to the designated escalation doctor.
This duration was set based on input from the hospital staff.
The operating doctor receives a notification about the escalation.
Both the operating and escalation doctors can then verify/correct the response.
Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick
(Figure 1F) for both the doctors, indicating task completion.
However, if either doctor selects ‘No’, the green tick appears only after that doctor provides a correction.
In instances where neither the operating nor escalation doctor verifies a question within six hours, both receive a notification indicating the pending status.
Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors, tagged with appropriate messages.
This additional reminder was added based on the doctor’s feedback.
This workflow is mirrored for the patient coordinator and escalation coordinator.
3.2.4. KB Update Process
While verification remains vital in the medical context, we aim to reduce the effort required from already burdened medical professionals in verifying and editing CataractBot responses. To achieve this, we devised a mechanism to minimize repetitive edits performed by experts. This process involves using the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, we understand that certain responses, such as those specific to individual patients (e.g., “when to reach the hospital for my surgery?”), should not be added to the KB. Thus, we appointed a senior cataract surgery doctor as the ‘KB update expert’. Each night at 8 pm, this expert automatically receives an online spreadsheet via email containing columns such as ‘Question asked’, ‘Bot’s response’, ‘Expert’s correction’, ‘Bot’s Corrected Response’, and ‘Should Update KB?’. Their task involves reviewing each row and determining whether the information should be added to the CataractBot KB. They respond with a ‘Yes’ or ‘No’ in the ‘Should Update KB?’ column. If marked ‘Yes’, the KB update expert can modify the ‘Bot’s Corrected Response’ column as needed. At 3 am daily, we extract ‘Question asked’ and ‘Bot’s Corrected Response’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document. Additionally, we prompt the large language model (LLM) to prioritize this ‘expert-FAQ’ document within the KB. This ensures that the bot learns from these updates, gradually enhancing its knowledge over time. Consequently, when the bot encounters a similar question in the future, it can provide an accurate response without requiring further corrections, thus effectively reducing the workload of doctors and coordinators.
3.3. Implementation Details
The CataractBot system mainly relies on these five services (Figure 3): (a) LLM (Large Language Model): for response generation, (b) Vector Database: for storing and retrieving the custom knowledge base, (c) Language Technologies: for translation, and for converting speech-to-text and text-to-speech, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). Considering the medical context, instead of relying on GPT-4’s default knowledge, CataractBot utilized a curated knowledge base by the hospital staff. This custom KB comprises of various Cataract surgery and Sankara Eye Hospital-related documents, including Consent Form, SOP (Standard Operating Procedure), FAQs (Frequently Asked Questions), Pre- and Post-Operative Guidelines, Hospital Information, and Patient Discharge Do’s and Dont’s. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question asked by patient/attendant, GPT-4 classifies the query type (medical, logistical, small talk, or other). For a medical/logistical question, CataractBot employs a RAG (Retrieval Augmented Generation) (Lewis et al., 2020) approach. This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to extract an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message. Moreover, GPT-4 is prompted to give priority to the ‘expert-FAQ’ document–vetted by medical professionals–when it is among the top data chunks. The prompt also instructs that the response should adhere to WhatsApp’s message limit of 700 characters.
This prompt underwent several iterations. We engaged a doctor and a patient coordinator at the hospital to record questions posed by patients/attendants scheduled for cataract surgery over a week. A total of 153 questions were compiled. We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve the content of existing documents and contribute additional documents to enrich the KB. This process was iterated three times, to ensure comprehensive coverage and validate the effectiveness of our GPT-4 prompt with the RAG approach in extracting pertinent responses. Despite their effectiveness, we recognized the KB’s susceptibility to outdated or incomplete information. Hence, expert edits to bot’s answers, confirmed by the KB update expert, were consistently incorporated to regularly enhance and update the KB.
Additionally, we utilized GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 flawlessly executed this task.
Discussing language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, the CataractBot system adhered to the standard approach (ait, [n. d.]) of translating local Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective local Indic languages for output. Moreover, to facilitate speech as input and audio as output, CataractBot used Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.)
During testing, we encountered instances where GPT-4 was busy, resulting in response failures due to timeouts. To guarantee responses for every question, we implemented a queuing system. Each message from patients/attendants entered a queue and was removed only after receiving a response from GPT-4. Furthermore, to ensure expert verification, we employed multiple CronJobs for notifications and escalations. Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) gets logged in a cloud storage for further analysis.
4. Study Design
To evaluate the effectiveness of an LLM-powered expert-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 in collaboration with a local eye hospital, Sankara Eye Hospital. Sankara Eye Hospital is one of the leading tertiary eye care and teaching institutions in Bengaluru, India. It has 18 eye doctors (15 of whom perform cataract surgeries regularly) and over 10 optometrists. The hospital typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital.
4.1. Procedure for Patients and Attendants
As per the hospital’s protocol, once a patient elects to proceed with cataract surgery following the doctor’s recommendation, the patient and their attendant are directed by the doctor to meet a patient coordinator. The coordinator helps with general counseling, guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements. Toward the end of this interaction, the patient coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages (English, Hindi, Kannada, Tamil, or Telugu), (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as they would likely have minimal informational needs), and (e) being a patient of one of the three participating operating doctors. If the patient meets these criteria, the patient and their attendant are directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using CataractBot for around two weeks, engaging in two interviews, and completing a chatbot usability form. Upon confirming their willingness to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received a set of ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question to the bot, either by choosing from suggested questions (Figure 2F) or by typing/speaking a new question in their preferred language. Upon receiving a response, participants were briefed on the three icons (Figure 1E,F): question mark, green tick, and red cross. Throughout this process, the researcher encouraged them to clarify any doubts regarding the bot’s usage. Following this onboarding process, patients and their attendants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to ask questions to CataractBot.
Participants were instructed to use the bot from onboarding to one week post-surgery, after which their access was automatically revoked. Given that cataract surgeries were typically scheduled within a week of meeting with the patient coordinator, each patient/attendant was expected to use the bot for a maximum of two weeks. We conducted two interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery. These semi-structured interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. These specific days—surgery day and seven days post-surgery—were chosen because they coincided with the patient’s required hospital visits. After each interview, participants were requested to fill out a chatbot usability form. They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness. These adapted questions were sourced from metrics outlined in prior usability evaluations of healthcare bots (Abd-Alrazaq et al., 2020). Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while the exit interviews (seven days post-surgery) were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with surgery preparations and the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
4.2. Procedure for Doctors and Coordinators
In addition to recruiting patients and attendants, we recruited four doctors, each regularly performing cataract surgeries. Among them, three served as operating doctors, one as an escalation doctor and KB Update expert. Additionally, two patient coordinators were recruited, both specializing in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor, escalation coordinator, and KB update expert held senior positions in the hospital and carried administrative responsibilities. The KB update expert is part of the research team, and was actively involved in curating the knowledge base. This expert played a pivotal role in recruiting all other doctors and coordinators individually through one-on-one discussions.
To ensure preparedness, all participating doctors and coordinators attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and participants were onboarded onto the platform. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed several bugs, and feedback from doctors and coordinators was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in tagged WhatsApp messages (Figure 1E).
After using CataractBot for two weeks or longer, the participating doctors and coordinators were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes. None of the participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation.
4.3. Participants
A total of 31 patient-attendant pairs (Table 1), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate. One patient was not accompanied by an attendant. The average age of the 19 participating patients (3 females) was 58.88.01 years. They exhibited a diverse distribution in preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and highest education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 participating attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors (three as operating doctors, one as escalation doctor, and one as KB update expert) and two patient coordinators (one as coordinator and one as escalation coordinator) participated in the study (Table 2). Note: The roles of escalation doctor and KB update expert were performed by the same doctor.
| PId | Age | Sex | Language |
AId | Age | Sex | Language |
Days |
| A1 | 43 | M | English | Masters | Yes | Yes | 7 | 9 | |||||||||||||||
| A2 | 31 | F | Kannada | Grade 12 | Yes | Yes | 3 | 13 | |||||||||||||||
| P3 | 56 | M | Tamil | Grade 10 | A3 | 35 | M | English | Bachelors | Yes | No | 4 | 6 | ||||||||||
| P4 | 52 | M | English | Grade 12 | A4 | 39 | F | English | Grade 12 | Yes | No | 4 | 8 | ||||||||||
| P5 | 67 | M | Kannada | Grade 10 | A5 | 40 | M | Kannada | Bachelors | Yes | No | 5 | 4 | ||||||||||
| A6 | 35 | M | Telugu | Grade 10 | Yes | No | 1 | 2 | |||||||||||||||
| P7 | 58 | M | Telugu | Grade 10 | A7 | 25 | M | English | Masters | Yes | Yes | 23 | 17 | ||||||||||
| A8 | 39 | M | Hindi | Bachelors | Yes | No | 6 | 6 | |||||||||||||||
| P9 | 69 | M | English | Masters | A9 | 38 | F | English | Masters | Yes | Yes | 6 | 14 | ||||||||||
| A10 | 71 | M | English | Bachelors | Yes | Yes | 26 | 15 | |||||||||||||||
| A11 | 32 | M | Hindi | Grade 10 | Yes | No | 14 | 6 | |||||||||||||||
| P12 | 65 | M | Hindi | Bachelors | A12 | 31 | F | English | Masters | Yes | Yes | 2 | 6 | ||||||||||
| P13 | 69 | M | Tamil | Grade 12 | A13 | 35 | M | English | Masters | Yes | Yes | 25 | 18 | ||||||||||
| P14 | 56 | F | Kannada | Grade 10 | A14 | 33 | F | English | Bachelors | Yes | Yes | 22 | 30 | ||||||||||
| A15 | 57 | M | English | Grade 12 | Yes | No | 4 | 4 | |||||||||||||||
| P16 | 57 | F | Hindi | Grade 10 | A16 | 25 | M | English | Masters | Yes | No | 1 | 1 | ||||||||||
| A17 | 39 | M | English | Bachelors | Yes | Yes | 2 | 8 | |||||||||||||||
| P18 | 56 | M | English | Grade 12 | A18 | 23 | M | English | Masters | Yes | Yes | 27 | 24 | ||||||||||
| P19 | 41 | M | English | Masters | A19 | 36 | F | English | Bachelors | Yes | No | 19 | 15 | ||||||||||
| P20 | 64 | M | Tamil | Grade 12 | A20 | 60 | F | Tamil | Grade 12 | Yes | No | 4 | 16 | ||||||||||
| A21 | 43 | M | English | Masters | Yes | No | 15 | 19 | |||||||||||||||
| A22 | 35 | F | English | Bachelors | Yes | No | 7 | 11 | |||||||||||||||
| P23 | 54 | M | English | Bachelors | 7 | 8 | |||||||||||||||||
| P24 | 66 | M | Kannada | Bachelors | A24 | 37 | M | English | Masters | No | No | 1 | 1 | ||||||||||
| A25 | 40 | F | Tamil | Grade 10 | No | No | 7 | 3 | |||||||||||||||
| P26 | 65 | M | Hindi | Bachelors | A26 | 34 | M | English | Bachelors | No | No | 19 | 15 | ||||||||||
| P27 | 64 | M | Hindi | Bachelors | A27 | 35 | M | English | Bachelors | No | No | 18 | 14 | ||||||||||
| A28 | 40 | M | English | Masters | No | No | 1 | 1 | |||||||||||||||
| P29 | 45 | F | Kannada | Grade 10 | A29 | 29 | F | English | Bachelors | No | No | 15 | 10 | ||||||||||
| P30 | 50 | M | Hindi | Masters | A30 | 25 | F | English | Bachelors | No | No | 5 | 3 | ||||||||||
| P31 | 63 | M | Kannada | Bachelors | A31 | 32 | M | English | Bachelors | No | No | 15 | 36 |
| Id | Role | Age | Sex | Languages | Highest Education | Experience |
| D1 | Operating Doctor | 42 | F | English, Hindi, Kannada, Tamil, Telugu | Masters | 10+ years |
| D2 | Operating Doctor | 44 | F | English, Hindi, Kannada, Tamil, Telugu, Malyalam | Masters | 15+ years |
| D3 | Operating Doctor | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| D4 | Escalation Doctor; KB update expert | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| C1 | Coordinator | 33 | F | English, Hindi, Kannada, Tamil, Telugu | Bachelors | 10+ years |
| C2 | Escalation Coordinator | 36 | F | English, Hindi, Kannada, Tamil, Malyalam | Bachelors | 15+ years |
4.4. Data collection and analysis
We performed a mixed-method analysis to systematically analyze the collected data comprising of usage logs, interview transcripts, and usability survey responses. CataractBot interaction logs and usability survey data were quantitatively analyzed. Descriptive statistics and statistical tests (such as t-tests and ANOVAs) were used to analyze the count and type of questions asked, verification styles and response times of doctors and coordinators, edits performed by the KB update expert, and responses to the usability survey. Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied.
The first author translated and transcribed all interviews (totalling 11.5 hours) with patients, attendants, doctors and coordinators, into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and recent interviews, fostering a nuanced understanding of bot usage and practices while identifying emerging themes of interest. For the analysis of interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage”, “Features”, and “Privacy and Trust”) that we present in our findings below.
Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyzed their usage log data separately.
4.5. Positionality
All authors are of Indian origin, and currently reside in Bengaluru, India. Among them, two identify as female and the rest as male. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly. Their involvement spanned designing the study, shaping research questions, providing feedback on initial CataractBot versions, contributing to knowledge base curation, and serving as the KB update expert. Another author is affiliated with the Quality Control and Patient Safety team, who ensured that any risk of harm to the patients was mitigated. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in the Indian context. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot, guided our study design, and shaped our data analysis.
5. Findings
Overall, 31 patient-attendant222We use the terms ‘participants’ or ‘information-seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators. pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair. Among the 19 participating patients, only 12 interacted with the bot by sending one or more messages, while all 30 participating attendants, except one, messaged the bot. Patients sent a total of 89 messages (7.45.0 message/patient), while attendants sent 254 messages (8.86.8 message/attendant). Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others. Doctors directly approved 69.8% (157) of the bot-generated answers to medical questions and provided corrections for 29.3% (66) of the answers, while patient coordinators directly approved 58.6% (51) of logistical answers and provided corrections for 40.2% (35) of the answers. Note: Only two medical questions and one logistical question did not receive any response from the experts, while for eight questions (five medical and three logistical) an expert responded with ‘No’ to the poll without providing a correction. Below, in our findings, we delve into how, why, and when participants used CataractBot, examine the impact of novel features, and conclude by addressing privacy and trust concerns in the context of a chatbot in a medical use case.
5.1. CataractBot Usage
5.1.1. WHY
Patients and attendants used CataractBot for various reasons, such as seeking clarification on answers provided by the expert, addressing questions they had forgotten to ask the doctor or coordinator, inquiring about matters they felt less comfortable asking the expert directly, resolving emergency situations, and obtaining status updates. In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it (or a similar bot) in the future.
Ten participants highlighted the bot’s helpfulness in recalling information provided by the doctor/coordinator during in-person consultations, either due to forgetfulness or incomplete understanding. As A13 stated: “We are human minds… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed, “I need not remember everything.” (A20). The value of persistent information in the context of a text-based bot has been underscored in prior research (Jain et al., 2018). Despite knowing the answers to certain questions, six participants mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17).
Doctors highlighted that many patients/attendants now come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits to the hospital: “When you’re going to a surgery… there are ten things to ask. And when you are in front (of the doctor), you’ll remember only two things. Most of the patients nowadays come with a list of questions.” (D1). CataractBot has the potential to alleviate this burden of remembering every important question. Also, this flexibility of using CataractBot at one’s convenience–“as and when I get anything on my mind to ask” (P17)–was noted. The asynchronous nature of the conversation is beneficial, as information seekers do not have to constantly wait, either physically or over a call, for an expert’s response. A majority of participants (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 158.7168.4 minutes; we found verification without edits (151.1171.4 minutes) was faster than verification with edits (174.5161.4 minutes) which followed by 191.3172.4 minutes to provide the correction for the bot. This asynchronous feature not only benefited patients but also motivated experts to engage with the bot. The minimal time required to verify and edit an answer is another motivating factor.
“I can do it anytime… I do not have to think about phrasing and grammar… The bot is going to make it simpler and readable.” – D2.
Three participants noted the challenge of obtaining “precise information” from external sources like Google, prompting them to use the bot. Extracting accurate and relevant information from Google search results and formulating effective search queries is non-trivial (ter Hofstede et al., 1996), particularly for individuals with low (tech-)literacy. Moreover, due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), such individuals often hesitate to ask questions. Experts acknowledged that educated individuals have access to alternative information sources, and are comfortable posing questions. In contrast, “less educated, local language speaking patients” are the ones for whom CataractBot “can do wonders” (D3).
“They have lots of doubts… even if I talk to them in their native language, they hesitate to ask… because it sounds silly… Could be because of the lower socio-economic background, or education.” – D1.
Five participants, including three with Grade-12 education, exclusively relied on CataractBot for information (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online. Another expressed that inquiring with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature:
“(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” – A4.
Although explicitly mentioned during onboarding that CataractBot should not be used during emergencies, three patients/attendants attempted to use it to address emergencies, such as allergic reactions and taking the wrong medications. This raised concerns among two experts as CataractBot is not equipped to diagnose or handle emergencies. Expert responses are not immediate, and patient queries may lack the necessary details for the expert to provide an appropriate response without further clarification.
“If the patient has a little redness… and is supposed to take a flight… the bot may say “use the same medication no need to worry”, or it says “it’s an emergency come to the hospital, cancel the flight… In either scenario, it’s going to be difficult, because you don’t know what the redness is due to.” – D1.
The doctor suggested that the bot should be capable of detecting emergency situations, refrain from providing a response, and instead advise users to call the hospital.
Seven participants emphasized that the bot saved them time, as it eliminated the need to visit or call the hospital. Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as A4 stated:
“Yesterday she (coordinator) called… she told to come tomorrow, that’s all. I was about to ask a few questions but then she cut the call. Then I called back. It was busy, may be because she was talking to others.” – A4.
Despite the compelling reasons for using the bot, a few participants used it minimally. Out of the 49 participants, 9 did not use the bot at all, while 7 sent 1-2 messages to the bot. Six participants mentioned that their in-person interactions with the doctor and patient coordinator at Sankara Eye Hospital sufficed to address their queries, reducing their reliance on the bot. Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This perception could stem from their previous experience with cataract surgery, undertaken six months or more ago for their other eye. The patient coordinator added that abundant cataract surgery related information is readily available online, unlike specialised surgeries. This suggests that CataractBot has the potential to play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers. Additionally, a similar bot could prove valuable for more complex (surgical) procedures.
Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and/or experiencing delays in responses. A total of 36 (10.5%) messages pertained to requests for status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, lacking access to patient details and resulting in IDK responses for the non-generic questions. In such instances, the patient coordinator later provided appropriate responses. While participants acknowledged the appropriateness of schedule-related IDK responses due to the lack of access to patient details, for other queries, they suggested providing a “tentative response” rather than a blank IDK. For example, for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response, adhering to the cautious approach advised by hospital experts due to medical and legal concerns. However, A10 wanted a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context. E.g., A21 asked a complex, personalized, and high-stakes question: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check.”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor would know her condition… I don’t want to come and visit the doctor [again]… it will waste my 2-3 hours.” (A21).
For time-sensitive questions, A1 and A9 desired “quick response” from the bot. The delay in expert verification led them to perceive the bot as ineffective. “The dosage… every 2 hours take 1 drop. If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” (A1). This was reflected in the usability survey as well, with 27.2% of participants selecting ‘Disagree’ or ‘Strongly Disagree’ regarding the bot’s grasp of medical knowledge.
Experts used CataractBot driven by their sincere commitment to help patients. As D2 put it “I want to help my patients… I am least bothered about the timing. If he is having a problem, he should be answered… after all, we are living for the patient’s sake.”. Moreover, the asynchronous nature of the bot provided experts with the flexibility to offer verifications at their convenience. Finally, experts appreciated the intermediary role of the bot. The bot-in-the-middle model introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
5.1.2. HOW
The workflow employed by our patients and attendants varied, with time-sensitivity being a crucial factor. For time-insensitive questions, participants tended to consume the unverified response, wait for the expert’s verification, and later, consume the (updated) verified response. Our participants demonstrated a clear understanding of the bot’s workflow and expressed being “okay to wait” since the response was being verified by “busy doctors”.
“I know that it will take its own time… as its doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right? If that is the case, then bot is not required. They can directly reply.” – A17.
A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities. They would ask a question, leave their phones, and check them later when free. In such cases, notifications played a key role in the information-seeking workflow. The notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory. For questions where participants received an unsatisfactory IDK response, they sought information from other sources, including Google and asking friends and family.
CataractBot provides both an immediate LLM-generated response and an asynchronous expert-verified response. For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the inherent delay associated with obtaining a verified response. Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions. At times, they first used Google and subsequently posed the same question to the bot, hoping for a quick, verified response from the expert. In only two instances involving critical questions, participants relied on the bot’s response and took interim actions before getting verification from the expert due to the time-sensitive nature of the questions. According to D4, without CataractBot, patients might have resorted to Google, risking following inaccurate advice, given the unavailability of doctors for immediate consultations. However, CataractBot improves the likelihood of receiving accurate responses, as it is utilizing a custom knowledge base. Irrespective of time sensitivity, certain questions such as “What time i should come?” (P19), or “I get operated on 5th jan, can i do exercises as alom vilom etc?” (P27), could only be adequately answered by hospital experts. Participants relied on CataractBot for such questions.
We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4:
“It takes only about 15-20 seconds to answer each question… they are easy. It was not really eating into my time where I had like hundred messages [to respond] every day… It’s much easier than replying to my mailbox.” – D4.
None of the doctors adhered to a specific time for responding; they verified queries while awaiting the next patient, during lunch or coffee breaks, post-OPD towards the end of the day after 4 pm, and/or after reaching home. The only exception was during days reserved for surgeries. All doctors emphasized a no-phone-use policy during surgical procedures, emphasizing the need for minimal messages on those days. Even the coordinators suggested adding a working-time parameter for CataractBot, restricting message delivery to the specified period.
As the experts (except one coordinator) used CataractBot on their personal WhatsApp accounts, it blurred the boundary between their home and work (similar to (Mols and Pridmore, 2021; Wang et al., 2020)). On a positive note, WhatsApp minimized the learning curve for experts. However, our experts pointed out that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp. I just see messages in the important groups, that’s it. Otherwise my WhatsApp has 5000+ [unread] messages… college groups, family groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list. Additionally, experts managed multiple patients simultaneously through CataractBot. It could potentially compromise the quality of patient care, as observed by Johansson et al. (2020a). For instance, an expert mistakenly tagged the wrong query while adding a correction, leading to confusion for the patient.
Learning Curve: A majority of participants found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question. Several factors contributed to the ease of learning, with participants’ familiarity with WhatsApp being the most significant. Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), facilitated a quick start. Encouraging patients to use the bot as part of onboarding and providing detailed instructions proved essential in minimizing the learning curve. Contrary to initial expectations that understanding the three icons might be complex, participants found it self-explanatory:
“That is very simple… Even my dad understood. I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… Even for the tick mark also you are sending notification. If you would have done it in the background without a message, he would have not noticed… If MY dad was able to use, I think anyone can use, because even attending a phone call is a bit difficult for him.” – A13.
Explicit notification messages related to the experts’ verification step were appreciated by multiple participants.
5.1.3. WHAT
To conduct a more detailed analysis of the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Note: Several questions fell into multiple categories. Figure 4 illustrates the relative significance of these categories before, during, and after the day of surgery. Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia” and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”). One day before the surgery, with the coordinator finalizing the schedule, questions related to the “schedule” peaked, as patients/attendants awaited calls from patient coordinators. On the day of surgery, with patients starting a course of eye drops to prevent infection and reduce inflammation, questions about medication and dosage rose in importance, as did those about the expected discharge procedures. After the surgery, participants were particularly concerned about pain management and sought guidance on dos-and-don’ts during recovery, including information on exercise and travel.
Regarding question structure, our participants minimally used double-barreled questions—single messages containing multiple questions or topics—unlike prior work (Li et al., 2023). This difference may be attributed to demographic variations, as Li et al. (2023) focused on doctors who aimed to maximize productivity by asking multiple questions together. Only four double-barreled questions were sent by three of our participants, and in each instance, the two questions were related. E.g., “What is the Valsalva maneuver and why is it harmful after cataract surgery?” (A21). The bot answered both questions in each case.
Anthropomorphization: A total of 39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot. Participants often addressed the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’. Many queries were prefaced with greetings (like ‘hi’) and included acknowledgements (like ‘okay’), contributing to the humanization of the bot. In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews: “How should it be during the day? Please let me know.” – P14, in Kannada. Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as “The date of the surgery is tomorrow. We will leave home at 8 am from Ayyappa Nagar.” (P29). Such messages not only caused confusion for the bot but also for the experts. The classifier erred in not classifying these messages as small talk; instead, it specifically labeled them as medical to ensure that user queries were not overlooked. The observed behaviors can, in part, be attributed to participants’ awareness that doctors and patient coordinators are on the other side reviewing these messages. Participants attempted to direct their queries or provide status updates to them.
5.1.4. WHO
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=-2.04, p¡0.05. Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert. This highlights the significance of the bot supporting multiple information-seeker and information-verifier roles. Further, in two cases, patients came accompanied with multiple attendants, suggesting the need for the bot to accommodate not just a patient-attendant pair, but rather, a flexible number of information-seekers per patient.
During our interviews, we identified several reasons for the patients to use the bot minimally, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants – for example, 2 attendants, who were software testers by profession, mentioned that they “kept pushing some questions” (A21) to identify bugs and help improve the system. Moreover, a subset of patients enrolled on the bot (7) placed trust in the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (or know) about what will be this, what will be that”.” – A3.
This aligns with prior findings indicating less-educated older populations may have reduced information needs (Ma et al., 2023). Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced interaction with the bot.
Patients found value in local language interaction and speech as input. A13 and A14 noted that the bot’s Indic language support helped their parents (patients) to use the bot independently.
“Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
Quantitatively within patients, P13 sent the maximum number of messages to CataractBot at 16, comprising 15 text messages and 1 tap on a related question. The second-highest number of messages (15 messages) was sent by P19, the youngest patient at 41 years and a well-educated background holding a Master’s degree. Following that, P26 and P31 sent 13 and 9 messages respectively, with the majority being audio messages (8 and 6 audio messages, respectively).
Interestingly, the involvement of additional stakeholders had a negative impact, as patients tended to offload the knowledge-seeking ‘work’ to their active attendants. For instance, “I have been using it. I told him [patient] what (answers) I got. So maybe he did not feel the need (to use it).” (A20).
Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the 223 answered medical questions, 76.2% were verified by operating doctors, and 23.7% by the escalation doctor. In contrast, only 26.7% of the 86 answered logistical questions were verified by the coordinator, and 73.2% by the escalation coordinator. This discrepancy occurred because the patient coordinator participated in the study using their work phone, which was only active from 9 am to 6 pm, as they left their work phone at the hospital. Moreover, they were on emergency leave for 11 days during the study period. Doctors mentioned having 3 or more days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours. Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators. While the latter had previous face-to-face meetings with patients and attendants, there was no such guarantee for escalation experts. As a result, escalation experts lacked knowledge of patients’ context, and hence provided generic responses.
5.1.5. WHEN
Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour. Notably, the maximum number of messages (56) were sent between 10-11 am, despite participants receiving reminder messages from CataractBot at 7:30 am and 4 pm daily. The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.”.
The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery, for a maximum total of 15 days. However, due to a bug in the code to disable the bot seven days post-surgery, the bot remained active. Moreover, 3 participants had their surgeries rescheduled due to increased blood sugar levels and travel, leading them to use the bot for more than seven days before their surgery. Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 5). Four participants requested the bot to be available beyond the seven days post-surgery period. They cited that post-operative medications and precautions would continue for a month, and having the bot available during that time would serve as reassurance and align with the expected level of service from the hospital.
A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery. For the remaining days, across 49 users, the total number of messages was less than 10 messages/day. The highest number of messages in a single day were sent on the day of the surgery (76), and 156 messages sent before the surgery day and 111 post the surgery day. We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05. Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05). This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety. This finding aligns with prior research indicating that patients and attendants tend to seek information closer to their doctor’s visit (Beisecker and Beisecker, 1990). The higher pre-surgery usage may also be due to the novelty effect. With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the scheduled surgery day (Figure 4), as anticipated. Despite the high pre-surgery usage, four participants mentioned that since they did not have access to CataractBot in the early stages of diagnosis, before scheduling the surgery, information-seeking patients and attendants had already clarified most of their questions from other sources, including Google, friends and family, and hospital staff. Particularly, three patients were scheduled for surgery in the next 2 days, so their pre-surgery usage was limited to less than 2 days.
In contrast, experts expressed the view that questions on the day of surgery might be unnecessary and suggested encouraging patients/attendants to talk directly with doctors and coordinators face-to-face:
“There are multiple questions asked on the day of surgery, “when he is going to be discharged?” Asking a question when you are in the hospital itself, looks a little too much… There is no need for you to use the bot… absolutely avoidable. You can walk 10 steps and go to the staff and ask.” – D1.
Regarding the usage patterns of experts, responses were verified throughout the day and even late at night (1 am to 2 am). The majority of responses (31) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. Also, the reminder message was sent to the experts at 4 pm. Note: Every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours was sent to both the operating and escalation experts. The second-highest responses (28) occurred at 1-2 pm, during their lunch break. This time-of-day usage pattern was described as:
“I use my mobile data only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
Experts refrained from using their phones in front of patients/attendants due to the negative perception associated with using phones for personal reasons. As C1 described: “If I see my phone (in front of a patient)… they will not understand what I am doing, they will think okay, simply they are watching videos or texting.”. In a busy hospital context, physical patients always took priority. Hospitals in India often face high patient loads. Experts are typically overwhelmed with more patients than they can handle. During working hours, there is usually a long queue of patients waiting to consult with experts, making it challenging for experts to use their phones comfortably to verify bot responses. This also contributed to many medical and logistical queries being escalated. We observed varying opinions on using the chatbot after work hours; two experts were reluctant to use the bot outside of working hours, while four were in favor of using it “anytime”.
5.2. Features
5.2.1. Expert-in-the-loop
Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others. Note: 13 questions were marked as ‘Send to Coordinator’ by one of the operating doctors. On manual analysis, the escalation doctor confirmed that all of them were indeed medical questions, such as “after how many days can I shave”, but were simple enough to be answered by the patient coordinator. For our analysis, we consider them as logistical questions, as they were finally answered by the patient coordinator. Doctors approved 69.8% (157) of the bot’s generated answers directly and provided corrections for 27.1% (61) of responses, while patient coordinators approved 58.6% (51) of responses to logistical questions directly and provided corrections for 36.8% (32) of responses. For 5 medical queries and 3 logistical queries, the expert answered the poll response with a ”No” but did not provide the correction. The high number of logistical corrections was mainly because a majority of logistical questions (62.1%) were about scheduling, such as “time of surgery” and “discharge time”, for which the LLM cannot provide an appropriate response as it was not connected with the hospital’s patient management system. The length of these corrections in characters varied with 154.4105.0 for medical questions and 51.042.6 for logistical questions.
Despite having a curated knowledge base and an expanding knowledge base with the addition of expert corrections after KB Update Expert approval, the bot provided IDK responses to 16.6% of questions. Even after excluding scheduling-related questions, 23.9% of responses were corrected by the expert. Further analysis revealed that their suggested edits across all questions mainly fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), redirecting the patient to check with other sources (8.8%), and removing extra unnecessary information (3.7%).
The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by our participants, serving as the key reason for their engagement with the bot.
“That’s why I gave my WhatsApp (number to the coordinators while onboarding). Because I can talk directly to the doctor.” – A2.
However, since participants were aware that these messages were being reviewed by doctors and coordinators, four participants used the bot to send direct messages to them instead of asking questions. For instance, P19 messaged “Give me call?”
Experts did not have access to the final updated answer sent to patients/attendants after their suggested edits. We made this design decision for CataractBot to limit verification to a single round, minimizing the workload on experts. However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understand and give that as an answer? I did not know.” (D3). In the CataractBot system, only the KB Update Expert had access to the updated answer sent to the information seeker.
Finally, asynchronous expert correction may not be effective in all instances, especially when patients ask critical time-sensitive questions, such as allergic reactions after surgery. In such cases, two patients suggested having a fallback option of a human agent available over a call to assist: “similar to Amazon… if the query is not resolved after two-three attempts, it should just ask you to contact [call] the coordinator.” (P21). Even the escalation doctor D4 suggested a similar solution. However, prior work (Weisz et al., 2019) suggests that in such solutions, users tend to bypass the automated agent to directly connect with the human agent over a voice call.
5.2.2. Input Modality
Patients and attendants were enabled with three ways to interact with CataractBot: typing messages in their preferred language, recording audio, or tapping one of the three related questions suggested by the bot. However, experts could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023). Out of the total 343 messages sent by patients and attendants, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions. The 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 245 messages (163 text, 17 audio, and 74 taps). One-way ANOVAs on message types (after Greenhouse-Geisser correction for sphericity violations) found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01), ‘text’ and ‘taps’ (t(29)=3.0, p¡.01), and not for ‘audio’ and ‘taps’ (t(29)=1.9, p=.11). The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12). Moreover, post-surgery limitations on screen use made speech input and audio output the preferred modality. 34 participants did not use the audio modality at all, while most participants (31) selected one of the related questions. As P7 stated: “I used the audio message once. After that there was no need for me… I just selected from the options. Click here (shows the phone) and automatically question will come, and click one question and get the answer.”. While patients and attendants found the related questions feature convenient and frequently used it, there were instances of dissatisfaction. On eight occasions, the answers to suggested questions were not available in the knowledge base, resulting in IDK responses. This issue arose because the CataractBot system use an LLM call to suggest three related questions based on the users’ last question, independent of the KB. This approach aimed to uncover gaps in the KB by not restrict questions to the curated KB, and introduce a personalized touch to the interaction by suggesting questions relevant to the ongoing conversation.
5.2.3. Multilinguality
Experts interacted with CataractBot only in English. Among our patients and attendants, the majority (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. The inclusion of four Indic languages, along with speech input functionality, ensured effective use of the bot by users with minimal literacy and tech proficiency. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. Patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and selected 14 related questions. A similar trend was observed among attendants.
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, due to variability in dialect across India, existing transcription and translation technologies have been found to be erroneous. It is a known issue and prior work (Jain et al., 2018) often relied on human wizards to correct errors caused by language technologies. For example, a query by P29 (Kannada) was transcribed and translated to “Krishna immediately said, how long is this kampu?”. D3, upon requesting and listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately. Experts faced challenged in “deciphering” such messages since they only received the English text, which was sometimes nonsensical. They suggested that they should be provided with the original audio recordings of patient queries. D4 further recommended that including audio recordings would help extract non-verbal cues (such as anxiety or the presence of other family members), enabling the expert to respond more appropriately.
It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot. We found three participants who sent a text message each by typing an Indic language query using Latin script. For one of them, the bot responded with an accurate response because the message contained a code-mix of Hindi and English:
A8: agar opration k baad pain ho raha to kya karna hai?
Bot: If the patient is experiencing pain after the surgery, they should report to the doctor immediately…
For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…” and the IDK response. Additionally, even in the selected language, participants with low literacy and language proficiency struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)). For example, “Tomorrow cataract operation When should come No phone call received yet?” (P18). Such questions were common with 65 questions, of which 26 were asked using audio in an Indic language. Further analysis revealed that the speech-to-text engine was mainly responsible for these errors (19 out of 26) due to the different dialects used by the users, while the remaining 7 were translation errors.
Despite these challenges, the robustness of the GPT-4 model enabled the chatbot to answer most (32) of such ill-formulated questions appropriately. For instance, in response to the question asked by P18, the bot replied, “The surgical list is usually prepared the previous evening. If you haven’t received a call from the inpatient department, please call and confirm your timing.”
5.2.4. KB Updation
Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB Update expert (D4) to be added to the knowledge base. In cases where the expert edits were accepted, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB Update expert in 49.3% of cases. These additional changes by D4 extended beyond information-related aspects, also addressing aspects of “tone and structure” (such as preferring active over passive voice). The changes made by the KB Update expert were minimal, with an average relative Hamming distance of 28.2%. This hints that the bot effectively incorporated the expert edit suggestions into the final bot response.
All our experts noticed an improvement in the completeness and accuracy of CataractBot’s generated answers over time, resulting in a reduction of their workload: “With time, I think the answers did get better… I am mostly just validating and saying yes.” (D3). However, there was an instance where the KB Update expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the KB. This led CataractBot provide incorrect information to three users. This error was manually corrected in the KB by a researcher on the team. Despite efforts to expand the KB, we identified limited knowledge in three medical areas—types of lenses and their comparisons, lab tests and prerequisites, and details related to prescribed medicines—and one logistical area—insurance tie-ups. Addressing it requires addition of new documents to the custom KB. Hence, D4 suggested the need for a simple way to upload documents to further expand the KB.
5.3. Privacy and Trust
5.3.1. Personalization vs Privacy
In the context of evolving LLM technology, there is an ongoing discussion about privacy concerns, as LLMs have been shown to memorize training data (including personally identifiable information (PII) like emails and phone numbers) and leak it during inference (Borkar, 2023). To address these concerns, CataractBot is intentionally designed to be non-specific to individual patients, thus preserving patient privacy. A few participants (3) appreciated that their medical records were private.
“Using personal health information is… violation of the patient’s personal rights. See, if that (health data) is added… external agencies can extract that information and they might use it for some advertising purposes… Personal information or any diagnosis related to any patient should not be shared.” – A17.
Three educated participants (holding Master’s degrees) considered personalization as “an unnecessary add-on”. They raised concerns about potential complications associated with sharing PII data with LLMs, both in terms of legislation and “privacy threads”. A1, e.g., remarked “I would definitely not want my mom’s data to be out there, all over the internet.”. Moreover, the lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two three people I know are going to have their cataract surgery soon… I’ll use it for them.” (A11).
Since CataractBot neither had access to patient medical records nor was integrated with the hospital’s patient management system, it was not privy to patient-specific medical and logistical details, which resulted in IDK responses for 20 queries related to surgery schedules, updates during the procedure, and medication dosage pre-/post-surgery. During interviews, five participants expressed frustration with the IDK responses, and two of them discontinued using the chatbot after encountering a few such responses.
Participants expected personalized answers to their queries and were disappointed with generic responses. There existed an expectation mismatch, fueled by the recent hype and promises surrounding LLMs. Participants anticipated more advanced capabilities from such a system. In contrast, a participant working in the field of AI had a more balanced perspective: “See, I understand, it is just a machine… It cannot surpass or bypass or completely replace humans.” (A17). Another participant, A21, highlighted that providing more information in the chat messages could improve the bot’s ability to deliver a more personalized response. However, they acknowledged the challenge in “setting context every time”.
Eight participants asked for personalized responses from the chatbot, with little concern for privacy. They felt that a personalized chatbot could save time and energy, eliminating the need to seek information from other sources or visit the hospital. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s like nothing personal, to be hidden or something. All have common health issues.” – A14. Moreover, two participants, due to the low literacy and tech-literacy, questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
Educated participants were willing to share their personal medical data if the benefits outweighed the associated costs.
“If you’re traveling to Dubai and you say “I won’t show you my passport”… they will say “Thank you so much, you can’t board the plane.”… [Similarly] if you don’t want to share YOUR information [to the bot], the bot should remark “If you ask any personal questions the bot will not have access so you may have to call or visit the doctor or hospital.”… somewhere we have to compromise.” – A21.
Additionally, two participants mentioned that personalization, along with the sharing of patient medical data, should be an “opt-in” feature. Individuals should conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized CataractBot would offer more relevant responses and minimize IDK responses.
With each query, CataractBot provides experts with patient details such as name, age, gender, and surgery date (Figure 1A). D3 found this feature “(useful) while giving post-operative information, because (we) won’t remember patient metadata”. Four doctors and patient coordinators were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses. Such integration would not only reduce their workload by addressing medication and logistic-related questions accurately, but it will also enhance user trust. D3 suggested collecting at least basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “Should I eat and come?”… I don’t know what to tell.” (D3). Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of the participants provide detailed information due to time constraints. C2 suggested integrating the bot with the hospitals’ EHR (Electronic Health Record) system. However, two doctors (D1 and D2) voiced concerns about CataractBot accessing patient data, citing privacy issues and skepticism about the LLMs capability to understand complex patient records accurately. “There are a lot of variables involved, including the surgeon, the particular patient, the particular case, the tests… Its too much information to process.” (D1). In addition, the KB Update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for all experts and negatively impacting chatbot adoption among experts.
5.3.2. Bot-driven vs User-driven
CataractBot operates as a user-driven platform from the perspective of patients and attendants, while it functions as a bot-driven system for experts. Information seekers initiate conversations by asking questions and receive responses. In contrast, the bot initiates and drives verification-related conversations with experts. Patients and attendants appreciated that CataractBot “does not disturb” them, providing information solely based on the questions asked. However, a few participants suggested the bot should proactively offer relevant information about the surgery. Specifically, four attendants wished regular updates from the chatbot during the surgery, particularly while waiting anxiously outside for the procedure to conclude. Similarly, expert D4 suggested that the chatbot should notify patients: “Your discharge summary is ready, you can come and pick it up.”. A9 recommended reminders from the chatbot to ensure patients “take eye drops timely” post-surgery.
5.3.3. Reliability and Trust
Six patients/attendants explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery. As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2). This contributed to participants’ trust in the bot. Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot.
While the hospital’s reputation instilled trust in the bot, participants gained confidence in the bot’s responses due to the human experts verifying answers, denoted by the green tick icon on verified responses (Figure 1F). As A12 stated: “Basically the trust only came when we saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”. This ensured that the bot’s response were more reliable and “not fake” (P18), in contrast to a few participants who complained about receiving false information from Google searches.
Although the expert-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence. Four participants noted that every time an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers (Figure 2E) reduced. For instance, “After some point when doctors said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert. This could be due to the bot’s limited custom KB or the prompting strategy favoring caution, as per doctors recommendation. A9 pointed: “It said that severe cough is dangerous on the surgery day as it increases eye pressure. But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”.
5.3.4. Accountability
CataractBot provides a persistent written record of expert-patient communication, carrying several implications. Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Moreover, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can use the same information to be shared with their doctor at their place, “Okay, this is what was told to us.” So that [communication] becomes a little easier.” (D1). However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4). Furthermore, coordinators identified the system’s potential to identify logistical bottlenecks in the hospital. For example, if multiple patients are inquiring about their surgery time, it may indicate a lapse in timely communication, leading them to reach out to the bot.
6. Discussion
In this paper, we propose CataractBot, an LLM-powered WhatsApp-based expert-in-the-loop conversational agent designed to address the information needs of patients undergoing cataract surgery. Our user study revealed positive evidence of its usability among different stakeholders, including patients, attendants, doctors, and coordinators. Patients and attendants appreciated CataractBot for its ability to answer questions anytime and provide verified responses, contributing to the establishment of trust. Additionally, it streamlined communication, saving time by reducing the need for hospital calls/visits. Notably, individuals with low literacy and low tech literacy found it easy to use because of the robust multilingual and multimodal support. Experts commended CataractBot for acting as a facilitator, creating a layer between experts and patients. The asynchronous communication feature provided them the flexibility to verify LLM-generated responses at their convenience.
We discuss key design implications below.
6.1. Design Implications
6.1.1. Personalization
Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information (Yang et al., 2023), they may also raise privacy concerns among certain users, potentially reducing usage. Additionally, personalized responses may increase the workload for experts, as they would need to refer to patient-specific documents for verification purposes. Many of these verification edits by experts may not be suitable for inclusion in the KB due to there patient-specific nature. Given our demographic of individuals with low literacy and low tech-literacy, individuals might require assistance in making an informed decision about consenting to the use of their health data (Okolo et al., 2021). Also, adopting a binary approach–either sharing all medical records or none–may not align with user preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access, in line with the philosophy of the National Digital Health Mission (NDHM) (National Health Authority and Family Welfare, 2020) in India.
To achieve this, a potential solution could be to ask users to manually upload images/PDFs of relevant documents (such as consultation sheets, discharge summaries, and test reports) they wish the bot to access. This approach empowers users to control the information available to the chatbot for a more personalized response. Currently, CataractBot operates as a reactive, user-driven system. However, with personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s treatment journey, providing advice on comorbidities and medication reminders.
6.1.2. Reduce Workload for Experts
Deploying CataractBot in the hospital increased the workload for patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates). To alleviate their burden, we propose ways to streamline CataractBot-related tasks and to use CataractBot to support them in their day-to-day activities. First, to simplify patient onboarding, we propose installing QR codes with a concise CataractBot introduction at key locations in the hospital. Users can scan the code, self-onboard by providing necessary information (language preference, operating doctor, and surgery date), and instantly engage with the bot. Users should receive a brief tutorial video on interaction guidelines, emphasizing that the bot is not intended for emergencies. Second, CataractBot should be utilized to deliver general information about the surgery and provide status updates to patients and attendants, thereby assisting experts in managing their other responsibilities. Supporting routine procedures, as recommended by Li et al. (2023), would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time. Third, to reduce the verification load for experts, we propose a system to pre-verify frequently asked questions automatically based on experts’ repeated confirmations of correctness. Fourth, we recommend introducing a feature that allows experts to ‘pause’ CataractBot by specifying their ‘unavailable until ¡date time¿’ status in natural language. This not only helps prevent disturbances for experts during surgery days, outside working hours, or when on leave, but also ensures that information seekers are better informed of the experts’ availability, also suggested by Yang et al. (2023). Finally, regarding audio messages, the lack of non-verbal cues hindered experts’ ability to deliver care (similar to (Li et al., 2023)). To address the limitations of translation and transcription technologies, the information-seeker’s original audio clip and transcribed query should be shared with the experts.
6.1.3. Usefulness
Prior work found that patients often seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention (Attfield et al., 2006). However, access to CataractBot is currently provided post-scheduling of the surgery. Introducing the bot before surgery scheduling might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts. For consultation-related information needs, a similar bot connected with the user’s family doctor could be enabled to address generic medical information needs. Based on our findings, we recommend keeping CataractBot enabled for a month post-surgery. During bot interactions, a few participants expressed frustration with IDK responses (similar to previous findings (Xiao et al., 2023)), leading to discontinuation of usage. To address this, we suggest enhancing IDK responses by providing general information about the topic, offering tentative recommendations with disclaimers, and guiding users to alternative information sources. Lastly, considering the visual capabilities of recent LLM models like GPT-4V (a multimodal model), CataractBot should be enhanced to handle image-based inputs, aligning with recommendations from previous work (Jain et al., 2018). For instance, patients could upload images of medications instead of providing written details. While this could improve the user experience, it may also increase the workload for experts.
6.1.4. Scaling
In this work, the use of CataractBot was limited to a small sample of 49 patients and attendants over 2 months, while Sankara Eye Hospital performs over 50 cataract surgeries daily. Scaling CataractBot to cater to all cataract patients would require several modifications. These include implementing a QR code-based self-onboarding system, establishing a more comprehensive custom knowledge base from the outset, streamlining the verification process involving multiple doctors and coordinators for each patient (considering the potential non-participation of some experts), and improving the KB update process by engaging multiple KB Update Experts. Even within our short deployment, we noticed the knowledge base becoming outdated and the KB Update expert adding incorrect information. For longitudinal deployments, we recommend engaging multiple hospital staff in managing the KB to reduce reliance on a single expert, and implementing a simple method for the staff to add new documents to the KB. Recognizing the impracticality of manual updates in rapidly changing environments (Xiao et al., 2023), we propose automatic updates for specific information sources, such as standard operating procedures.
Cataract surgery, being short, straightforward, commonplace, highly successful, and extensively documented, resulted in minimal queries from participants to CataractBot. A chatbot similar to CataractBot could provide more value to information seekers undergoing complex and less common treatments, like glaucoma surgery or radiation therapy for head and neck cancer. With the existing framework for constructing experts-in-the-loop bot, creating such bots should be feasible333We will open-source our codebase to help the research community build and deploy their own end-to-end expert bots.. Furthermore, the applicability of such a system need not be confined to the medical domain; other fields, such as legal or agriculture (Jain et al., 2018), could benefit from similar chatbots. In these scenarios, end-users (clients or farmers) could receive synchronous responses from an LLM operating on a custom knowledge base and subsequently obtain verified responses asynchronously from experts (lawyers or agricultural experts).
6.2. Limitations
We acknowledge several limitations of this work. First, the results should be interpreted considering the specificities of the study, including the relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. The positive user responses and other conclusions should be validated by future, preferably longitudinal studies. Second, we need to conduct system-level analysis, such as scalability testing, power measurement, etc., in future work before scaling the proposed solution to the masses. Third, CataractBot uses responses from the expert-in-the-loop after approval by the KB Update expert to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB. This issue needs to be resolved and requires further investigation. Finally, although participants showed high trust in the initial encounter, IDK responses and inaccurate information in daily usage can lead to distrust of the system. We also note the potential risk of disseminating inaccurate medical information at scale with chatbot technologies. Therefore, the knowledge base should be carefully developed, reviewed, and updated by experts.
7. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. To tackle this, we propose a novel solution—a LLM-powered expert-in-the-loop chatbot framework, that utilizes RAG (Retrieval Augmented Generation) over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, squeezed time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
Acknowledgements.
Many thanks to all the participants for their time and patience.References
- (1)
- ait ([n. d.]) [n. d.]. An AI service for real-time document and text translation. https://www.microsoft.com/en-us/ai/ai-translator Accessed: 2024-02-01.
- Abd-Alrazaq et al. (2020) Alaa Abd-Alrazaq, Zeineb Safi, Mohannad Alajlani, Jim Warren, Mowafa Househ, and Kerstin Denecke. 2020. Technical metrics used to evaluate health care chatbots: scoping review. Journal of medical Internet research 22, 6 (2020), e18301.
- Ahuja et al. (2023) Kabir Ahuja, Harshita Diddee, Rishav Hada, Millicent Ochieng, Krithika Ramesh, Prachi Jain, Akshay Nambi, Tanuja Ganu, Sameer Segal, Maxamed Axmed, Kalika Bali, and Sunayana Sitaram. 2023. MEGA: Multilingual Evaluation of Generative AI. arXiv:2303.12528 [cs.CL]
- Attfield et al. (2006) Simon J. Attfield, Anne Adams, and Ann Blandford. 2006. Patient information needs: pre- and post-consultation. Health Informatics Journal 12, 2 (2006), 165–177. https://doi.org/10.1177/1460458206063811 arXiv:https://doi.org/10.1177/1460458206063811 PMID: 17023406.
- Au Yeung et al. (2023) Joshua Au Yeung, Zeljko Kraljevic, Akish Luintel, Alfred Balston, Esther Idowu, Richard J Dobson, and James T Teo. 2023. AI chatbots not yet ready for clinical use. Frontiers in Digital Health 5 (2023), 60.
- Beisecker and Beisecker (1990) Analee E. Beisecker and Thomas D. Beisecker. 1990. Patient Information-Seeking Behaviors When Communicating with Doctors. Medical Care 28, 1 (1990), 19–28. http://www.jstor.org/stable/3765618
- Bickmore et al. (2009) Timothy W Bickmore, Laura M Pfeifer, and Brian W Jack. 2009. Taking the time to care: empowering low health literacy hospital patients with virtual nurse agents. In Proceedings of the SIGCHI conference on human factors in computing systems. 1265–1274.
- Borkar (2023) Jaydeep Borkar. 2023. What can we learn from Data Leakage and Unlearning for Law? arXiv preprint arXiv:2307.10476 (2023).
- Byrne and Heyman (1997) G Byrne and R Heyman. 1997. Patient anxiety in the accident and emergency department. J. Clin. Nurs. 6, 4 (July 1997), 289–295.
- Cajander and Grünloh (2019a) Åsa Cajander and Christiane Grünloh. 2019a. Electronic Health Records Are More Than a Work Tool: Conflicting Needs of Direct and Indirect Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300865
- Cajander and Grünloh (2019b) Åsa Cajander and Christiane Grünloh. 2019b. Electronic Health Records Are More Than a Work Tool: Conflicting Needs of Direct and Indirect Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300865
- Christalle et al. (2019) Eva Christalle, Jördis Zill, Wiebke Frerichs, Martin Härter, Yvonne Nestoriuc, Jörg Dirmaier, and Isabelle Scholl. 2019. Assessment of patient information needs: A systematic review of measures. PLOS ONE 14 (01 2019), e0209165. https://doi.org/10.1371/journal.pone.0209165
- Civan-Hartzler et al. (2010) Andrea Civan-Hartzler, David W. McDonald, Chris Powell, Meredith M. Skeels, Marlee Mukai, and Wanda Pratt. 2010. Bringing the field into focus: user-centered design of a patient expertise locator. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 1675–1684. https://doi.org/10.1145/1753326.1753577
- Clarke et al. (2016a) Martina A Clarke, Joi L Moore, Linsey M Steege, Richelle J Koopman, Jeffery L Belden, Shannon M Canfield, Susan E Meadows, Susan G Elliott, and Min Soon Kim. 2016a. Health information needs, sources, and barriers of primary care patients to achieve patient-centered care: A literature review. Health Informatics Journal 22, 4 (2016), 992–1016. https://doi.org/10.1177/1460458215602939 arXiv:https://doi.org/10.1177/1460458215602939 PMID: 26377952.
- Clarke et al. (2016b) Martina A Clarke, Joi L Moore, Linsey M Steege, Richelle J Koopman, Jeffery L Belden, Shannon M Canfield, Susan E Meadows, Susan G Elliott, and Min Soon Kim. 2016b. Health information needs, sources, and barriers of primary care patients to achieve patient-centered care: A literature review. Health Informatics Journal 22, 4 (2016), 992–1016. https://doi.org/10.1177/1460458215602939 arXiv:https://doi.org/10.1177/1460458215602939 PMID: 26377952.
- Dahlén and Janson (2002) Inger Dahlén and Christer Janson. 2002. Anxiety and depression are related to the outcome of emergency treatment in patients with obstructive pulmonary disease. Chest 122, 5 (Nov. 2002), 1633–1637.
- Duffy et al. (2004) F Duffy, Geoffrey Gordon, Gerald Whelan, Kathy Cole-Kelly, Richard Frankel, Natalie Buffone, Stephanie Lofton, MaryAnne Wallace, Leslie Goode, and Lynn Langdon. 2004. Assessing Competence in Communication and Interpersonal Skills: The Kalamazoo II Report. Academic medicine : journal of the Association of American Medical Colleges 79 (07 2004), 495–507. https://doi.org/10.1097/00001888-200406000-00002
- Eriksson-Liebon et al. (2021) Magda Eriksson-Liebon, Susanne Roos, and Ingrid Hellström. 2021. Patients’ expectations and experiences of being involved in their own care in the emergency department: A qualitative interview study. Journal of clinical nursing 30, 13-14 (2021), 1942–1952.
- Fadhil (2018) Ahmed Fadhil. 2018. A conversational interface to improve medication adherence: towards AI support in patient’s treatment. arXiv preprint arXiv:1803.09844 (2018).
- Ferré et al. (2020) Fabrice Ferré, Nicolas Boeschlin, Bruno Bastiani, Adeline Castel, Anne Ferrier, Laetitia Bosch, Fabrice Muscari, Matt Kurrek, Olivier Fourcade, Antoine Piau, and Vincent Minville. 2020. Improving Provision of Preanesthetic Information Through Use of the Digital Conversational Agent “MyAnesth”: Prospective Observational Trial. J Med Internet Res 22, 12 (4 Dec 2020), e20455. https://doi.org/10.2196/20455
- Findlater et al. (2009) Leah Findlater, Ravin Balakrishnan, and Kentaro Toyama. 2009. Comparing Semiliterate and Illiterate Users’ Ability to Transition from Audio+text to Text-Only Interaction. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Boston, MA, USA) (CHI ’09). Association for Computing Machinery, New York, NY, USA, 1751–1760. https://doi.org/10.1145/1518701.1518971
- Jain et al. (2018) Mohit Jain, Pratyush Kumar, Ishita Bhansali, Q. Vera Liao, Khai Truong, and Shwetak Patel. 2018. FarmChat: A Conversational Agent to Answer Farmer Queries. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 2, 4, Article 170 (dec 2018), 22 pages. https://doi.org/10.1145/3287048
- Jo et al. (2023) Eunkyung Jo, Daniel A Epstein, Hyunhoon Jung, and Young-Ho Kim. 2023. Understanding the benefits and challenges of deploying conversational AI leveraging large language models for public health intervention. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–16.
- Johansson et al. (2020a) Anders Johansson, Monica Larsson, and Bodil Ivarsson. 2020a. General practitioners’ experiences of digital written patient dialogues: a pilot study using a mixed method. Journal of Primary Care & Community Health 11 (2020), 2150132720909656.
- Johansson et al. (2020b) Anders Johansson, Monica Larsson, and Bodil Ivarsson. 2020b. Patients’ experiences with a digital primary health care concept using written dialogues: a pilot study. Journal of Primary Care & Community Health 11 (2020), 2150132720910564.
- Joseph-Williams et al. (2014) Natalie Joseph-Williams, Glyn Elwyn, and Adrian Edwards. 2014. Knowledge is not power for patients: A systematic review and thematic synthesis of patient-reported barriers and facilitators to shared decision making. Patient Education and Counseling 94, 3 (2014), 291–309. https://doi.org/10.1016/j.pec.2013.10.031
- Joshi et al. (2011) Anirudha Joshi, Girish Dalvi, Manjiri Joshi, Prasad Rashinkar, and Aniket Sarangdhar. 2011. Design and evaluation of Devanagari virtual keyboards for touch screen mobile phones. In Proceedings of the 13th International Conference on Human Computer Interaction with Mobile Devices and Services. 323–332.
- Kinnersley et al. (2007) P Kinnersley, A Edwards, K Hood, N Cadbury, R Ryan, H Prout, D Owen, F Macbeth, P Butow, and C Butler. 2007. Interventions before consultations for helping patients address their information needs. Cochrane Database Syst. Rev. 2010, 3 (July 2007), CD004565.
- Klasnja et al. (2010a) Predrag Klasnja, Andrea Civan Hartzler, Kent T. Unruh, and Wanda Pratt. 2010a. Blowing in the Wind: Unanchored Patient Information Work during Cancer Care. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 193–202. https://doi.org/10.1145/1753326.1753355
- Klasnja et al. (2010b) Predrag Klasnja, Andrea Civan Hartzler, Kent T. Unruh, and Wanda Pratt. 2010b. Blowing in the wind: unanchored patient information work during cancer care. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 193–202. https://doi.org/10.1145/1753326.1753355
- Kumar et al. (2023) Harsh Kumar, Yiyi Wang, Jiakai Shi, Ilya Musabirov, Norman AS Farb, and Joseph Jay Williams. 2023. Exploring the Use of Large Language Models for Improving the Awareness of Mindfulness. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–7.
- Kurtz (2002) Suzanne M. Kurtz. 2002. Doctor-Patient Communication: Principles and Practices. Canadian Journal of Neurological Sciences / Journal Canadien des Sciences Neurologiques 29, S2 (2002), S23–S29. https://doi.org/10.1017/S0317167100001906
- Larasati et al. (2023) Retno Larasati, Anna De Liddo, and Enrico Motta. 2023. Meaningful Explanation Effect on User’s Trust in an AI Medical System: Designing Explanations for Non-Expert Users. ACM Trans. Interact. Intell. Syst. 13, 4, Article 30 (dec 2023), 39 pages. https://doi.org/10.1145/3631614
- Leong et al. (2018) Zeina Atrash Leong, Michael S. Horn, Lisa Thaniel, and Emily Meier. 2018. Inspiring AWE: Transforming Clinic Waiting Rooms into Informal Learning Environments with Active Waiting Education. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3173672
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 9459–9474. https://proceedings.neurips.cc/paper_files/paper/2020/file/6b493230205f780e1bc26945df7481e5-Paper.pdf
- Li et al. (2023) Brenna Li, Tetyana Skoropad, Puneet Seth, Mohit Jain, Khai Truong, and Alex Mariakakis. 2023. Constraints and Workarounds to Support Clinical Consultations in Synchronous Text-based Platforms. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 342, 17 pages. https://doi.org/10.1145/3544548.3581014
- Lotfi et al. (2019) Mojgan Lotfi, Vahid Zamanzadeh, Leila Valizadeh, and Mohammad Khajehgoodari. 2019. Assessment of nurse–patient communication and patient satisfaction from nursing care. Nursing open 6, 3 (2019), 1189–1196.
- Ma et al. (2023) Xiumei Ma, Yunxing Liu, Pengfei Zhang, Rongtao Qi, and Fanbo Meng. 2023. Understanding online health information seeking behavior of older adults: A social cognitive perspective. Frontiers in Public Health 11 (2023), 1147789.
- Moden (2022) Rebecca Moden. 2022. Blank Page Syndrome And How To Beat It. Retrieved Sept 1, 2023 from https://www.ool.co.uk/blog/blank-page-syndrome-and-how-to-beat-it/
- Mols and Pridmore (2021) Anouk Mols and Jason Pridmore. 2021. Always available via WhatsApp: Mapping everyday boundary work practices and privacy negotiations. Mobile Media & Communication 9, 3 (2021), 422–440. https://doi.org/10.1177/2050157920970582
- Montagna et al. (2023) Sara Montagna, Stefano Ferretti, Lorenz Cuno Klopfenstein, Antonio Florio, and Martino Francesco Pengo. 2023. Data decentralisation of LLM-based chatbot systems in chronic disease self-management. In Proceedings of the 2023 ACM Conference on Information Technology for Social Good. 205–212.
- National Health Authority and Family Welfare (2020) Ministry of Health National Health Authority and Government of India Family Welfare. 2020. National Digital Health Mission: Strategy Overview. https://www.niti.gov.in/sites/default/files/2023-02/ndhm_strategy_overview.pdf
- Okolo et al. (2021) Chinasa T. Okolo, Srujana Kamath, Nicola Dell, and Aditya Vashistha. 2021. “It Cannot Do All of My Work”: Community Health Worker Perceptions of AI-Enabled Mobile Health Applications in Rural India. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 701, 20 pages. https://doi.org/10.1145/3411764.3445420
- OpenAI et al. (2023) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mo Bavarian, Jeff Belgum, Irwan Bello, Jake Berdine, Gabriel Bernadett-Shapiro, Christopher Berner, Lenny Bogdonoff, Oleg Boiko, Madelaine Boyd, Anna-Luisa Brakman, Greg Brockman, Tim Brooks, Miles Brundage, Kevin Button, Trevor Cai, Rosie Campbell, Andrew Cann, Brittany Carey, Chelsea Carlson, Rory Carmichael, Brooke Chan, Che Chang, Fotis Chantzis, Derek Chen, Sully Chen, Ruby Chen, Jason Chen, Mark Chen, Ben Chess, Chester Cho, Casey Chu, Hyung Won Chung, Dave Cummings, Jeremiah Currier, Yunxing Dai, Cory Decareaux, Thomas Degry, Noah Deutsch, Damien Deville, Arka Dhar, David Dohan, Steve Dowling, Sheila Dunning, Adrien Ecoffet, Atty Eleti, Tyna Eloundou, David Farhi, Liam Fedus, Niko Felix, Simón Posada Fishman, Juston Forte, Isabella Fulford, Leo Gao, Elie Georges, Christian Gibson, Vik Goel, Tarun Gogineni, Gabriel Goh, Rapha Gontijo-Lopes, Jonathan Gordon, Morgan Grafstein, Scott Gray, Ryan Greene, Joshua Gross, Shixiang Shane Gu, Yufei Guo, Chris Hallacy, Jesse Han, Jeff Harris, Yuchen He, Mike Heaton, Johannes Heidecke, Chris Hesse, Alan Hickey, Wade Hickey, Peter Hoeschele, Brandon Houghton, Kenny Hsu, Shengli Hu, Xin Hu, Joost Huizinga, Shantanu Jain, Shawn Jain, Joanne Jang, Angela Jiang, Roger Jiang, Haozhun Jin, Denny Jin, Shino Jomoto, Billie Jonn, Heewoo Jun, Tomer Kaftan, Łukasz Kaiser, Ali Kamali, Ingmar Kanitscheider, Nitish Shirish Keskar, Tabarak Khan, Logan Kilpatrick, Jong Wook Kim, Christina Kim, Yongjik Kim, Hendrik Kirchner, Jamie Kiros, Matt Knight, Daniel Kokotajlo, Łukasz Kondraciuk, Andrew Kondrich, Aris Konstantinidis, Kyle Kosic, Gretchen Krueger, Vishal Kuo, Michael Lampe, Ikai Lan, Teddy Lee, Jan Leike, Jade Leung, Daniel Levy, Chak Ming Li, Rachel Lim, Molly Lin, Stephanie Lin, Mateusz Litwin, Theresa Lopez, Ryan Lowe, Patricia Lue, Anna Makanju, Kim Malfacini, Sam Manning, Todor Markov, Yaniv Markovski, Bianca Martin, Katie Mayer, Andrew Mayne, Bob McGrew, Scott Mayer McKinney, Christine McLeavey, Paul McMillan, Jake McNeil, David Medina, Aalok Mehta, Jacob Menick, Luke Metz, Andrey Mishchenko, Pamela Mishkin, Vinnie Monaco, Evan Morikawa, Daniel Mossing, Tong Mu, Mira Murati, Oleg Murk, David Mély, Ashvin Nair, Reiichiro Nakano, Rajeev Nayak, Arvind Neelakantan, Richard Ngo, Hyeonwoo Noh, Long Ouyang, Cullen O’Keefe, Jakub Pachocki, Alex Paino, Joe Palermo, Ashley Pantuliano, Giambattista Parascandolo, Joel Parish, Emy Parparita, Alex Passos, Mikhail Pavlov, Andrew Peng, Adam Perelman, Filipe de Avila Belbute Peres, Michael Petrov, Henrique Ponde de Oliveira Pinto, Michael, Pokorny, Michelle Pokrass, Vitchyr Pong, Tolly Powell, Alethea Power, Boris Power, Elizabeth Proehl, Raul Puri, Alec Radford, Jack Rae, Aditya Ramesh, Cameron Raymond, Francis Real, Kendra Rimbach, Carl Ross, Bob Rotsted, Henri Roussez, Nick Ryder, Mario Saltarelli, Ted Sanders, Shibani Santurkar, Girish Sastry, Heather Schmidt, David Schnurr, John Schulman, Daniel Selsam, Kyla Sheppard, Toki Sherbakov, Jessica Shieh, Sarah Shoker, Pranav Shyam, Szymon Sidor, Eric Sigler, Maddie Simens, Jordan Sitkin, Katarina Slama, Ian Sohl, Benjamin Sokolowsky, Yang Song, Natalie Staudacher, Felipe Petroski Such, Natalie Summers, Ilya Sutskever, Jie Tang, Nikolas Tezak, Madeleine Thompson, Phil Tillet, Amin Tootoonchian, Elizabeth Tseng, Preston Tuggle, Nick Turley, Jerry Tworek, Juan Felipe Cerón Uribe, Andrea Vallone, Arun Vijayvergiya, Chelsea Voss, Carroll Wainwright, Justin Jay Wang, Alvin Wang, Ben Wang, Jonathan Ward, Jason Wei, CJ Weinmann, Akila Welihinda, Peter Welinder, Jiayi Weng, Lilian Weng, Matt Wiethoff, Dave Willner, Clemens Winter, Samuel Wolrich, Hannah Wong, Lauren Workman, Sherwin Wu, Jeff Wu, Michael Wu, Kai Xiao, Tao Xu, Sarah Yoo, Kevin Yu, Qiming Yuan, Wojciech Zaremba, Rowan Zellers, Chong Zhang, Marvin Zhang, Shengjia Zhao, Tianhao Zheng, Juntang Zhuang, William Zhuk, and Barret Zoph. 2023. GPT-4 Technical Report. arXiv:arXiv:2303.08774
- Radionova et al. (2023) Natalia Radionova, Eylem Ög, Anna-Jasmin Wetzel, Monika A Rieger, and Christine Preiser. 2023. Impacts of Symptom Checkers for Laypersons’ Self-diagnosis on Physicians in Primary Care: Scoping Review. Journal of Medical Internet Research 25 (2023), e39219.
- Roller et al. (2020) Stephen Roller, Emily Dinan, Naman Goyal, Da Ju, Mary Williamson, Yinhan Liu, Jing Xu, Myle Ott, Kurt Shuster, Eric M Smith, et al. 2020. Recipes for building an open-domain chatbot. arXiv preprint arXiv:2004.13637 (2020).
- Singh et al. (2017) Kamaljeet Singh, Arshi Misbah, Pranav Saluja, and Arun Kumar Singh. 2017. Review of manual small-incision cataract surgery. Indian journal of ophthalmology 65, 12 (2017), 1281.
- Skeels et al. (2010) Meredith M. Skeels, Kenton T. Unruh, Christopher Powell, and Wanda Pratt. 2010. Catalyzing Social Support for Breast Cancer Patients. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 173–182. https://doi.org/10.1145/1753326.1753353
- Stamenova et al. (2020) Vess Stamenova, Payal Agarwal, Leah Kelley, Jamie Fujioka, Megan Nguyen, Michelle Phung, Ivy Wong, Nike Onabajo, R Sacha Bhatia, and Onil Bhattacharyya. 2020. Uptake and patient and provider communication modality preferences of virtual visits in primary care: a retrospective cohort study in Canada. BMJ open 10, 7 (2020), e037064.
- Statista (2022) Statista. 2022. Most used social media and messaging apps in India 2022. https://www.statista.com/statistics/1235799/india-most-used-social-media-and-messaging-apps/ Accessed: 2024-02-01.
- Tang and Lansky (2005) Paul C Tang and David Lansky. 2005. The missing link: bridging the patient-provider health information gap. Health Aff. (Millwood) 24, 5 (Sept. 2005), 1290–1295.
- Tang and Newcomb (1998) P C Tang and C Newcomb. 1998. Informing patients: a guide for providing patient health information. J. Am. Med. Inform. Assoc. 5, 6 (Nov. 1998), 563–570.
- ter Hofstede et al. (1996) A. H. M. ter Hofstede, H. A. Proper, and Th. P. van der Weide. 1996. Query Formulation as an Information Retrieval Problem. Comput. J. 39, 4 (01 1996), 255–274. https://doi.org/10.1093/comjnl/39.4.255 arXiv:https://academic.oup.com/comjnl/article-pdf/39/4/255/1916704/390255.pdf
- Tongue et al. (2005) John R. Tongue, Howard R. Epps, and Laura L. Forese. 2005. Communication skills for patient-centered care : Research-based, easily learned techniques for medical interviews that benefit orthopaedic surgeons and their patients. Journal of Bone and Joint Surgery, American Volume 87 (2005), 652–658. https://api.semanticscholar.org/CorpusID:71035586
- Tseng et al. (2023) Yuan-Chi Tseng, Weerachaya Jarupreechachan, and Tuan-He Lee. 2023. Understanding the Benefits and Design of Chatbots to Meet the Healthcare Needs of Migrant Workers. 7, CSCW2, Article 315 (oct 2023), 34 pages. https://doi.org/10.1145/3610106
- Wang et al. (2020) Ding Wang, Santosh D. Kale, and Jacki O’Neill. 2020. Please Call the Specialism: Using WeChat to Support Patient Care in China. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3313831.3376274
- Wang et al. (2021b) Dakuo Wang, Liuping Wang, Zhan Zhang, Ding Wang, Haiyi Zhu, Yvonne Gao, Xiangmin Fan, and Feng Tian. 2021b. “Brilliant AI Doctor” in Rural Clinics: Challenges in AI-Powered Clinical Decision Support System Deployment. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Yokohama¡/city¿, ¡country¿Japan¡/country¿, ¡/conf-loc¿) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 697, 18 pages. https://doi.org/10.1145/3411764.3445432
- Wang et al. (2021a) Liuping Wang, Dakuo Wang, Feng Tian, Zhenhui Peng, Xiangmin Fan, Zhan Zhang, Mo Yu, Xiaojuan Ma, and Hongan Wang. 2021a. CASS: Towards Building a Social-Support Chatbot for Online Health Community. Proc. ACM Hum.-Comput. Interact. 5, CSCW1, Article 9 (apr 2021), 31 pages. https://doi.org/10.1145/3449083
- Weisz et al. (2019) Justin D. Weisz, Mohit Jain, Narendra Nath Joshi, James Johnson, and Ingrid Lange. 2019. BigBlueBot: teaching strategies for successful human-agent interactions. In Proceedings of the 24th International Conference on Intelligent User Interfaces (Marina del Ray, California) (IUI ’19). Association for Computing Machinery, New York, NY, USA, 448–459. https://doi.org/10.1145/3301275.3302290
- Wilcox et al. (2010) Lauren Wilcox, Dan Morris, Desney Tan, and Justin Gatewood. 2010. Designing Patient-Centric Information Displays for Hospitals. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 2123–2132. https://doi.org/10.1145/1753326.1753650
- Williams and Moser (2019) Michael Williams and Tami Moser. 2019. The art of coding and thematic exploration in qualitative research. International Management Review 15, 1 (2019), 45–55.
- Xiao et al. (2023) Ziang Xiao, Q. Vera Liao, Michelle Zhou, Tyrone Grandison, and Yunyao Li. 2023. Powering an AI Chatbot with Expert Sourcing to Support Credible Health Information Access. In Proceedings of the 28th International Conference on Intelligent User Interfaces (Sydney, NSW, Australia) (IUI ’23). Association for Computing Machinery, New York, NY, USA, 2–18. https://doi.org/10.1145/3581641.3584031
- Yadav et al. (2019) Deepika Yadav, Prerna Malik, Kirti Dabas, and Pushpendra Singh. 2019. Feedpal: Understanding Opportunities for Chatbots in Breastfeeding Education of Women in India. Proc. ACM Hum.-Comput. Interact. 3, CSCW, Article 170 (nov 2019), 30 pages. https://doi.org/10.1145/3359272
- Yang et al. (2023) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Shao Zhang, Ethan Rogers, Stephen Intille, Nawar Shara, Dakuo Wang, et al. 2023. Talk2Care: Facilitating Asynchronous Patient-Provider Communication with Large-Language-Model. arXiv preprint arXiv:2309.09357 (2023). arXiv:2309.09357 [cs.CL]
- YM et al. (2023) Pradyumna YM, Vinod Ganesan, Dinesh Kumar Arumugam, Meghna Gupta, Nischith Shadagopan, Tanay Dixit, Sameer Segal, Pratyush Kumar, Mohit Jain, and Sriram Rajamani. 2023. PwR: Exploring the Role of Representations in Conversational Programming. arXiv:2309.09495 [cs.HC]
Appendix A Appendix
A.1. Cataract Bot Implementation Details
A.1.1. Cataract Bot LLM Prompts
The CataractBot system leverages LLM (GPT-4 in our case) for these four tasks:
Response Generation: For every medical and logistical question asked by the patient/attendant, the system performs a vector search on the Knowledge Base (KB) to extract the three most relevant data chunks related to the query. The LLM is then prompted (Table 3) to extract an answer for the query from these data chunks. Note: Question Classification is part of the same LLM call to improve efficiency.
Related Questions Generation: For every medical and logistical question, the system prompts LLM to generate three related questions based on the preceding query. Note: The 72 character limit is due to the WhatsApp’s message limit in interactive suggestions.
Final Response Generation: If the expert marks the initial LLM-generated response as incorrect or incomplete, the system prompts LLM to generate the final response by merging both the initial response and the expert’s edit suggestion.
Shorten Response: If the generated response exceeds WhatsApp’s message limit of 700 characters, the system prompts LLM to summarize it within the specified character limit.
| System Prompt | Query Prompt | ||||||||||||||||||||||||||||||||||||

### v2

@@
 Abstract.
-The healthcare landscape is evolving, with patients seeking more reliable information about their health conditions, treatment options, and potential risks.
+The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options.
 Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information.
-Patients primarily trust doctors and hospital staff, highlighting the need for expert-endorsed health information.
-However, the pressure on experts has led to reduced communication time, impacting information sharing.
-To address this gap, we propose CataractBot, an experts-in-the-loop chatbot powered by large language models (LLMs).
-Developed in collaboration with a tertiary eye hospital in India, CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously.
-CataractBot features multimodal support and multilingual capabilities.
-In an in-the-wild deployment study with 49 participants, CataractBot proved valuable, providing anytime accessibility, saving time, and accommodating diverse literacy levels.
-Trust was established through expert verification.
+Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information.
+However, increased patient loads on experts has led to reduced communication time, impacting information sharing.
+To address this gap, we develop CataractBot, an experts-in-the-loop chatbot powered by LLMs, in collaboration with an eye hospital in India.
+CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously.
+It has multimodal and multilingual capabilities.
+In an in-the-wild deployment study with 55 participants, CataractBot proved valuable, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors.
+Users reported that their trust in the system was established through expert verification.
 Broadly, our results could inform future work on designing expert-mediated LLM bots.
 1. Introduction
-The evolving landscape of healthcare is witnessing a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007).
-It signifies a departure from the traditional passive patient role toward a more informed, empowered, and engaged approach to healthcare, resulting in an increased demand for information.
-Patients and their caregivers seek accessible, comprehensive, and reliable information about their health conditions, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019).
+The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information.
+Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019).
 To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010).
-In particular, for major treatments like surgery or radiotherapy, this demand for information becomes more crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases.
+In particular, for major treatments like surgery or radiotherapy, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases.
 Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002).
-Conversely, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
+Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
 Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007).
-This poses a challenge in discerning trustworthy sources from misleading ones, leading to potential misconceptions.
-Additionally, patients frequently encounter difficulties finding information online at an appropriate level–ranging from oversimplified to overly technical–and harbor skepticism toward its reliability (Xiao et al., 2023).
-As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although locating them can be challenging (Civan-Hartzler et al., 2010).
+This poses a challenge in discerning trustworthy sources from misleading ones.
+Patients also frequently encounter difficulties finding information online at an appropriate level, ranging from oversimplified to overly technical (Xiao et al., 2023).
+As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although identifying them can be challenging (Civan-Hartzler et al., 2010).
 Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a).
 Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998).
-Study participants noted that being better informed increased their understanding of treatment plans, motivation to adhere to them, and satisfaction with hospital visits.
 However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing.
-In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a).
-Even providing all potentially relevant information can pose challenges as it can overwhelm patients with cognitive overload (Christalle et al., 2019).
+In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a).
 Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide.
+However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019).
 Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004).
@@
 Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information.
-In collaboration with the medical team at a tertiary eye hospital in Bangalore, India, we exemplified our approach by developing CataractBot.
+In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot.
 This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants.
-Recognizing the limitations of generic LLMs in the medical domain, we utilized RAG (Retrieval-Augmented Generation) using a knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses.
-CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and updates to the knowledge base based on expert-provided edits to minimize future expert intervention.
-In particular, for high-stakes medical settings, access to reliable and trustworthy information is crucial, hence, each response that patients (or their attendants) receive undergoes scrutiny and potential editing by experts (doctors or patient coordinators).
-CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantaneously and subsequently providing verification at the experts’ convenience.
-With that, the paper aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and their attendants) undergoing surgical treatment?
-(RQ2) How does the inclusion of an expert-in-the-loop influence the perceived trustworthiness of the information provided by the chatbot?
-(RQ3) How are such chatbots integrated into the workflow of doctors, patient coordinators, patients, and their attendants?
-To answer them, we conducted an in-the-wild deployment study involving 49 information seekers (19 patients, 30 attendants) using CataractBot for at least two weeks—a week before and after the surgery.
-Additionally, four doctors and two patient coordinators served as experts in this study.
-Our findings revealed patients and attendants highly appreciated CataractBot for its anytime accessibility, alleviating hesitations associated with power differences in asking questions, saving time by minimizing hospital visits/calls, and accommodating individuals with low literacy and tech proficiency through multilingual and multimodal support.
-Trust in the system was established through rigorous verification performed by doctors and coordinators.
-Experts praised the bot as a facilitator, introducing a layer between them and patients, and providing flexibility through asynchronous communication.
-Moreover, experts noted improved responses over time, likely due to continuous knowledge base updates.
-Our main contributions are: (1) The conception and development of a specialized chatbot111Our code is available at https://github.com/microsoft/BYOeB., designed in close collaboration with doctors and hospital staff, to assist patients undergoing Cataract surgery with their information needs.
-(2) The deployment of this chatbot CataractBot in a real-world setting, among 31 Cataract surgery patients and their attendants at a tertiary eye hospital in India.
-Four doctors and two patient coordinators served as experts, actively verifying and refining responses generated by the chatbot.
-(3) Drawing from a comprehensive mixed-methods analysis of gathered data, our study not only presents pivotal insights specific to CataractBot but also offers broader implications of LLM-powered expert-in-the-loop chatbots.
+Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (RAG) over a Knowledge Base (KB) curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses.
+CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a KB update framework based on expert-provided edits to minimize future expert intervention.
+CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
+Our work aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and attendants) undergoing surgery?
+(RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage?
+(RQ3) How can CataractBot be integrated into the current doctor-patient workflow?
+To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use the terms ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators.
+(19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators).
+Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support.
+Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators.
+Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication.
+Moreover, experts noted improved responses over time, likely due to continuous KB updates.
+Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://removed_for_anonymity., designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs.
+(2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants at a tertiary eye hospital in India.
+Four doctors and two patient coordinators served as experts, actively verifying and correcting responses generated by the bot.
+(3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
 2. Related Work
-Our work is primarily informed by two areas of relevant research: solutions to address patients’ information needs and conversational agents used in healthcare settings.
-2.1. Addressing Patients’ Information Needs
-Prior work has focused on understanding the diverse informational requirements of patients, particularly in the context of healthcare decision-making (Tseng et al., 2023; Xiao et al., 2023; Beisecker and Beisecker, 1990).
-Patients and their attendants exhibit a pattern of seeking information throughout the various stages of a patient’s journey, such as to prepare for doctor consultations (Radionova et al., 2023; Leong et al., 2018), to verify diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009), to assess the need for clinical interventions (Yang et al., 2023; Tseng et al., 2023), and to aid in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018).
-Studies emphasize that satisfying these diverse information needs enhances patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998).
-To fulfill these diverse informational needs, patients, while trusting physicians for their clinical expertise, often turn to the Internet for the ease of access to information (Xiao et al., 2023).
-However, the challenge lies in identifying relevant and appropriate information, with instances of individuals consuming inaccurate information (Clarke et al., 2016b).
-Information seekers also resort to consulting experienced patients within their social network who can share valuable knowledge and insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010).
-Recent work in the field of Computer-Human Interaction has explored novel avenues to address the medical information gaps.
+Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, including those powered by LLMs, in healthcare settings.
+2.1. Patients’ Information Needs
+Patients and their attendants seek information throughout the various stages of their medical journey, such as preparing for doctor consultations (Radionova et al., 2023; Leong et al., 2018), verifying diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009; Christalle et al., 2019), assessing the need for clinical interventions (Yang et al., 2024; Tseng et al., 2023), and aiding in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018).
+Studies emphasize that satisfying these diverse information needs improves patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998).
+Patients primarily trust doctors with their medical queries (Ding et al., 2020).
+However, traditional in-person consultations pose difficulties, including high costs (Bodenheimer and Smith, 2013), long wait times, and a strong hierarchy in the patient-physician relationship (Fochsen et al., 2006).
+Patients often miss opportunities to ask additional (clarification) questions during these brief encounters, as they face information overload (Ferré et al., 2020; Kinnersley et al., 2007).
+Doctors also struggle with time due to staff shortages straining healthcare ecosystems (Bodenheimer and Smith, 2013; Akerkar and Bichile, 2004).
+Thus, patients and caregivers are increasingly turning to the Internet for easier access to medical information (Xiao et al., 2023; Chandwani and Kulkarni, 2016).
+However, the challenge lies in identifying relevant and appropriate information.
+Individuals struggle to understand complex health information or may even consume inaccurate information, making the process frustrating (Zhang, 2010; Arora et al., 2008; Clarke et al., 2016b).
+Information seekers may also consult experienced patients within their social networks who can share valuable insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010).
+In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014).
+Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
+Recent work in the field of Computer-Human Interaction has explored novel avenues to address these medical information gaps.
 For instance, Leong et al.
-(2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, while Wilcox et al.
+(2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, Wilcox et al.
 (2010) and Bickmore et al.
-(2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details based on their records.
-Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005).
+(2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details, and Pfeifer Vardoulakis et al.
+(2012) propose using mobile phones to present dynamic, interactive reports on patients’ progress and care plans throughout their emergency department stay.
+Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005; Chung, 2017).
 Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
-In summary, patients require reliable, relevant, and timely information.
-However, challenges arise in accessing such information from busy or hesitant doctors and hospital staff.
-This dichotomy results in patients either receiving generic, unverified information from the internet and social circles or obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts.
-Our solution aims to strike a balance by providing a free, generic, high-quality answer instantaneously using Large Language Models (LLMs), which is later verified (and corrected if needed) by a medical professional with minimal increase in their workload.
+In summary, patients need reliable, relevant, and timely information.
+However, accessing this from busy doctors and hospital staff is challenging.
+This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles or if lucky, obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts.
+Our solution aims to strike a balance by providing a generic, high-quality answer instantly using LLMs, which is later verified by a medical professional with minimal increase to their workload.
 2.2. Chatbots in Healthcare
-Recently, researchers have been exploring chatbots as a means to provide easy and direct access to information, due to their ubiquity, minimal learning curve, and ability to offer a more personalized experience through natural language conversation (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020).
-Chatbots have been found to enhance the user experience by enabling users to engage in longer conversations and build complex queries message by message while retaining conversation context (Xiao et al., 2023).
-Stamenova et al.
-(2020) reveal that patients prefer messaging over video consultations, emphasizing that chatbots can help promote continuity in the treatment.
-Chatbots have been applied in various healthcare settings by the research community, including mental health support (Kumar et al., 2023; Jo et al., 2023), appointment scheduling, seeking information for chronic disease management (Montagna et al., 2023), assistance for new mothers (Yadav et al., 2019), and public health interventions to support socially isolated individuals (Jo et al., 2023).
-Commercial offerings, such as Babylon Health111Babylon Health: https://www.emed.com/uk, Ada222Ada: https://ada.com/, and Florence333Florence: https://florence.chat/, allow individuals to inquire about their health issues.
-All of these bots are fully automated, with some even using LLMs, and the involvement of experts is limited to the data curation step (Xiao et al., 2023).
-Hence, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
-Beyond standalone chatbots, prior work has explored chat as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners.
-Synchronous messaging requires both the patient and the doctor to be simultaneously active, resembling an in-person consultation, allowing instant feedback and coherent dialogue.
-On the other hand, asynchronous messaging platforms like WhatsApp, WeChat (Wang et al., 2020), or email enable communication between the patient and doctor based on their convenience, providing flexibility.
-Previous research indicates that both doctors and patients prefer asynchronous texts over instant calls (audio or video) (Stamenova et al., 2020; Johansson et al., 2020a).
-In our work, we combine synchronous and asynchronous modes of communication, leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode.
+Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023).
+They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023).
+Researchers have applied chatbots in various healthcare settings (Wutz et al., 2023), including mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023), appointment scheduling (Ittarat et al., 2023), seeking information for chronic disease management (Montagna et al., 2023), and assistance for new mothers (Yadav et al., 2019).
+Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk.
+Ada: https://ada.com/.
+Florence: https://florence.chat/., allow individuals to inquire about their health issues.
+All of these bots are fully automated, with some even using LLMs.
+However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023).
+Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
+LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018).
+Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020), and few-shot learning.
+Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022).
+Seitz et al.
+(2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport.
+Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
+Given our specific focus on cataract surgery, we also examined literature at the intersection of ophthalmology and LLMs.
+Bernstein et al.
+(2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists.
+Based on similar findings, Ittarat et al.
+(2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions.
+Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients.
+Our work contributes to this body of literature.
+Beyond chatbots, prior work has explored “chat” as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners.
+Synchronous messaging requires both the patient and the doctor to be simultaneously active, allowing instant feedback and coherent dialogue.
+On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties, providing flexibility.
+Previous research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023).
+In our work, we combine both approaches—leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode.
 This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.
+Finally, the use of AI in real-world healthcare deployments is limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021).
+To avoid such shortcomings, as recommended by Thieme et al.
+(2023b, a), we designed and studied our system within a tertiary eye hospital in India and validated its utility to understand its integration into clinical workflows.
+3. Formative Study
+The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital.
+To identify key chatbot requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023.
+From these interviews, we aimed to understand their current workflow, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps.
+Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience.
+They participated in the study voluntarily without compensation.
+The interviews were conducted in English and lasted around 45 minutes each.
+All sessions were audio-recorded and transcribed by the interviewer.
+We conducted a thematic analysis to identify these key design requirements for the chatbot.
+Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context.
+While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances.
+For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness.
+Similarly, cultural considerations significantly influence pre- and post-surgery questions.
+As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance.
+To ensure accurate and contextual responses, our chatbot relies on a Knowledge Base (KB) curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
+Trustworthiness: Access to trustworthy information is crucial during medical treatment.
+As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff.
+Even when patients learn something from the internet or friends, they often confirm that information with their doctor.
+E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens.
+Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data.
+This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
+Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses.
+Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors.
+Due to busy schedules of doctors and coordinators, immediate responses are not always feasible.
+This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs.
+To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom KB curated by hospital staff.
+Subsequently, these responses gets reviewed and refined by doctors/coordinators.
+Moreover, doctors mentioned their unavailability for several hours while performing surgeries.
+To ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
+Usability: After cataract surgery, patients are expected to reference their discharge summaries for any queries, but this can be challenging due to their temporarily reduced vision.
+Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role immediate family members play in monitoring the patient’s healthcare.
+Coupled with doctors’ input (“Hardly 10% of questions are asked by patients; attendants are the one who ask.”), this underscored the necessity for the bot to be usable by the four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators.
+To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal similar to (Jain et al., 2018; Findlater et al., 2009).
+We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022), despite the richer feature set of Telegram or the development flexibility of a custom chatbot.
+This choice aims to minimize the learning curve and ensure quick onboarding without requiring additional app installations.
+Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s KB, minimizing similar edits in future.
 4. CataractBot System Design
-3.1. Design Requirements
-Drawing from our discussions with doctors and patient coordinators, literature review, and understanding of the Indian healthcare ecosystem, we have identified several key requirements for the chatbot.
-Accuracy: The chatbot’s capability to answer both medical and logistical questions with accuracy is a fundamental requirement.
-Providing precise, contextually relevant, and verified responses is paramount given the medical context.
-While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it might not be tailored to India-specific or hospital-specific nuances.
-For instance, while the phacoemulsification method is prevalent globally for cataract surgeries, India often favors manual small incision-based procedures due to their cost-effectiveness (Singh et al., 2017).
-Similarly, cultural considerations significantly impact advice on pre- and post-surgery choices, e.g., questions related to diet necessitates inclusion and exclusion of Indian food options for accurate guidance.
-To ensure accurate and contextual response, our chatbot generates responses from a knowledge base curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
-Trustworthiness: In medical surgical treatment, access to reliable trustworthy information is crucial.
-Previous research (Xiao et al., 2023) highlights how patients often encounter misinformation from Internet-based (medical) sources, leading to skepticism towards them.
-Consequently, they tend to rely on the experiences of friends and family members who have undergone similar treatments, although such resources can sometimes be challenging to locate (Civan-Hartzler et al., 2010).
-Typically, patients place their utmost trust in the operating doctor and the hospital staff responsible for their care during the procedure (Larasati et al., 2023).
-Hence, each response the patient (or their attendant) receives is vetted and, if necessary, edited by doctors and patient coordinators to ensure trustworthiness.
-Timeliness: Prompt responses are important as delays often lead patients to seek information elsewhere.
-Patient coordinators revealed that patients frequently reach out to them (over a phone call) for relaying medical queries to their consulting doctors.
-However, due to busy schedules of doctors and coordinators, immediate responses are not always feasible.
-That often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation.
-To counter this, we decided the chatbot should provide instantaneous responses to every patient query using a custom knowledge base curated by hospital staff.
-Subsequently, these responses were reviewed and refined by doctors/coordinators later.
-Usability: The developed system must be usable by patients, their attendants, doctors and patient coordinators.
-Thus, we chose WhatsApp as our platform to develop CataractBot, as it is the most widely used messaging platform in India (Statista, 2022).
-Despite Telegram’s richer feature set or a custom chatbot offering more development flexibility, we opted for WhatsApp to minimize the learning curve and ensure seamless onboarding without requiring additional app installations.
-Furthermore, due to the frequent use of WhatsApp among Indians (Statista, 2022), we anticipated higher adoption and usage of a WhatsApp-based chatbot.
-To cater to the diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, we made the chatbot multilingual—supporting five languages (English, Hindi, Kannada, Tamil, and Telugu)—and multimodal (offering both text and speech inputs and outputs), similar to (Jain et al., 2018; Findlater et al., 2009).
+Here, we describe the CataractBot system (Figure 3), including its key components–input language and modality, response generation and verification, and escalation and reminders–and provide a detailed account of its implementation.
 4.1. Components
-From our discussions with doctors and hospital staff, we have identified four key stakeholders in the cataract surgery ecosystem: the patient, their attendants, the doctor, and the patient coordinator.
-Once onboarded on CataractBot, a patient (and their attendant) scheduled for cataract surgery within the next 7 days can ask any question related to the surgical treatment and receive a response verified by the hospital staff.
-In addition, we have a few more roles: an escalation doctor, an escalation patient coordinator, and a KB (Knowledge Base) update expert.
-Below, we describe the key components of our CataractBot system.
 4.1.1. Input Language and Modality
-The Sankara Eye Hospital situated in Bangalore (in the state of Karnataka), known as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states.
+Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states.
+Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011).
 To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka).
-Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language via WhatsApp.
-Among these messages is an option allowing users to modify their preferred language using a drop-down menu at any time in the future.
-Note: To converse in a specific language, users must ensure their preferred language is selected before interacting.
-For example, if a user has chosen English as their preferred language, they cannot begin conversing with CataractBot in Kannada.
-However, they can navigate to the welcome messages, switch their preferred language to Kannada, and then engage with the bot in Kannada.
-To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, both text and speech inputs are enabled (Jain et al., 2018; Findlater et al., 2009).
-Users can ask questions by holding the mic button on WhatsApp, speaking in their preferred language, and sending a voice message.
-For every voice message, CataractBot not only provides a text response but also sends an audio message (Figure 2D).
-Given that the doctor and patient coordinator may not be proficient in all five languages, their interactions with the bot for verifying and editing messages are exclusively conducted in English.
-Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot addresses this by providing a set of three frequently asked questions in a drop-down menu along with the welcome message.
-Additionally, after every medical and logistical question from the patient/attendant, the bot’s response includes a “What to do next?” prompt, offering three related questions in a drop-down menu based on the preceding query (Figure 2F).
+As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English.
+Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language.
+Among these messages is an option allowing users to modify their preferred language in future.
+To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs.
+For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
+Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message.
+Additionally, every bot’s response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F).
 This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time.
-Users are free to select one of the suggested questions or pose a different query to continue the conversation.
 4.1.2. Response Generation, Verification and Icons
-Upon receiving a message from a patient/attendant, CataractBot classifies it as a medical question, logistical question, small talk, or other queries.
-For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response.
-In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response.
+Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk, or ‘other’, and provides a response in real-time.
+For medical and logistical questions, the bot strictly employs the KB curated by the hospital’s medical team to generate an appropriate response.
+The response includes a question mark icon
+as a ‘reaction’ indicating the unverified status (Figure 2E).
+In instances where the custom KB lacks an answer, the bot responds with a template “I don’t know” response.
 For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
-For miscellaneous questions unrelated to medical or logistical matters, e.g., about sports or weather, the bot responds with a template “I don’t know” message.
-The patient/attendant receives these responses in real-time from CataractBot.
-Specifically, for medical and logistical questions, the response includes a question mark icon
-as a reaction indicating the unverified status (Figure 2E).
-For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, patient demographics, and citations of the documents used to generate the response.
-The doctor is then prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
+For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
+As Rajashekar et al. (2024) found that citations improve trust in LLM-generated responses, this message also includes citations of the documents used to generate the response.
+The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
 Tapping ‘Yes’ replaces the question mark icon
-in the user’s received answer with a green tick icon
-, confirming verification (Figure 1F). Additionally, the bot notifies the user that the answer has been verified, tagging that particular response (Figure 1G).
+in the information seeker’s received answer with a green tick icon
+, confirming verification (Figure 1F). Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G).
 Tapping ‘No’ replaces the question mark
 with a red cross icon
-, indicating an incorrect answer (Figure 1C), and asks the user to await a corrected response from the doctor (Figure 1D).
-The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they need to offer a correction in free form.
+, indicating an incorrect answer (Figure 1C), and asks the information seeker to await a corrected response from the doctor (Figure 1D).
+The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they can offer a correction in free form text.
 The CataractBot system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
 (Figure 1E).
+This new response is not sent to the doctor, to minimize their workload by limiting verification to a single round.
 If a question is misclassified, i.e., a logistical question is sent to the doctor, they have the option to ‘Send to Patient Coordinator’.
-Similar to the doctor’s workflow, for logistical questions, the patient coordinators verify and provide corrections to the bot’s response.
-In deciding how to deliver chatbot responses to patients, we deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction.
-Through our discussions with hospital staff, it became clear that providing real-time responses was crucial.
+Similar to the doctor’s workflow, the patient coordinators verify and provide corrections to the bot’s response for logistical questions.
+We deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction.
+Through our
+formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial.
 Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
 4.1.3. Escalation Mechanism
-The CataractBot system employs an intricate escalation mechanism to ensure swift verification of every query posed by the patient/attendant.
+The CataractBot system employs an intricate escalation and reminder mechanism to ensure swift verification.
 In the doctor’s WhatsApp interface, unanswered questions are marked with a question mark icon
-(Figure 2E), enabling them to easily identify pending queries.
-If an answer is not verified by the operating doctor within three hours, it automatically escalates to the designated escalation doctor.
-This duration was set based on input from the hospital staff.
+(Figure 2E) and answered questions with a green tick
+, enabling them to easily identify pending queries.
+If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor.
 The operating doctor receives a notification about the escalation.
 Both the operating and escalation doctors can then verify/correct the response.
-Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick
-(Figure 1F) for both the doctors, indicating task completion.
+Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick (Figure 1F) for both the doctors, indicating task completion.
 However, if either doctor selects ‘No’, the green tick appears only after that doctor provides a correction.
-In instances where neither the operating nor escalation doctor verifies a question within six hours, both receive a notification indicating the pending status.
-Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors, tagged with appropriate messages.
-This additional reminder was added based on the doctor’s feedback.
-This workflow is mirrored for the patient coordinator and escalation coordinator.
+If neither the operating nor escalation doctor verifies a question within six hours, both receive a reminder notification indicating the pending status.
+Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors.
+This additional reminder was added based on the doctors’ feedback.
+This workflow is mirrored for the patient coordinator and escalation patient coordinator.
 4.1.4. KB Update Process
-While verification remains vital in the medical context, we aim to reduce the effort required from already burdened medical professionals in verifying and editing CataractBot responses.
-To achieve this, we devised a mechanism to minimize repetitive edits performed by experts.
-This process involves using the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future.
-However, we understand that certain responses, such as those specific to individual patients (e.g., “when to reach the hospital for my surgery?”), should not be added to the KB.
-Thus, we appointed a senior cataract surgery doctor as the ‘KB update expert’.
-Each night at 8 pm, this expert automatically receives an online spreadsheet via email containing columns such as ‘Question asked’, ‘Bot’s response’, ‘Expert’s correction’, ‘Bot’s Corrected Response’, and ‘Should Update KB?’.
-Their task involves reviewing each row and determining whether the information should be added to the CataractBot KB.
-They respond with a ‘Yes’ or ‘No’ in the ‘Should Update KB?’ column.
-If marked ‘Yes’, the KB update expert can modify the ‘Bot’s Corrected Response’ column as needed.
-At 3 am daily, we extract ‘Question asked’ and ‘Bot’s Corrected Response’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document.
-Additionally, we prompt the large language model (LLM) to prioritize this ‘expert-FAQ’ document within the KB.
-This ensures that the bot learns from these updates, gradually enhancing its knowledge over time.
-Consequently, when the bot encounters a similar question in the future, it can provide an accurate response without requiring further corrections, thus effectively reducing the workload of doctors and coordinators.
+To minimize experts’ labor, we use the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future.
+However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the KB as they are not generalizable.
+We appointed a senior cataract surgeon as the ‘KB update expert’, who received a spreadsheet via email at 8pm daily.
+Their task involves reviewing each row and determining whether the information should be added to the KB (by responding with a ‘Yes’/‘No’ in the ‘Should Update KB?’ column), and if so, modifying as needed the ‘Final Answer for KB’ column, containing the bot’s updated answer based on the expert’s correction.
+At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for KB’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document.
+Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the KB.
+This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
 4.2. Implementation Details
-The CataractBot system mainly relies on these five services (Figure 3): (a) LLM (Large Language Model): for response generation, (b) Vector Database: for storing and retrieving the custom knowledge base, (c) Language Technologies: for translation, and for converting speech-to-text and text-to-speech, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
+The CataractBot system relies on these five components (Figure 3): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom KB, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
 We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023).
-Considering the medical context, instead of relying on GPT-4’s default knowledge, CataractBot utilized a curated knowledge base by the hospital staff.
-This custom KB comprises of various Cataract surgery and Sankara Eye Hospital-related documents, including Consent Form, SOP (Standard Operating Procedure), FAQs (Frequently Asked Questions), Pre- and Post-Operative Guidelines, Hospital Information, and Patient Discharge Do’s and Dont’s.
+The custom KB comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information.
 These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings.
-Upon receiving a question asked by patient/attendant, GPT-4 classifies the query type (medical, logistical, small talk, or other).
-For a medical/logistical question, CataractBot employs a RAG (Retrieval Augmented Generation) (Lewis et al., 2020) approach.
+Upon receiving a question, GPT-4 classifies the query type.
+For a medical/logistical question, CataractBot employs a RAG (Lewis et al., 2020) approach.
 This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query.
-GPT-4 is then prompted (full prompt available in Appendix A) to extract an answer for the query from these data chunks.
+GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks.
 If the answer is not present, the bot responds with an “I don’t know” message.
-Moreover, GPT-4 is prompted to give priority to the ‘expert-FAQ’ document–vetted by medical professionals–when it is among the top data chunks.
-The prompt also instructs that the response should adhere to WhatsApp’s message limit of 700 characters.
-This prompt underwent several iterations.
-We engaged a doctor and a patient coordinator at the hospital to record questions posed by patients/attendants scheduled for cataract surgery over a week.
-A total of 153 questions were compiled.
+The prompt underwent several iterations.
+To validate its effectiveness, a doctor and a patient coordinatorrecorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week.
 We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator.
-This not only helped refine the LLM prompt, but also led the medical staff to improve the content of existing documents and contribute additional documents to enrich the KB.
-This process was iterated three times, to ensure comprehensive coverage and validate the effectiveness of our GPT-4 prompt with the RAG approach in extracting pertinent responses.
-Despite their effectiveness, we recognized the KB’s susceptibility to outdated or incomplete information.
-Hence, expert edits to bot’s answers, confirmed by the KB update expert, were consistently incorporated to regularly enhance and update the KB.
-Additionally, we utilized GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response.
+This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the KB.
+This process was iterated thrice.
+Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response.
 This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input.
-Our experience showed that GPT-4 flawlessly executed this task.
-Discussing language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023).
-To bridge the language gap, the CataractBot system adhered to the standard approach (ait, [n.
-d.]) of translating local Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator.
-It then translated GPT-4’s English responses back into the respective local Indic languages for output.
-Moreover, to facilitate speech as input and audio as output, CataractBot used Azure speech-to-text and text-to-speech models.
-(Note: GPT-4 lacks support for speech input.)
-During testing, we encountered instances where GPT-4 was busy, resulting in response failures due to timeouts.
-To guarantee responses for every question, we implemented a queuing system.
-Each message from patients/attendants entered a queue and was removed only after receiving a response from GPT-4.
-Furthermore, to ensure expert verification, we employed multiple CronJobs for notifications and escalations.
-Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) gets logged in a cloud storage for further analysis.
+Our experience showed that GPT-4 executed this task well.
+Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023).
+To bridge the language gap, our system adhered to the standard approach (ait, [n.
+d.]) of translating Indic languages into English for input using Azure AI Translator.
+It then translated GPT-4’s English responses back into the respective Indic languages for output.
+To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models.
+(Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) get logged in a cloud storage for further analysis.
 5. Study Design
-To evaluate the effectiveness of an LLM-powered expert-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 in collaboration with a local eye hospital, Sankara Eye Hospital.
-Sankara Eye Hospital is one of the leading tertiary eye care and teaching institutions in Bengaluru, India.
-It has 18 eye doctors (15 of whom perform cataract surgeries regularly) and over 10 optometrists.
-The hospital typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily.
+To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital.
+This is one of the leading tertiary eye care and teaching institutions in Bangalore, India.
+It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily.
 This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital.
+None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
 5.1. Procedure for Patients and Attendants
-As per the hospital’s protocol, once a patient elects to proceed with cataract surgery following the doctor’s recommendation, the patient and their attendant are directed by the doctor to meet a patient coordinator.
-The coordinator helps with general counseling, guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements.
-Toward the end of this interaction, the patient coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages (English, Hindi, Kannada, Tamil, or Telugu), (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as they would likely have minimal informational needs), and (e) being a patient of one of the three participating operating doctors.
-If the patient meets these criteria, the patient and their attendant are directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
-The researcher introduced them to CataractBot and outlined the study’s protocol which involved using CataractBot for around two weeks, engaging in two interviews, and completing a chatbot usability form.
-Upon confirming their willingness to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form.
+As per the hospital’s protocol, once a patient opts for cataract surgery based on their doctor’s recommendation, the patient and their attendant needs to meet a patient coordinator.
+The coordinator helps with general guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements.
+Toward the end of this interaction, the coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon.
+If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
+The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks, engaging in an interview pre and post-surgery, and completing a chatbot usability form.
+If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form.
 This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education).
-Upon form submission, participants received a set of ‘welcome messages’ from CataractBot.
-The researcher instructed them to ask a trial question to the bot, either by choosing from suggested questions (Figure 2F) or by typing/speaking a new question in their preferred language.
-Upon receiving a response, participants were briefed on the three icons (Figure 1E,F): question mark, green tick, and red cross.
-Throughout this process, the researcher encouraged them to clarify any doubts regarding the bot’s usage.
-Following this onboarding process, patients and their attendants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to ask questions to CataractBot.
-Participants were instructed to use the bot from onboarding to one week post-surgery, after which their access was automatically revoked.
-Given that cataract surgeries were typically scheduled within a week of meeting with the patient coordinator, each patient/attendant was expected to use the bot for a maximum of two weeks.
-We conducted two interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery.
-These semi-structured interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses.
+Upon form submission, participants received ‘welcome messages’ from CataractBot.
+The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language.
+Upon receiving a response, participants were briefed on the icons and expert verification system.
+Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage.
+After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “Feel free to ask any cataract surgery related questions” until one week post surgery, when their access was automatically revoked.
+We conducted two semi-structured interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery.
+These specific days were chosen because they coincided with the patient’s required hospital visits.
+The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses.
 Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery.
-These specific days—surgery day and seven days post-surgery—were chosen because they coincided with the patient’s required hospital visits.
 After each interview, participants were requested to fill out a chatbot usability form.
-They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness.
-These adapted questions were sourced from metrics outlined in prior usability evaluations of healthcare bots (Abd-Alrazaq et al., 2020).
+They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.2).
+These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots; previous work used scales that included aspects not relevant to our study, such as the chatbot’s visual design (Abd-Alrazaq et al., 2020; Holmes et al., 2023).
 Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
 All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages.
-(Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while the exit interviews (seven days post-surgery) were conducted either in person at the hospital or through pre-scheduled telephone calls.
-This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with surgery preparations and the procedure itself taking 45 minutes.
+(Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while seven days post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls.
+This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking 45 minutes.
 Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time.
@@
 5.2. Procedure for Doctors and Coordinators
-In addition to recruiting patients and attendants, we recruited four doctors, each regularly performing cataract surgeries.
-Among them, three served as operating doctors, one as an escalation doctor and KB Update expert.
-Additionally, two patient coordinators were recruited, both specializing in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator.
-The escalation doctor, escalation coordinator, and KB update expert held senior positions in the hospital and carried administrative responsibilities.
-The KB update expert is part of the research team, and was actively involved in curating the knowledge base.
-This expert played a pivotal role in recruiting all other doctors and coordinators individually through one-on-one discussions.
-To ensure preparedness, all participating doctors and coordinators attended one of two one-hour training sessions conducted by three researchers.
-During these sessions, CataractBot was demonstrated, and participants were onboarded onto the platform.
+In addition to patients and attendants, we recruited four doctors and two patient coordinators.
+Among the doctors, three served as operating doctors, and one as both an escalation doctor and KB update expert.
+Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator.
+The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities.
+To ensure preparedness, all participating experts attended one of two one-hour training sessions conducted by three researchers.
+During these sessions, CataractBot was demonstrated, and they were onboarded onto the platform.
 The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit.
-The first session revealed several bugs, and feedback from doctors and coordinators was collected to enhance the bot.
-For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in tagged WhatsApp messages (Figure 1E).
-After using CataractBot for two weeks or longer, the participating doctors and coordinators were interviewed by the first author.
+The first session revealed a few bugs, and feedback was collected to enhance the bot.
+For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E).
+After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author.
 These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients.
 The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.
-None of the participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation.
 5.3. Participants
 A total of 31 patient-attendant pairs (Table 1), consisting of 19 patients and 30 attendants, took part in our study.
-Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate.
+Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate but their attendants used the bot.
 One patient was not accompanied by an attendant.
 The average age of the 19 participating patients (3 females) was 58.88.01 years.
-They exhibited a diverse distribution in preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and highest education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters).
+Although females comprised 32.7% of our study participants, only 15.8% of the patients were female.
+This aligns with prior work showing a gender disparity in accessing cataract surgery in India (Prasad et al., 2020; Ye et al., 2020).
+Patients had diverse preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters).
 In contrast, the attendants were younger, predominantly fluent in English, and well-educated.
@@
 Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation.
-Additionally, four doctors (three as operating doctors, one as escalation doctor, and one as KB update expert) and two patient coordinators (one as coordinator and one as escalation coordinator) participated in the study (Table 2).
-Note: The roles of escalation doctor and KB update expert were performed by the same doctor.
+Additionally, four doctors and two patient coordinators participated in the study (Table 2).
+Note: There was a partial overlap between participants in this study and the formative study, with two doctors and one coordinator participating in both.
 | PId | Age | Sex | Language |
@@
 Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied.
-The first author translated and transcribed all interviews (totalling 11.5 hours) with patients, attendants, doctors and coordinators, into English soon after they were conducted.
-Throughout the study, all authors regularly engaged in discussions to review observations and recent interviews, fostering a nuanced understanding of bot usage and practices while identifying emerging themes of interest.
-For the analysis of interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line.
+The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted.
+Throughout the study, all authors regularly engaged in discussions to review observations and interviews.
+To analyze interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line.
 Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)).
@@
 5.5. Positionality
-All authors are of Indian origin, and currently reside in Bengaluru, India.
+All authors are of Indian origin, and currently reside in Bangalore, India.
 Among them, two identify as female and the rest as male.
-One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly.
-Their involvement spanned designing the study, shaping research questions, providing feedback on initial CataractBot versions, contributing to knowledge base curation, and serving as the KB update expert.
-Another author is affiliated with the Quality Control and Patient Safety team, who ensured that any risk of harm to the patients was mitigated.
+One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital.
+Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the KB, and ensuring that any risk of harm to the patients was mitigated.
+The ophthalmologist also served as the KB update expert.
 Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development.
@@
 6. Findings
-Overall, 31 patient-attendant222We use the terms ‘participants’ or ‘information-seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators.
-pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair.
-Among the 19 participating patients, only 12 interacted with the bot by sending one or more messages, while all 30 participating attendants, except one, messaged the bot.
-Patients sent a total of 89 messages (7.45.0 message/patient), while attendants sent 254 messages (8.86.8 message/attendant).
-Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others.
-Doctors directly approved 69.8% (157) of the bot-generated answers to medical questions and provided corrections for 29.3% (66) of the answers, while patient coordinators directly approved 58.6% (51) of logistical answers and provided corrections for 40.2% (35) of the answers.
-Note: Only two medical questions and one logistical question did not receive any response from the experts, while for eight questions (five medical and three logistical) an expert responded with ‘No’ to the poll without providing a correction.
-Below, in our findings, we delve into how, why, and when participants used CataractBot, examine the impact of novel features, and conclude by addressing privacy and trust concerns in the context of a chatbot in a medical use case.
+Overall, 31 patient-attendant pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair.
+Out of the 343 messages, 225 were medical questions, 87 logistical, 27 small talk, and 4 others.
+Doctors directly approved (said ‘Yes’ to) 69.8% (157) of the bot-generated answers to medical questions, while patient coordinators directly approved 58.6% (51) of logistical answers.In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it in the future.
+Below, in our findings, we delve into how participants used CataractBot (answering RQ1 and RQ3), examine the impact of novel features introduced in CataractBot (answering RQ2), and explore privacy and trust implications of using an LLM-powered chatbot (answering RQ1).
 6.1. CataractBot Usage
-5.1.1. WHY
-Patients and attendants used CataractBot for various reasons, such as seeking clarification on answers provided by the expert, addressing questions they had forgotten to ask the doctor or coordinator, inquiring about matters they felt less comfortable asking the expert directly, resolving emergency situations, and obtaining status updates.
-In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it (or a similar bot) in the future.
-Ten participants highlighted the bot’s helpfulness in recalling information provided by the doctor/coordinator during in-person consultations, either due to forgetfulness or incomplete understanding.
-As A13 stated: “We are human minds… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed, “I need not remember everything.” (A20).
-The value of persistent information in the context of a text-based bot has been underscored in prior research (Jain et al., 2018).
-Despite knowing the answers to certain questions, six participants mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17).
-Doctors highlighted that many patients/attendants now come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits to the hospital: “When you’re going to a surgery… there are ten things to ask.
-And when you are in front (of the doctor), you’ll remember only two things.
-Most of the patients nowadays come with a list of questions.” (D1).
-CataractBot has the potential to alleviate this burden of remembering every important question.
-Also, this flexibility of using CataractBot at one’s convenience–“as and when I get anything on my mind to ask” (P17)–was noted.
-The asynchronous nature of the conversation is beneficial, as information seekers do not have to constantly wait, either physically or over a call, for an expert’s response.
-A majority of participants (86.4%) agreed that the bot responded quickly.
-On average, the bot took 11.810.2 seconds to respond.
-Expert verification took 158.7168.4 minutes; we found verification without edits (151.1171.4 minutes) was faster than verification with edits (174.5161.4 minutes) which followed by 191.3172.4 minutes to provide the correction for the bot.
-This asynchronous feature not only benefited patients but also motivated experts to engage with the bot.
-The minimal time required to verify and edit an answer is another motivating factor.
-“I can do it anytime… I do not have to think about phrasing and grammar… The bot is going to make it simpler and readable.” – D2.
-Three participants noted the challenge of obtaining “precise information” from external sources like Google, prompting them to use the bot.
-Extracting accurate and relevant information from Google search results and formulating effective search queries is non-trivial (ter Hofstede et al., 1996), particularly for individuals with low (tech-)literacy.
-Moreover, due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), such individuals often hesitate to ask questions.
-Experts acknowledged that educated individuals have access to alternative information sources, and are comfortable posing questions.
+6.1.1. Reasons for (Not) Using CataractBot
+Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates.
+Doctors stated that most information seekers now-a-days come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits.
+CataractBot can alleviate the burden of remembering every question.
+Ten participants highlighted that the bot helped in recalling information provided by the doctor/coordinator during in-person consultations, which was either forgotten or incompletely understood.
+As A13 stated: “We are humans… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed (similar to (Jain et al., 2018)), “I need not remember everything.” (A20).
+Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), individuals often hesitate to ask questions.
+Experts acknowledged that educated individuals have access to other information sources, and are comfortable posing questions.
 In contrast, “less educated, local language speaking patients” are the ones for whom CataractBot “can do wonders” (D3).
-“They have lots of doubts… even if I talk to them in their native language, they hesitate to ask… because it sounds silly… Could be because of the lower socio-economic background, or education.” – D1.
 Five participants, including three with Grade-12 education, exclusively relied on CataractBot for information (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online.
-Another expressed that inquiring with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature:
-“(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” – A4.
-Although explicitly mentioned during onboarding that CataractBot should not be used during emergencies, three patients/attendants attempted to use it to address emergencies, such as allergic reactions and taking the wrong medications.
-This raised concerns among two experts as CataractBot is not equipped to diagnose or handle emergencies.
-Expert responses are not immediate, and patient queries may lack the necessary details for the expert to provide an appropriate response without further clarification.
-“If the patient has a little redness… and is supposed to take a flight… the bot may say “use the same medication no need to worry”, or it says “it’s an emergency come to the hospital, cancel the flight… In either scenario, it’s going to be difficult, because you don’t know what the redness is due to.” – D1.
-The doctor suggested that the bot should be capable of detecting emergency situations, refrain from providing a response, and instead advise users to call the hospital.
+Another expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4).
+Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), in spite of knowing certain answers.
+The asynchronous feature was helpful, as the information seekers did not have to constantly wait, either physically or over a call, for an expert’s response.
 Seven participants emphasized that the bot saved them time, as it eliminated the need to visit or call the hospital.
-Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as A4 stated:
-“Yesterday she (coordinator) called… she told to come tomorrow, that’s all.
-I was about to ask a few questions but then she cut the call.
-Then I called back.
-It was busy, may be because she was talking to others.” – A4.
-Despite the compelling reasons for using the bot, a few participants used it minimally.
-Out of the 49 participants, 9 did not use the bot at all, while 7 sent 1-2 messages to the bot.
-Six participants mentioned that their in-person interactions with the doctor and patient coordinator at Sankara Eye Hospital sufficed to address their queries, reducing their reliance on the bot.
+Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4).
+Also, information seekers found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question (Appendix A.2), mainly due to their familiarity with WhatsApp.
+Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), lowered barriers to initiating conversation.
+Experts used CataractBot driven by their commitment to help patients.
+“I want to help my patients… I am least bothered about the timing. If he is having a problem, he should be answered… after all, we are living for the patient’s sake.” – D2.
+Experts also appreciated the intermediary role of the bot, which introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
+Despite the compelling reasons to use the bot, out of the 49 information seekers, 8 did not use the bot sending two or less messages.
+Six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries, reducing their reliance on the bot.
 Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions.
-This perception could stem from their previous experience with cataract surgery, undertaken six months or more ago for their other eye.
-The patient coordinator added that abundant cataract surgery related information is readily available online, unlike specialised surgeries.
-This suggests that CataractBot has the potential to play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers.
-Additionally, a similar bot could prove valuable for more complex (surgical) procedures.
-Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and/or experiencing delays in responses.
-A total of 36 (10.5%) messages pertained to requests for status updates—seeking information about the surgery time and discharge time.
-Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, lacking access to patient details and resulting in IDK responses for the non-generic questions.
-In such instances, the patient coordinator later provided appropriate responses.
-While participants acknowledged the appropriateness of schedule-related IDK responses due to the lack of access to patient details, for other queries, they suggested providing a “tentative response” rather than a blank IDK.
-For example, for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response, adhering to the cautious approach advised by hospital experts due to medical and legal concerns.
-However, A10 wanted a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”.
+This perception stemmed from their previous experience with cataract surgery of their other eye.
+This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers.
+Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and experiencing delays in experts verifying responses.
+A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time.
+Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in IDK responses for such questions.
+The patient coordinator later provided appropriate responses.
+While participants acknowledged the appropriateness of IDK responses due to the lack of access to patient details, they suggested providing a “tentative response” if possible.
+E.g., for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response.
+However, A10 expected a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”.
 From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context.
-E.g., A21 asked a complex, personalized, and high-stakes question: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check.”.
-This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor would know her condition… I don’t want to come and visit the doctor [again]… it will waste my 2-3 hours.” (A21).
+E.g., A21 asked: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check”.
+This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21).
 For time-sensitive questions, A1 and A9 desired “quick response” from the bot.
@@
 “The dosage… every 2 hours take 1 drop.
-If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” (A1).
-This was reflected in the usability survey as well, with 27.2% of participants selecting ‘Disagree’ or ‘Strongly Disagree’ regarding the bot’s grasp of medical knowledge.
-Experts used CataractBot driven by their sincere commitment to help patients.
-As D2 put it “I want to help my patients… I am least bothered about the timing.
-If he is having a problem, he should be answered… after all, we are living for the patient’s sake.”.
-Moreover, the asynchronous nature of the bot provided experts with the flexibility to offer verifications at their convenience.
-Finally, experts appreciated the intermediary role of the bot.
-The bot-in-the-middle model introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
-5.1.2. HOW
-The workflow employed by our patients and attendants varied, with time-sensitivity being a crucial factor.
-For time-insensitive questions, participants tended to consume the unverified response, wait for the expert’s verification, and later, consume the (updated) verified response.
-Our participants demonstrated a clear understanding of the bot’s workflow and expressed being “okay to wait” since the response was being verified by “busy doctors”.
-“I know that it will take its own time… as its doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right?
+If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” – A1.
+6.1.2. Workflow
+Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour.
+The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.” The workflow employed by participants varied, with time-sensitivity being a crucial factor.
+A majority of participants (86.4%) agreed that the bot responded quickly.
+On average, the bot took 11.810.2 seconds to respond.
+Expert verification took 162.9172.3 minutes.
+We found verification without edits (151.1147.1 minutes) to be significantly faster than verification with edits (191.3172.4 minutes) with t(265)=2.7, p¡0.05.
+For time-insensitive questions, participants were “okay to wait” for the expert’s verification, since the response was being verified by “busy doctors”.
+This demonstrated their clear understanding of the bot’s workflow.
+“I know that it will take its own time… as it’s doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right?
 If that is the case, then bot is not required.
@@
 A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities.
-They would ask a question, leave their phones, and check them later when free.
-In such cases, notifications played a key role in the information-seeking workflow.
-The notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
-For questions where participants received an unsatisfactory IDK response, they sought information from other sources, including Google and asking friends and family.
-CataractBot provides both an immediate LLM-generated response and an asynchronous expert-verified response.
-For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the inherent delay associated with obtaining a verified response.
-Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions.
-At times, they first used Google and subsequently posed the same question to the bot, hoping for a quick, verified response from the expert.
-In only two instances involving critical questions, participants relied on the bot’s response and took interim actions before getting verification from the expert due to the time-sensitive nature of the questions.
-According to D4, without CataractBot, patients might have resorted to Google, risking following inaccurate advice, given the unavailability of doctors for immediate consultations.
-However, CataractBot improves the likelihood of receiving accurate responses, as it is utilizing a custom knowledge base.
-Irrespective of time sensitivity, certain questions such as “What time i should come?” (P19), or “I get operated on 5th jan, can i do exercises as alom vilom etc?” (P27), could only be adequately answered by hospital experts.
-Participants relied on CataractBot for such questions.
-We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4:
-“It takes only about 15-20 seconds to answer each question… they are easy.
-It was not really eating into my time where I had like hundred messages [to respond] every day… It’s much easier than replying to my mailbox.” – D4.
-None of the doctors adhered to a specific time for responding; they verified queries while awaiting the next patient, during lunch or coffee breaks, post-OPD towards the end of the day after 4 pm, and/or after reaching home.
-The only exception was during days reserved for surgeries.
-All doctors emphasized a no-phone-use policy during surgical procedures, emphasizing the need for minimal messages on those days.
-Even the coordinators suggested adding a working-time parameter for CataractBot, restricting message delivery to the specified period.
-As the experts (except one coordinator) used CataractBot on their personal WhatsApp accounts, it blurred the boundary between their home and work (similar to (Mols and Pridmore, 2021; Wang et al., 2020)).
-On a positive note, WhatsApp minimized the learning curve for experts.
-However, our experts pointed out that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp.
-I just see messages in the important groups, that’s it.
+They would ask a question, leave their phones, and check the answer later when free.
+In such cases, notifications played a key role as the notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
+If participants received an IDK response, they sought information from other sources, including Google and asking friends and family.
+For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the delay in obtaining an expert verified response.
+Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27).
+According to D4, without CataractBot, patients might have resorted to “Google” and risked following inaccurate advice, given the unavailability of doctors for immediate consultations.
+CataractBot improves the likelihood of receiving accurate responses, as it utilizes a custom KB.
+We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.” The majority of responses (31, 10.8%) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm.
+The second-highest responses (28, 9.8%) occurred at 1-2 pm, during their lunch break.
+This usage pattern was described as:
+“I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
+As five experts used CataractBot on their personal WhatsApp accounts, it minimized the learning curve but also blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020).
+Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp.
+I just see messages in the important groups.
 Otherwise my WhatsApp has 5000+ [unread] messages… college groups, family groups.” (D2).
 Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
-Additionally, experts managed multiple patients simultaneously through CataractBot.
-It could potentially compromise the quality of patient care, as observed by Johansson et al.
-(2020a).
-For instance, an expert mistakenly tagged the wrong query while adding a correction, leading to confusion for the patient.
-Learning Curve: A majority of participants found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question.
-Several factors contributed to the ease of learning, with participants’ familiarity with WhatsApp being the most significant.
-Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), facilitated a quick start.
-Encouraging patients to use the bot as part of onboarding and providing detailed instructions proved essential in minimizing the learning curve.
-Contrary to initial expectations that understanding the three icons might be complex, participants found it self-explanatory:
-“That is very simple… Even my dad understood.
-I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… Even for the tick mark also you are sending notification.
-If you would have done it in the background without a message, he would have not noticed… If MY dad was able to use, I think anyone can use, because even attending a phone call is a bit difficult for him.” – A13.
-Explicit notification messages related to the experts’ verification step were appreciated by multiple participants.
-5.1.3. WHAT
-To conduct a more detailed analysis of the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)).
+6.1.3. Question Type and Style
+To analyze the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)).
 Note: Several questions fell into multiple categories.
-Figure 4 illustrates the relative significance of these categories before, during, and after the day of surgery.
-Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia” and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”).
-One day before the surgery, with the coordinator finalizing the schedule, questions related to the “schedule” peaked, as patients/attendants awaited calls from patient coordinators.
-On the day of surgery, with patients starting a course of eye drops to prevent infection and reduce inflammation, questions about medication and dosage rose in importance, as did those about the expected discharge procedures.
-After the surgery, participants were particularly concerned about pain management and sought guidance on dos-and-don’ts during recovery, including information on exercise and travel.
-Regarding question structure, our participants minimally used double-barreled questions—single messages containing multiple questions or topics—unlike prior work (Li et al., 2023).
-This difference may be attributed to demographic variations, as Li et al.
-(2023) focused on doctors who aimed to maximize productivity by asking multiple questions together.
-Only four double-barreled questions were sent by three of our participants, and in each instance, the two questions were related.
-E.g., “What is the Valsalva maneuver and why is it harmful after cataract surgery?” (A21).
-The bot answered both questions in each case.
-Anthropomorphization: A total of 39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot.
-Participants often addressed the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’.
+Figure 4A illustrates the relative significance of these categories before, on, and after the day of surgery.
+Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia”, “lens”, and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”, “documents to bring for surgery”, and “take blood thinners on the day of surgery”).
+One day before the surgery, questions related to the “schedule” peaked, as information seekers awaited calls from patient coordinators about their surgery time.
+On the day of surgery, questions focused on discharge procedures and post-surgery medication and dosage.
+After the surgery, participants were concerned about pain management and sought guidance on dos-and-don’ts during recovery (e.g., “washing hair”, “doing yoga”, and “watching TV”).
+39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot with addressing the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’.
 Many queries were prefaced with greetings (like ‘hi’) and included acknowledgements (like ‘okay’), contributing to the humanization of the bot.
-In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews: “How should it be during the day?
-Please let me know.” – P14, in Kannada.
-Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as “The date of the surgery is tomorrow.
-We will leave home at 8 am from Ayyappa Nagar.” (P29).
-Such messages not only caused confusion for the bot but also for the experts.
-The classifier erred in not classifying these messages as small talk; instead, it specifically labeled them as medical to ensure that user queries were not overlooked.
-The observed behaviors can, in part, be attributed to participants’ awareness that doctors and patient coordinators are on the other side reviewing these messages.
-Participants attempted to direct their queries or provide status updates to them.
-5.1.4. WHO
-Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=-2.04, p¡0.05.
+In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews.
+Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as P29 telling CataractBot that “The date of the surgery is tomorrow.
+We will leave home at 8 am from Ayyappa Nagar.”
+6.1.4. Multiple Stakeholders
+Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=2.04, p¡0.05 (Figure 4C).
 Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert.
-This highlights the significance of the bot supporting multiple information-seeker and information-verifier roles.
-Further, in two cases, patients came accompanied with multiple attendants, suggesting the need for the bot to accommodate not just a patient-attendant pair, but rather, a flexible number of information-seekers per patient.
-During our interviews, we identified several reasons for the patients to use the bot minimally, particularly related to demographics.
-Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants – for example, 2 attendants, who were software testers by profession, mentioned that they “kept pushing some questions” (A21) to identify bugs and help improve the system.
-Moreover, a subset of patients enrolled on the bot (7) placed trust in the hospital to provide them with all the required information and did not actively seek information themselves.
+This highlights the significance of the bot supporting multiple information seeker and information verifier roles.
+During our interviews, we identified a few reasons for the patients to minimally use the bot, particularly related to demographics.
+Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants These patients placed trust in the hospital to provide them with all the required information and did not actively seek information themselves.
 For instance:
-“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (or know) about what will be this, what will be that”.” – A3.
-This aligns with prior findings indicating less-educated older populations may have reduced information needs (Ma et al., 2023).
-Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced interaction with the bot.
-Patients found value in local language interaction and speech as input. A13 and A14 noted that the bot’s Indic language support helped their parents (patients) to use the bot independently.
-“Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
-Quantitatively within patients, P13 sent the maximum number of messages to CataractBot at 16, comprising 15 text messages and 1 tap on a related question.
-The second-highest number of messages (15 messages) was sent by P19, the youngest patient at 41 years and a well-educated background holding a Master’s degree.
-Following that, P26 and P31 sent 13 and 9 messages respectively, with the majority being audio messages (8 and 6 audio messages, respectively).
-Interestingly, the involvement of additional stakeholders had a negative impact, as patients tended to offload the knowledge-seeking ‘work’ to their active attendants.
-For instance, “I have been using it.
-I told him [patient] what (answers) I got.
-So maybe he did not feel the need (to use it).” (A20).
-Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts.
-Among the 223 answered medical questions, 76.2% were verified by operating doctors, and 23.7% by the escalation doctor.
-In contrast, only 26.7% of the 86 answered logistical questions were verified by the coordinator, and 73.2% by the escalation coordinator.
-This discrepancy occurred because the patient coordinator participated in the study using their work phone, which was only active from 9 am to 6 pm, as they left their work phone at the hospital.
-Moreover, they were on emergency leave for 11 days during the study period.
-Doctors mentioned having 3 or more days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours.
+“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. – A3.
+Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced bot interaction.
+Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—for instance, “I have been using it.
+I told him [the patient] what (answers) I got.
+So maybe he did not feel the need (to use it).” (A20)—so the involvement of additional stakeholders had a partially negative impact..
+Moving on to expert stakeholders, both the doctors and patient coordinators highlighted the significant role of escalation experts.
+Among the medical questions, 76.2% were verified by operating doctors and 23.7% by the escalation doctor.
+Doctors mentioned having 2-3 days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations.
 Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators.
@@
 As a result, escalation experts lacked knowledge of patients’ context, and hence provided generic responses.
-5.1.5. WHEN
-Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour.
-Notably, the maximum number of messages (56) were sent between 10-11 am, despite participants receiving reminder messages from CataractBot at 7:30 am and 4 pm daily.
-The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.”.
-The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery, for a maximum total of 15 days.
-However, due to a bug in the code to disable the bot seven days post-surgery, the bot remained active.
-Moreover, 3 participants had their surgeries rescheduled due to increased blood sugar levels and travel, leading them to use the bot for more than seven days before their surgery.
-Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 5).
-Four participants requested the bot to be available beyond the seven days post-surgery period.
-They cited that post-operative medications and precautions would continue for a month, and having the bot available during that time would serve as reassurance and align with the expected level of service from the hospital.
-A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery.
-For the remaining days, across 49 users, the total number of messages was less than 10 messages/day.
-The highest number of messages in a single day were sent on the day of the surgery (76), and 156 messages sent before the surgery day and 111 post the surgery day.
+Among the different stakeholders, bot interactions varied.
+For patients and attendants, the bot operates as a user-driven platform where they initiate conversations by asking questions.
+In contrast, the bot functions as a bot-driven system for experts, initiating and driving verification-related conversations.
+Patients and attendants appreciated that CataractBot “does not disturb” them and provides information solely based on the questions asked.
+However, a few participants suggested the bot should proactively offer relevant information.
+For example, four attendants wished for regular updates from the chatbot during the surgery while waiting anxiously outside.
+A9 wanted reminders from the chatbot to ensure patients “take eye drops on time” post-surgery.
+6.1.5. Question Time and Response Time
+The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery.
+However, due to a bug in the code, the bot was never disabled.
+Moreover, 3 participants had their surgeries rescheduled, leading them to use the bot for a longer period.
+Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 4B).
+A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages sent on the day of the surgery (76).
+156 messages were sent before the surgery day and 111 post the surgery day.
 We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05.
 Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05).
-This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety.
-This finding aligns with prior research indicating that patients and attendants tend to seek information closer to their doctor’s visit (Beisecker and Beisecker, 1990).
+This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety, aligning with prior research (Beisecker and Beisecker, 1990).
 The higher pre-surgery usage may also be due to the novelty effect.
-With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the scheduled surgery day (Figure 4), as anticipated.
-Despite the high pre-surgery usage, four participants mentioned that since they did not have access to CataractBot in the early stages of diagnosis, before scheduling the surgery, information-seeking patients and attendants had already clarified most of their questions from other sources, including Google, friends and family, and hospital staff.
-Particularly, three patients were scheduled for surgery in the next 2 days, so their pre-surgery usage was limited to less than 2 days.
-In contrast, experts expressed the view that questions on the day of surgery might be unnecessary and suggested encouraging patients/attendants to talk directly with doctors and coordinators face-to-face:
-“There are multiple questions asked on the day of surgery, “when he is going to be discharged?” Asking a question when you are in the hospital itself, looks a little too much… There is no need for you to use the bot… absolutely avoidable.
-You can walk 10 steps and go to the staff and ask.” – D1.
-Regarding the usage patterns of experts, responses were verified throughout the day and even late at night (1 am to 2 am).
-The majority of responses (31) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm.
-Also, the reminder message was sent to the experts at 4 pm.
-Note: Every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours was sent to both the operating and escalation experts.
-The second-highest responses (28) occurred at 1-2 pm, during their lunch break.
-This time-of-day usage pattern was described as:
-“I use my mobile data only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
-Experts refrained from using their phones in front of patients/attendants due to the negative perception associated with using phones for personal reasons.
-As C1 described: “If I see my phone (in front of a patient)… they will not understand what I am doing, they will think okay, simply they are watching videos or texting.”.
-In a busy hospital context, physical patients always took priority.
-Hospitals in India often face high patient loads.
-Experts are typically overwhelmed with more patients than they can handle.
-During working hours, there is usually a long queue of patients waiting to consult with experts, making it challenging for experts to use their phones comfortably to verify bot responses.
-This also contributed to many medical and logistical queries being escalated.
-We observed varying opinions on using the chatbot after work hours; two experts were reluctant to use the bot outside of working hours, while four were in favor of using it “anytime”.
+With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the surgery day (Figure 4B), as anticipated.
+In contrast, experts expressed that questions on the day of surgery might be unnecessary:
+“Multiple questions were asked on the day of surgery, “when he is going to be discharged?” … You are in the hospital… Why to use the bot!… You can walk 10 steps and go to the staff and ask.” – D1.
+Regarding the usage patterns of experts, responses were verified throughout the day.
+The only exception was during days reserved for surgeries.
+All doctors emphasized a no-phone-use policy during surgeries, stressing the need for minimal messages on those days.
+A coordinator suggested adding a working-time parameter for CataractBot to restrict message delivery to specified periods.
+Experts refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons.
+During working hours, the long queue of waiting patients made it challenging for experts to use their phones.
+These contributed to several queries being escalated.
+Opinions on using the chatbot after work hours varied: two experts were reluctant, while four were in favor of using it “anytime”.
 6.2. Features
-5.2.1. Expert-in-the-loop
-Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others.
-Note: 13 questions were marked as ‘Send to Coordinator’ by one of the operating doctors.
-On manual analysis, the escalation doctor confirmed that all of them were indeed medical questions, such as “after how many days can I shave”, but were simple enough to be answered by the patient coordinator.
-For our analysis, we consider them as logistical questions, as they were finally answered by the patient coordinator.
-Doctors approved 69.8% (157) of the bot’s generated answers directly and provided corrections for 27.1% (61) of responses, while patient coordinators approved 58.6% (51) of responses to logistical questions directly and provided corrections for 36.8% (32) of responses.
-For 5 medical queries and 3 logistical queries, the expert answered the poll response with a ”No” but did not provide the correction.
-The high number of logistical corrections was mainly because a majority of logistical questions (62.1%) were about scheduling, such as “time of surgery” and “discharge time”, for which the LLM cannot provide an appropriate response as it was not connected with the hospital’s patient management system.
-The length of these corrections in characters varied with 154.4105.0 for medical questions and 51.042.6 for logistical questions.
-Despite having a curated knowledge base and an expanding knowledge base with the addition of expert corrections after KB Update Expert approval, the bot provided IDK responses to 16.6% of questions.
-Even after excluding scheduling-related questions, 23.9% of responses were corrected by the expert.
-Further analysis revealed that their suggested edits across all questions mainly fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), redirecting the patient to check with other sources (8.8%), and removing extra unnecessary information (3.7%).
-The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by our participants, serving as the key reason for their engagement with the bot.
-“That’s why I gave my WhatsApp (number to the coordinators while onboarding). Because I can talk directly to the doctor.” – A2.
-However, since participants were aware that these messages were being reviewed by doctors and coordinators, four participants used the bot to send direct messages to them instead of asking questions.
-For instance, P19 messaged “Give me call?”
-Experts did not have access to the final updated answer sent to patients/attendants after their suggested edits.
-We made this design decision for CataractBot to limit verification to a single round, minimizing the workload on experts.
-However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understand and give that as an answer?
-I did not know.” (D3).
+6.2.1. Experts-in-the-loop
+The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by participants, serving as the key reason for their engagement with the bot.
+For instance, “That’s why I gave my WhatsApp (number to onboard the bot), because I can talk directly to the doctor.” – A2.
+Contrary to initial expectations that understanding the icons might be complex, participants found it self-explanatory:
+“That’s simple… Even my dad understood.
+I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… If MY dad was able to use, I think anyone can.” – A13.
+However, since participants knew that these messages were being reviewed by doctors and coordinators, four used the bot to send direct messages to them instead of asking questions.
+For instance, P19 messaged “Dr.
+Give me a call?”
+The length of expert corrections in characters varied significantly (t(91)=5.2, p¡0.05) with 152.4104.9 for medical questions and 50.942.6 for logistical questions.
+Further analysis revealed that their suggested edits fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), asking the patient to visit the hospital (8.8%), and removing extra unnecessary information (3.7%).
+Finally, experts did not have access to the final updated answer sent to information seekers after their suggested edits.
+This design decision limited verification to a single round, minimizing the workload of experts.
+However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understood and give that as an answer?
+I donno.” (D3).
 In the CataractBot system, only the KB update expert had access to the updated answer sent to the information seeker.
-Finally, asynchronous expert correction may not be effective in all instances, especially when patients ask critical time-sensitive questions, such as allergic reactions after surgery.
-In such cases, two patients suggested having a fallback option of a human agent available over a call to assist: “similar to Amazon… if the query is not resolved after two-three attempts, it should just ask you to contact [call] the coordinator.” (P21).
-Even the escalation doctor D4 suggested a similar solution.
-However, prior work (Weisz et al., 2019) suggests that in such solutions, users tend to bypass the automated agent to directly connect with the human agent over a voice call.
 6.2.2. Input Modality
-Patients and attendants were enabled with three ways to interact with CataractBot: typing messages in their preferred language, recording audio, or tapping one of the three related questions suggested by the bot.
-However, experts could only interact with CataractBot using text.
-This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023).
-Out of the total 343 messages sent by patients and attendants, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions.
-The 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 245 messages (163 text, 17 audio, and 74 taps).
-One-way ANOVAs on message types (after Greenhouse-Geisser correction for sphericity violations) found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=.46).
-Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01), ‘text’ and ‘taps’ (t(29)=3.0, p¡.01), and not for ‘audio’ and ‘taps’ (t(29)=1.9, p=.11).
+Out of the total 343 messages sent by information seekers, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions.
+19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 4C).
+One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46).
+Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants.
 The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated.
@@
 For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
-Moreover, post-surgery limitations on screen use made speech input and audio output the preferred modality.
-34 participants did not use the audio modality at all, while most participants (31) selected one of the related questions.
+Post-surgery limitations on screen use also made speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery.
+34 participants did not use the audio modality at all, while 31 selected one of the related questions.
 As P7 stated: “I used the audio message once.
 After that there was no need for me… I just selected from the options.
-Click here (shows the phone) and automatically question will come, and click one question and get the answer.”.
-While patients and attendants found the related questions feature convenient and frequently used it, there were instances of dissatisfaction.
-On eight occasions, the answers to suggested questions were not available in the knowledge base, resulting in IDK responses.
-This issue arose because the CataractBot system use an LLM call to suggest three related questions based on the users’ last question, independent of the KB.
-This approach aimed to uncover gaps in the KB by not restrict questions to the curated KB, and introduce a personalized touch to the interaction by suggesting questions relevant to the ongoing conversation.
+Click here (shows the phone) and automatically questions will come, and click a question and get the answer.”.
+While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction.
+On eight instances, the answers to suggested questions resulted in IDK responses.
+This issue arose because the CataractBot system uses an LLM call to suggest three related questions based on the users’ last question, independent of the KB.
+This approach aimed to uncover gaps in the KB by not restricting questions to the curated KB, and introduce personalization by suggesting questions relevant to the ongoing conversation.
+With respect to experts, they could only interact with CataractBot using text.
+This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023).
 6.2.3. Multilinguality
@@
 None of the participants changed their initially selected language.
-The inclusion of four Indic languages, along with speech input functionality, ensured effective use of the bot by users with minimal literacy and tech proficiency.
 Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions.
-Patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions.
-In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and selected 14 related questions.
+In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions.
+In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total.
 A similar trend was observed among attendants.
+Patients found value in local language interaction and speech as input, helping them use the bot independently: “Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful.
+Because now, for elders… it is very hard to use apps and things.” – A13.
 The audio queries in Indic languages were transcribed and translated to English using Azure AI services.
-However, due to variability in dialect across India, existing transcription and translation technologies have been found to be erroneous.
-It is a known issue and prior work (Jain et al., 2018) often relied on human wizards to correct errors caused by language technologies.
-For example, a query by P29 (Kannada) was transcribed and translated to “Krishna immediately said, how long is this kampu?”.
-D3, upon requesting and listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately.
-Experts faced challenged in “deciphering” such messages since they only received the English text, which was sometimes nonsensical.
-They suggested that they should be provided with the original audio recordings of patient queries.
-D4 further recommended that including audio recordings would help extract non-verbal cues (such as anxiety or the presence of other family members), enabling the expert to respond more appropriately.
-It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot.
-We found three participants who sent a text message each by typing an Indic language query using Latin script.
-For one of them, the bot responded with an accurate response because the message contained a code-mix of Hindi and English:
+However, this is not highly accurate due to the variability in dialect across India.
+It is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors.
+For example, a query by P29 (Kannada) was transcribed and translated to “I got kampu?”.
+D3, upon listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately.
+Experts faced challenges in “deciphering” such messages since they only received the English transcription.
+They suggested providing the original audio recordings of patient queries.
+D4 added that this would also help extract non-verbal cues, enabling them to respond more appropriately.
+It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot due to current translation technology limitations.
+We found three participants did that.
+For one of them, the bot responded accurately as the message contained a code-mix of Hindi and English:
 A8: agar opration k baad pain ho raha to kya karna hai?
 Bot: If the patient is experiencing pain after the surgery, they should report to the doctor immediately…
-For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…” and the IDK response.
+For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…”.
 Additionally, even in the selected language, participants with low literacy and language proficiency struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)).
-For example, “Tomorrow cataract operation When should come No phone call received yet?” (P18).
+For example, “Tomorrow operation When should come No phone call received yet?” (P18).
 Such questions were common with 65 questions, of which 26 were asked using audio in an Indic language.
-Further analysis revealed that the speech-to-text engine was mainly responsible for these errors (19 out of 26) due to the different dialects used by the users, while the remaining 7 were translation errors.
+Further analysis revealed that the transcription engine was mainly responsible for these errors (19 out of 26), while the remaining 7 were translation errors.
 Despite these challenges, the robustness of the GPT-4 model enabled the chatbot to answer most (45) of such ill-formulated questions appropriately.
-For instance, in response to the question asked by P18, the bot replied, “The surgical list is usually prepared the previous evening.
-If you haven’t received a call from the inpatient department, please call and confirm your timing.”
 6.2.4. KB Updation
-Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB Update expert (D4) to be added to the knowledge base.
-In cases where the expert edits were accepted, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB Update expert in 49.3% of cases.
-These additional changes by D4 extended beyond information-related aspects, also addressing aspects of “tone and structure” (such as preferring active over passive voice).
-The changes made by the KB Update expert were minimal, with an average relative Hamming distance of 28.2%.
-This hints that the bot effectively incorporated the expert edit suggestions into the final bot response.
-All our experts noticed an improvement in the completeness and accuracy of CataractBot’s generated answers over time, resulting in a reduction of their workload: “With time, I think the answers did get better… I am mostly just validating and saying yes.” (D3).
+Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB update expert (D4) to be added to the KB.
+For 49.3% of these, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB update expert.
+These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active over passive voice and making responses more generally applicable across patients.
+This hints that the bot effectively incorporated the expert edit suggestions into the final response.
+Experts noticed improvements in the completeness and accuracy of CataractBot’s generated answers over time, thus reducting their workload: “With time, I think the answers did get better… I am mostly just saying yes.” (D3).
 However, there was an instance where the KB update expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the KB.
-This led CataractBot provide incorrect information to three users.
-This error was manually corrected in the KB by a researcher on the team.
-Despite efforts to expand the KB, we identified limited knowledge in three medical areas—types of lenses and their comparisons, lab tests and prerequisites, and details related to prescribed medicines—and one logistical area—insurance tie-ups.
-Addressing it requires addition of new documents to the custom KB.
-Hence, D4 suggested the need for a simple way to upload documents to further expand the KB.
+CataractBot provided this to three seekers, before it was fixed.
+Despite expanding the KB, we identified our bot’s limited knowledge in two medical areas—lens types and their comparisons, and lab tests—and one logistical area—insurance tie-ups.
+Addressing these required manually adding new documents.
+D4 suggested the need for a simple way for experts to upload such documents to further expand the KB.
 6.3. Privacy and Trust
 6.3.1. Personalization vs Privacy
-In the context of evolving LLM technology, there is an ongoing discussion about privacy concerns, as LLMs have been shown to memorize training data (including personally identifiable information (PII) like emails and phone numbers) and leak it during inference (Borkar, 2023).
-To address these concerns, CataractBot is intentionally designed to be non-specific to individual patients, thus preserving patient privacy.
-A few participants (3) appreciated that their medical records were private.
-“Using personal health information is… violation of the patient’s personal rights.
-See, if that (health data) is added… external agencies can extract that information and they might use it for some advertising purposes… Personal information or any diagnosis related to any patient should not be shared.” – A17.
+Three participants appreciated that their medical records were not accessible to bot.
+“Using personal health information [by the bot] is… a violation of patient’s rights. If their data is added… external agencies can extract that information and might use it for advertising.” – A17.
 Three educated participants (holding Master’s degrees) considered personalization as “an unnecessary add-on”.
-They raised concerns about potential complications associated with sharing PII data with LLMs, both in terms of legislation and “privacy threads”.
+They raised concerns about potential complications associated with sharing PII with LLMs, both in terms of legislation and “privacy threats”.
 A1, e.g., remarked “I would definitely not want my mom’s data to be out there, all over the internet.”.
-Moreover, the lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two three people I know are going to have their cataract surgery soon… I’ll use it for them.” (A11).
-Since CataractBot neither had access to patient medical records nor was integrated with the hospital’s patient management system, it was not privy to patient-specific medical and logistical details, which resulted in IDK responses for 20 queries related to surgery schedules, updates during the procedure, and medication dosage pre-/post-surgery.
-During interviews, five participants expressed frustration with the IDK responses, and two of them discontinued using the chatbot after encountering a few such responses.
-Participants expected personalized answers to their queries and were disappointed with generic responses.
-There existed an expectation mismatch, fueled by the recent hype and promises surrounding LLMs.
-Participants anticipated more advanced capabilities from such a system.
-In contrast, a participant working in the field of AI had a more balanced perspective: “See, I understand, it is just a machine… It cannot surpass or bypass or completely replace humans.” (A17).
-Another participant, A21, highlighted that providing more information in the chat messages could improve the bot’s ability to deliver a more personalized response.
-However, they acknowledged the challenge in “setting context every time”.
-Eight participants asked for personalized responses from the chatbot, with little concern for privacy.
-They felt that a personalized chatbot could save time and energy, eliminating the need to seek information from other sources or visit the hospital.
+Moreover, lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two people I know are going to have their cataract surgery soon… I’ll use this bot for them.” (A11).
+On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “precise” and “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses.
 This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021).
-Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s like nothing personal, to be hidden or something.
-All have common health issues.” – A14.
-Moreover, two participants, due to the low literacy and tech-literacy, questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
-Educated participants were willing to share their personal medical data if the benefits outweighed the associated costs.
-“If you’re traveling to Dubai and you say “I won’t show you my passport”… they will say “Thank you so much, you can’t board the plane.”… [Similarly] if you don’t want to share YOUR information [to the bot], the bot should remark “If you ask any personal questions the bot will not have access so you may have to call or visit the doctor or hospital.”… somewhere we have to compromise.” – A21.
-Additionally, two participants mentioned that personalization, along with the sharing of patient medical data, should be an “opt-in” feature.
-Individuals should conduct a cost-benefit analysis and make an informed choice based on their preferences.
+Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s nothing personal, to be hidden or something.
+All have common health issues.” (A14).
+Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
+Two educated participants were willing to share their personal medical data if the benefits outweighed the costs.
+Additionally, two participants wanted personalization, along with the sharing of patient medical data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences.
 They believed that a personalized CataractBot would offer more relevant responses and minimize IDK responses.
-With each query, CataractBot provides experts with patient details such as name, age, gender, and surgery date (Figure 1A).
-D3 found this feature “(useful) while giving post-operative information, because (we) won’t remember patient metadata”.
-Four doctors and patient coordinators were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses.
-Such integration would not only reduce their workload by addressing medication and logistic-related questions accurately, but it will also enhance user trust.
-D3 suggested collecting at least basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive.
-So, when they ask “Should I eat and come?”… I don’t know what to tell.” (D3).
-Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of the participants provide detailed information due to time constraints.
-C2 suggested integrating the bot with the hospitals’ EHR (Electronic Health Record) system.
-However, two doctors (D1 and D2) voiced concerns about CataractBot accessing patient data, citing privacy issues and skepticism about the LLMs capability to understand complex patient records accurately.
-“There are a lot of variables involved, including the surgeon, the particular patient, the particular case, the tests… Its too much information to process.” (D1).
-In addition, the KB Update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for all experts and negatively impacting chatbot adoption among experts.
-5.3.2. Bot-driven vs User-driven
-CataractBot operates as a user-driven platform from the perspective of patients and attendants, while it functions as a bot-driven system for experts.
-Information seekers initiate conversations by asking questions and receive responses.
-In contrast, the bot initiates and drives verification-related conversations with experts.
-Patients and attendants appreciated that CataractBot “does not disturb” them, providing information solely based on the questions asked.
-However, a few participants suggested the bot should proactively offer relevant information about the surgery.
-Specifically, four attendants wished regular updates from the chatbot during the surgery, particularly while waiting anxiously outside for the procedure to conclude.
-Similarly, expert D4 suggested that the chatbot should notify patients: “Your discharge summary is ready, you can come and pick it up.”.
-A9 recommended reminders from the chatbot to ensure patients “take eye drops timely” post-surgery.
+“If you don’t want to share your information [with the bot], it should say “If you ask personal questions, I won’t have access, so you may have to visit the hospital.”… Somewhere we have to compromise.” – A21.
+Five experts were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses.
+Such integration would not only reduce their workload, but also enhance user trust.
+D3 suggested collecting basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive.
+So, when they ask “What should I eat and come?”… I ain’t sure what to tell.” (D3).
+Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of our participants provided that information due to time constraints.
+However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately.
+“There are a lot of variables involved, including the surgeon, the particular patient, their case, tests… It’s too much information to process.” (D1).
+In addition, the KB update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
 6.3.2. Reliability and Trust
-Six patients/attendants explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery.
+Six information seekers explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery.
 As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2).
-This contributed to participants’ trust in the bot.
 Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot.
-While the hospital’s reputation instilled trust in the bot, participants gained confidence in the bot’s responses due to the human experts verifying answers, denoted by the green tick icon on verified responses (Figure 1F).
-As A12 stated: “Basically the trust only came when we saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”.
-This ensured that the bot’s response were more reliable and “not fake” (P18), in contrast to a few participants who complained about receiving false information from Google searches.
-Although the expert-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence.
-Four participants noted that every time an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers (Figure 2E) reduced.
-For instance, “After some point when doctors said it is invalid, I was like, okay, should I even trust the bot?” (P19).
+All these contributed to participants’ trust in the bot.
+With respect to bot’s responses, participants gained trust due to the human experts verifying answers (Figure 1F).
+As A12 stated: “Basically the trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”.
+The experts-in-the-loop ensured that the bot’s response were more reliable and “not fake” (P18).
+This is in contrast to a few participants who complained about receiving false information from Google search results.
+Although the experts-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence.
+Four participants noted that when an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers reduced.
+For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19).
 In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert.
-This could be due to the bot’s limited custom KB or the prompting strategy favoring caution, as per doctors recommendation.
+This could be attributed to the limited custom KB, the experts’ high standards for the tone and structure of answers, or our prompting strategy favoring caution.
 A9 pointed: “It said that severe cough is dangerous on the surgery day as it increases eye pressure.
-But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”.
+But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”
 6.3.3. Accountability
-CataractBot provides a persistent written record of expert-patient communication, carrying several implications.
+CataractBot provides a persistent written record of expert-patient communication, which has implications.
 Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery.
-Moreover, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can use the same information to be shared with their doctor at their place, “Okay, this is what was told to us.” So that [communication] becomes a little easier.” (D1).
+Also, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can share the same information with their doctor at their native place… So that [communication with another doctor] becomes a little easier.” (D1).
 However, it also increased the sense of accountability among experts, making them “somewhat nervous”.
 They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4).
-Furthermore, coordinators identified the system’s potential to identify logistical bottlenecks in the hospital.
-For example, if multiple patients are inquiring about their surgery time, it may indicate a lapse in timely communication, leading them to reach out to the bot.
 7. Discussion
-In this paper, we propose CataractBot, an LLM-powered WhatsApp-based expert-in-the-loop conversational agent designed to address the information needs of patients undergoing cataract surgery.
-Our user study revealed positive evidence of its usability among different stakeholders, including patients, attendants, doctors, and coordinators.
-Patients and attendants appreciated CataractBot for its ability to answer questions anytime and provide verified responses, contributing to the establishment of trust.
+In this paper, we propose CataractBot, an LLM-powered WhatsApp-based experts-in-the-loop chatbot to address the information needs of patients undergoing cataract surgery.
+Our user study revealed positive evidence of its usability among the different stakeholders.
+Patients and attendants appreciated CataractBot for its ability to answer questions anytime with verified responses, which helped establish trust.
 Additionally, it streamlined communication, saving time by reducing the need for hospital calls/visits.
-Notably, individuals with low literacy and low tech literacy found it easy to use because of the robust multilingual and multimodal support.
+Individuals with low literacy and low-tech literacy found it easy to use because of the multilingual and multimodal support.
 Experts commended CataractBot for acting as a facilitator, creating a layer between experts and patients.
-The asynchronous communication feature provided them the flexibility to verify LLM-generated responses at their convenience.
-We discuss key design implications below.
+The asynchronous communication feature provided them the flexibility to verify responses at their convenience.
+We identify critical insights, open questions, and design implications that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale.
+We discuss these aspects below.
+7.1. Chatbots in Healthcare and Beyond
+7.1.1. Experts-in-the-loop
+Several studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, obscured reasoning, and bias (Denecke et al., 2024; Au Yeung et al., 2023; Fogel and Kvedar, 2018; Bérubé et al., 2021; Iliashenko et al., 2019; Kocaballi et al., 2020).
+Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022).
+To address these concerns and comply with OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”, CataractBot incorporates experts-in-the-loop.
+This feature ensured reliability and fostered trust, as users knew their operating doctor was verifying the answers.
+However, instances where users received information from the bot that was subsequently marked as incorrect or incomplete by the expert, their trust in the initial automated answers was reduced.
+To address this issue, one design approach could be to eliminate the initial automated response to the patient and only send expert-verified answers.
+Although this might result in long wait times, which could discourage usage of the bot and prompt patients to seek information elsewhere, it is worth exploring.
+In the future, it would be valuable to compare these two approaches: a hybrid method of synchronous automated answers with asynchronous verification versus a fully asynchronous system with solely verified answers.
+7.1.2. Conversational Design
+Prior research indicates that synchronous telehealth systems can make patients feel rushed and unable to adequately express their concerns (Gordon et al., 2020; Bhat et al., 2021).
+In contrast, CataractBot introduces a crucial buffer in the conversation between patients and doctors, allowing them to use the system at their convenience.
+However, this asynchronicity also posed challenges for both parties.
+Patients were unsure exactly when they would receive a verified response, while experts received no further feedback from patients after providing a verification.
+To address this, first, patients and attendants should be made aware of the experts’ availability and informed about the expected delay in verification, particularly when questions are sent outside of working hours, on surgery days, or when experts are on leave.
+To achieve this, experts should be able to specify their availability in natural language on CataractBot.
+This will also prevent disturbing experts during their personal time.
+Second, to provide experts with an acknowledgement after an answer is verified, patients and attendants should be offered a poll: ‘Yes, my question was answered’—which would notify the expert, and ‘No, please give me a call’—which would alert the patient coordinator.
+This should be approached with caution, as prior work (Weisz et al., 2019; Yeh et al., 2022) suggests that users tend to bypass the automated agent in favour of directly connecting with a human agent via voice call.
+7.1.3. Knowledge Base (KB)
+For CataractBot, we curated a custom KB.
+Even within our short deployment, we noticed the KB to be inadequate.
+For longitudinal deployments, we recommend implementing a simple method for experts to add new documents directly to the KB.
+Recognizing the impracticality of manual updates (after approval by the KB update expert) in rapidly changing environments, we also propose automatic updates based on the consensus among doctors in the edited answers.
+Lastly, as the KB updates, a method is needed to evaluate CataractBot’s answers over time.
+Tan et al.
+(2024) found that GPT-4’s evaluations of AI-generated responses align closely with human clinician assessments, demonstrating the LLM’s potential not only for generating but also for validating answers.
+Incorporating an LLM as an evaluator has the potential to further improve the process of automatically expanding the KB while ensuring that the quality of the responses remains intact.
+7.1.4. Beyond Cataract Surgery
+Cataract surgery, with its brief duration, straightforward procedure, high success rate, and extensive documentation, provided an ideal low-stakes environment for building and testing our experts-in-the-loop chatbot.
+A similar chatbot could be even more valuable for information seekers undergoing complex, atypical, or long-term treatments, such as glaucoma surgery, cancer treatment, or childbirth.
+With our open-sourced framework for developing experts-in-the-loop bots, this should be feasible.
+Furthermore, the applicability of such a system need not be confined to the medical domain.
+Other fields, such as legal, finance, education, or agriculture, could benefit from similar chatbots.
+In these scenarios, end-users (e.g., defendants, taxpayers, students, or farmers) could receive synchronous responses from an LLM operating on a custom KB and subsequently obtain verified responses asynchronously from experts (e.g., lawyers, accountants, teachers/career counselors, or agricultural experts, respectively).
+Expert-mediated bots can even be used to upskill less-trained experts, such as patient coordinators or paralegals.
+For instance, a maternal and child health bot could serve as a valuable learning resource for community health workers in rural or semi-urban areas, who typically have minimal (medical) training.
+Such a bot can help them provide accurate information to patients and their families while also boosting their confidence and competence.
 7.2. Design Implications
 7.2.1. Personalization
-Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information (Yang et al., 2023), they may also raise privacy concerns among certain users, potentially reducing usage.
-Additionally, personalized responses may increase the workload for experts, as they would need to refer to patient-specific documents for verification purposes.
-Many of these verification edits by experts may not be suitable for inclusion in the KB due to there patient-specific nature.
-Given our demographic of individuals with low literacy and low tech-literacy, individuals might require assistance in making an informed decision about consenting to the use of their health data (Okolo et al., 2021).
+Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information, they may also raise privacy concerns among certain users, potentially reducing usage.
+Additionally, personalized responses could increase the workload of experts, as they would need to refer to specific patient documents for verification.
+Many of these verification edits by experts may not be suitable for inclusion in the KB due to their patient-specific nature.
+Given our demographic of low literate and low tech-literate users, they might require assistance in making informed decisions about consenting to the use of their health data (Okolo et al., 2021).
 Also, adopting a binary approach–either sharing all medical records or none–may not align with user preferences.
 Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access, in line with the philosophy of the National Digital Health Mission (NDHM) (National Health Authority and Family Welfare, 2020) in India.
-To achieve this, a potential solution could be to ask users to manually upload images/PDFs of relevant documents (such as consultation sheets, discharge summaries, and test reports) they wish the bot to access.
-This approach empowers users to control the information available to the chatbot for a more personalized response.
-Currently, CataractBot operates as a reactive, user-driven system.
-However, with personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s treatment journey, providing advice on comorbidities and medication reminders.
-6.1.2. Reduce Workload for Experts
-Deploying CataractBot in the hospital increased the workload for patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates).
+A solution could be to ask users to manually upload relevant documents (such as consultation sheets, discharge summaries, and test reports) or textually enter details they wish the bot to access.
+This approach can empower users to control the information available to the bot for a more personalized experience.
+Currently, CataractBot operates as a user-driven system.
+With personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s journey, e.g., providing medication reminders.
+7.2.2. Reduce Workload of Experts
+Deploying CataractBot in the hospital increased the workload of patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates).
 To alleviate their burden, we propose ways to streamline CataractBot-related tasks and to use CataractBot to support them in their day-to-day activities.
-First, to simplify patient onboarding, we propose installing QR codes with a concise CataractBot introduction at key locations in the hospital.
-Users can scan the code, self-onboard by providing necessary information (language preference, operating doctor, and surgery date), and instantly engage with the bot.
+First, to simplify patient onboarding, QR codes can be installed at key locations in the hospital.
+Users can scan the code, self-onboard by providing necessary information (such as language preference, operating doctor, and surgery date), and instantly engage with the bot.
 Users should receive a brief tutorial video on interaction guidelines, emphasizing that the bot is not intended for emergencies.
 Second, CataractBot should be utilized to deliver general information about the surgery and provide status updates to patients and attendants, thereby assisting experts in managing their other responsibilities.
-Supporting routine procedures, as recommended by Li et al.
-(2023), would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time.
+This would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time.
 Third, to reduce the verification load for experts, we propose a system to pre-verify frequently asked questions automatically based on experts’ repeated confirmations of correctness.
-Fourth, we recommend introducing a feature that allows experts to ‘pause’ CataractBot by specifying their ‘unavailable until ¡date time¿’ status in natural language.
-This not only helps prevent disturbances for experts during surgery days, outside working hours, or when on leave, but also ensures that information seekers are better informed of the experts’ availability, also suggested by Yang et al.
-(2023).
 Finally, regarding audio messages, the lack of non-verbal cues hindered experts’ ability to deliver care (similar to (Li et al., 2023)).
-To address the limitations of translation and transcription technologies, the information-seeker’s original audio clip and transcribed query should be shared with the experts.
+To address the limitations of translation and transcription technologies, the information seeker’s original audio clip should also be shared with the experts.
 7.2.3. Usefulness
-Prior work found that patients often seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention (Attfield et al., 2006).
-However, access to CataractBot is currently provided post-scheduling of the surgery.
-Introducing the bot before surgery scheduling might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts.
-For consultation-related information needs, a similar bot connected with the user’s family doctor could be enabled to address generic medical information needs.
-Based on our findings, we recommend keeping CataractBot enabled for a month post-surgery.
+Attfield et al.
+(2006) reported that patients seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention.
+However, access to CataractBot is currently provided only post-scheduling of the surgery.
+Introducing the bot before this might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts.
 During bot interactions, a few participants expressed frustration with IDK responses (similar to previous findings (Xiao et al., 2023)), leading to discontinuation of usage.
@@
 For instance, patients could upload images of medications instead of providing written details.
-While this could improve the user experience, it may also increase the workload for experts.
-6.1.4. Scaling
-In this work, the use of CataractBot was limited to a small sample of 49 patients and attendants over 2 months, while Sankara Eye Hospital performs over 50 cataract surgeries daily.
-Scaling CataractBot to cater to all cataract patients would require several modifications.
-These include implementing a QR code-based self-onboarding system, establishing a more comprehensive custom knowledge base from the outset, streamlining the verification process involving multiple doctors and coordinators for each patient (considering the potential non-participation of some experts), and improving the KB update process by engaging multiple KB Update Experts.
-Even within our short deployment, we noticed the knowledge base becoming outdated and the KB Update expert adding incorrect information.
-For longitudinal deployments, we recommend engaging multiple hospital staff in managing the KB to reduce reliance on a single expert, and implementing a simple method for the staff to add new documents to the KB.
-Recognizing the impracticality of manual updates in rapidly changing environments (Xiao et al., 2023), we propose automatic updates for specific information sources, such as standard operating procedures.
-Cataract surgery, being short, straightforward, commonplace, highly successful, and extensively documented, resulted in minimal queries from participants to CataractBot.
-A chatbot similar to CataractBot could provide more value to information seekers undergoing complex and less common treatments, like glaucoma surgery or radiation therapy for head and neck cancer.
-With the existing framework for constructing experts-in-the-loop bot, creating such bots should be feasible333We will open-source our codebase to help the research community build and deploy their own end-to-end expert bots..
-Furthermore, the applicability of such a system need not be confined to the medical domain; other fields, such as legal or agriculture (Jain et al., 2018), could benefit from similar chatbots.
-In these scenarios, end-users (clients or farmers) could receive synchronous responses from an LLM operating on a custom knowledge base and subsequently obtain verified responses asynchronously from experts (lawyers or agricultural experts).
+While this could improve the user experience for information seekers, it may also increase the workload for experts.
 7.3. Limitations
@@
 The positive user responses and other conclusions should be validated by future, preferably longitudinal studies.
-Second, we need to conduct system-level analysis, such as scalability testing, power measurement, etc., in future work before scaling the proposed solution to the masses.
-Third, CataractBot uses responses from the expert-in-the-loop after approval by the KB Update expert to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB.
-This issue needs to be resolved and requires further investigation.
-Finally, although participants showed high trust in the initial encounter, IDK responses and inaccurate information in daily usage can lead to distrust of the system.
-We also note the potential risk of disseminating inaccurate medical information at scale with chatbot technologies.
-Therefore, the knowledge base should be carefully developed, reviewed, and updated by experts.
+Second, system-level analyses, such as scalability testing and power measurement, are needed before scaling the proposed solution to the masses.
+Third, CataractBot uses expert responses to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB.
+This issue requires further investigation.
+Finally, we note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies.
+Therefore, stronger methods are needed to develop, review, and update the KB, as well as to regularly evaluate the accuracy of bot’s responses.
 8. Conclusion
 Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare.
-To tackle this, we propose a novel solution—a LLM-powered expert-in-the-loop chatbot framework, that utilizes RAG (Retrieval Augmented Generation) over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously.
+To tackle this, we propose a novel solution—a LLM-powered experts-in-the-loop chatbot framework, that utilizes RAG over a custom KB to provide synchronous responses, and utilizes experts to provide verified responses asynchronously.
 Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention.
-Patients not only trusted the chatbot’s response but were also willing to wait, understanding that their busy doctors were verifying them.
-Simultaneously, doctors and coordinators, despite their hectic schedules, squeezed time during breaks and at home to help patients by verifying LLM-generated responses.
+Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them.
+Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses.
 The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
@@
 arXiv:2303.12528 [cs.CL]
+- Akerkar and Bichile (2004) Shashank M Akerkar and Lata S Bichile. 2004. Doctor patient relationship: changing dynamics in the information age. Journal of postgraduate medicine 50, 2 (2004), 120–122.
+- Arora et al.
+(2008) Neeraj K Arora, Bradford W Hesse, Barbara K Rimer, Kasisomayajula Viswanath, Marla L Clayman, and Robert T Croyle.
+2008.
+Frustrated and confused: the American public rates its cancer-related information-seeking experiences.
+Journal of general internal medicine 23 (2008), 223–228.
 - Attfield et al.
@@
 Frontiers in Digital Health 5 (2023), 60.
+- Beasley (2009) Robert E Beasley. 2009. Short message service (SMS) texting symbols: A functional analysis of 10,000 cellular phone text messages. The Reading Matrix 9, 2 (2009).
 - Beisecker and Beisecker (1990) Analee E.
@@
 http://www.jstor.org/stable/3765618
+- Bernstein et al.
+(2023) Isaac A Bernstein, Youchen Victor Zhang, Devendra Govil, Iyad Majid, Robert T Chang, Yang Sun, Ann Shue, Jonathan C Chou, Emily Schehlein, Karen L Christopher, et al.
+2023.
+Comparison of ophthalmologist and large language model chatbot responses to online patient eye care questions.
+JAMA Network Open 6, 8 (2023), e2330320–e2330320.
+- Bérubé et al.
+(2021) Caterina Bérubé, Zsolt Ferenc Kovacs, Elgar Fleisch, and Tobias Kowatsch.
+2021.
+Reliability of Commercial Voice Assistants’ Responses to Health-Related Questions in Noncommunicable Disease Management: Factorial Experiment Assessing Response Rate and Source of Information.
+Journal of Medical Internet Research 23, 12 (2021), e32161.
+- Bhat et al.
+(2021) Karthik S Bhat, Mohit Jain, and Neha Kumar.
+2021.
+Infrastructuring Telehealth in (In)Formal Patient-Doctor Contexts.
+Proc.
+ACM Hum.-Comput.
+Interact.
+5, CSCW2, Article 323 (oct 2021), 28 pages.
+https://doi.org/10.1145/3476064
 - Bickmore et al.
@@
 1265–1274.
+- Bodenheimer and Smith (2013) Thomas S Bodenheimer and Mark D Smith.
+2013.
+Primary care: proposed solutions to the physician shortage without training more physicians.
+Health Affairs 32, 11 (2013), 1881–1886.
 - Borkar (2023) Jaydeep Borkar. 2023. What can we learn from Data Leakage and Unlearning for Law? arXiv preprint arXiv:2307.10476 (2023).
@@
 https://doi.org/10.1145/3290605.3300865
+- Chandwani and Kulkarni (2016) Rajesh Chandwani and Vaibhavi Kulkarni.
+2016.
+Who’s the Doctor?
+Physicians’ Perception of Internet Informed Patients in India.
+In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16).
+Association for Computing Machinery, New York, NY, USA, 3091–3102.
+https://doi.org/10.1145/2858036.2858500
 - Christalle et al.
@@
 https://doi.org/10.1371/journal.pone.0209165
+- Chung (2017) Chia-Fang Chung.
+2017.
+Supporting patient-provider communication and engagement with personal informatics data.
+In Proceedings of the 2017 ACM International Joint Conference on Pervasive and Ubiquitous Computing and Proceedings of the 2017 ACM International Symposium on Wearable Computers (Maui, Hawaii) (UbiComp ’17).
+Association for Computing Machinery, New York, NY, USA, 335–338.
+https://doi.org/10.1145/3123024.3123197
 - Civan-Hartzler et al.
@@
 2002), 1633–1637.
+- De Choudhury et al.
+(2014) Munmun De Choudhury, Meredith Ringel Morris, and Ryen W.
+White.
+2014.
+Seeking and sharing health information online: comparing search engines and social media.
+In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14).
+Association for Computing Machinery, New York, NY, USA, 1365–1376.
+https://doi.org/10.1145/2556288.2557214
+- Denecke et al.
+(2024) Kerstin Denecke, Richard May, and Octavio Rivera Romero.
+2024.
+Potential of Large Language Models in Health Care: Delphi Study.
+J Med Internet Res 26 (13 May 2024), e52399.
+https://doi.org/10.2196/52399
+- Desai and Chin (2023) Smit Desai and Jessie Chin.
+2023.
+OK Google, Let’s Learn: Using Voice User Interfaces for Informal Self-Regulated Learning of Health Topics among Younger and Older Adults.
+In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23).
+Association for Computing Machinery, New York, NY, USA, Article 847, 21 pages.
+https://doi.org/10.1145/3544548.3581507
+- Ding et al.
+(2020) Xianghua Ding, Xinning Gui, Xiaojuan Ma, Zhaofei Ding, and Yunan Chen.
+2020.
+Getting the Healthcare We Want: The Use of Online ”Ask the Doctor” Platforms in Practice.
+In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20).
+Association for Computing Machinery, New York, NY, USA, 1–13.
+https://doi.org/10.1145/3313831.3376699
 - Duffy et al.
@@
 https://doi.org/10.1145/1518701.1518971
+- Fochsen et al.
+(2006) Grethe Fochsen, Kirti Deshpande, and Anna Thorson.
+2006.
+Power Imbalance and Consumerism in the Doctor-Patient Relationship: Health Care Providers’ Experiences of Patient Encounters in a Rural District in India.
+Qualitative Health Research 16, 9 (2006), 1236–1251.
+https://doi.org/10.1177/1049732306293776 arXiv:https://doi.org/10.1177/1049732306293776 PMID: 17038755.
+- Fogel and Kvedar (2018) Alexander L Fogel and Joseph C Kvedar. 2018. Artificial intelligence powers digital medicine. NPJ digital medicine 1, 1 (2018), 5.
+- Gordon et al.
+(2020) Howard S.
+Gordon, Pooja Solanki, Barbara G.
+Bokhour, and Ravi K.
+Gopal.
+2020.
+“I’m Not Feeling Like I’m Part of the Conversation” Patients’ Perspectives on Communicating in Clinical Video Telehealth Visits.
+Journal of General Internal Medicine 35, 6 (01 Jun 2020), 1751–1758.
+https://doi.org/10.1007/s11606-020-05673-w
+- Grönvall and Verdezoto (2013) Erik Grönvall and Nervo Verdezoto.
+2013.
+Beyond self-monitoring: understanding non-functional aspects of home-based healthcare technology.
+In Proceedings of the 2013 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Zurich, Switzerland) (UbiComp ’13).
+Association for Computing Machinery, New York, NY, USA, 587–596.
+https://doi.org/10.1145/2493432.2493495
+- Henry et al.
+(2022) Katharine E Henry, Rachel Kornfield, Anirudh Sridharan, Robert C Linton, Catherine Groh, Tony Wang, Albert Wu, Bilge Mutlu, and Suchi Saria.
+2022.
+Human–machine teaming is key to AI adoption: clinicians’ experiences with a deployed machine learning system.
+NPJ digital medicine 5, 1 (2022), 97.
+- Holmes et al.
+(2023) Samuel Holmes, Raymond Bond, Anne Moorhead, Jane Zheng, Vivien Coates, and Michael McTear.
+2023.
+Towards Validating a Chatbot Usability Scale.
+In Design, User Experience, and Usability, Aaron Marcus, Elizabeth Rosenzweig, and Marcelo M.
+Soares (Eds.).
+Springer Nature Switzerland, Cham, 321–339.
+- Iliashenko et al.
+(2019) Oksana Iliashenko, Zilia Bikkulova, and Alissa Dubgorn.
+2019.
+Opportunities and challenges of artificial intelligence in healthcare.
+In E3S Web of Conferences, Vol.
+110.
+EDP Sciences, 02028.
+- Ittarat et al.
+(2023) Mantapond Ittarat, Wisit Cheungpasitporn, and Sunee Chansangpetch.
+2023.
+Personalized Care in Eye Health: Exploring Opportunities, Challenges, and the Road Ahead for Chatbots.
+Journal of Personalized Medicine 13, 12 (2023), 1679.
 - Jain et al.
@@
 323–332.
+- Kapania et al.
+(2022) Shivani Kapania, Oliver Siy, Gabe Clapper, Azhagu Meena SP, and Nithya Sambasivan.
+2022.
+”Because AI is 100% right and safe”: User Attitudes and Sources of AI Authority in India.
+In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿New Orleans¡/city¿, ¡state¿LA¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’22).
+Association for Computing Machinery, New York, NY, USA, Article 158, 18 pages.
+https://doi.org/10.1145/3491102.3517533
 - Kinnersley et al.
@@
 https://doi.org/10.1145/1753326.1753355
+- Kocaballi et al.
+(2020) A Baki Kocaballi, Kiran Ijaz, Liliana Laranjo, Juan C Quiroz, Dana Rezazadegan, Huong Ly Tong, Simon Willcock, Shlomo Berkovsky, and Enrico Coiera.
+2020.
+Envisioning an artificial intelligence documentation assistant for future primary care consultations: A co-design study with general practitioners.
+Journal of the American Medical Informatics Association 27, 11 (08 2020), 1695–1704.
+https://doi.org/10.1093/jamia/ocaa131 arXiv:https://academic.oup.com/jamia/article-pdf/27/11/1695/34363907/ocaa131.pdf
 - Kumar et al.
@@
 https://doi.org/10.1017/S0317167100001906
-- Larasati et al.
-(2023) Retno Larasati, Anna De Liddo, and Enrico Motta.
-2023.
-Meaningful Explanation Effect on User’s Trust in an AI Medical System: Designing Explanations for Non-Expert Users.
-ACM Trans.
-Interact.
-Intell.
-Syst.
-13, 4, Article 30 (dec 2023), 39 pages.
-https://doi.org/10.1145/3631614
 - Leong et al.
@@
 Frontiers in Public Health 11 (2023), 1147789.
+- Makarem and Antoun (2016) Nisrine N.
+Makarem and Jumana Antoun.
+2016.
+Email communication in a developing country: different family physician and patient perspectives.
+Libyan Journal of Medicine 11, 1 (2016), 32679.
+https://doi.org/10.3402/ljm.v11.32679 arXiv:https://doi.org/10.3402/ljm.v11.32679 PMID: 28349842.
+- Malki et al.
+(2024) Lisa Mekioussa Malki, Dilisha Patel, and Aneesha Singh.
+2024.
+”The Headline Was So Wild That I Had To Check”: An Exploration of Women’s Encounters With Health Misinformation on Social Media.
+Proc.
+ACM Hum.-Comput.
+Interact.
+8, CSCW1, Article 128 (apr 2024), 26 pages.
+https://doi.org/10.1145/3637405
+- Mamykina et al.
+(2008) Lena Mamykina, Elizabeth Mynatt, Patricia Davidson, and Daniel Greenblatt.
+2008.
+MAHI: investigation of social scaffolding for reflective thinking in diabetes management.
+In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Florence, Italy) (CHI ’08).
+Association for Computing Machinery, New York, NY, USA, 477–486.
+https://doi.org/10.1145/1357054.1357131
 - Moden (2022) Rebecca Moden. 2022. Blank Page Syndrome And How To Beat It. Retrieved Sept 1, 2023 from https://www.ool.co.uk/blog/blank-page-syndrome-and-how-to-beat-it/
@@
 https://www.niti.gov.in/sites/default/files/2023-02/ndhm_strategy_overview.pdf
+- Office of the Registrar General & Census Commissioner (2011) Government of India Office of the Registrar General & Census Commissioner, India; Ministry of Home Affairs.
+2011.
+Census of India 2011: C-16: Population by mother tongue, Karnataka - 2011.
+https://censusindia.gov.in/nada/index.php/catalog/10208/study-description Accessed: 2024-07-24.
 - Okolo et al.
@@
 https://doi.org/10.1145/3411764.3445420
+- OpenAI (2024) OpenAI. 2024. Usage Policies. https://openai.com/policies/usage-policies/ Accessed: 2024-06-18.
 - OpenAI et al.
@@
 arXiv:arXiv:2303.08774
+- Osman Andersen et al.
+(2021) Tariq Osman Andersen, Francisco Nunes, Lauren Wilcox, Elizabeth Kaziunas, Stina Matthiesen, and Farah Magrabi.
+2021.
+Realizing AI in Healthcare: Challenges Appearing in the Wild.
+In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (, Yokohama, Japan,) (CHI EA ’21).
+Association for Computing Machinery, New York, NY, USA, Article 108, 5 pages.
+https://doi.org/10.1145/3411763.3441347
+- Perrier et al.
+(2015) Trevor Perrier, Nicola Dell, Brian DeRenzi, Richard Anderson, John Kinuthia, Jennifer Unger, and Grace John-Stewart.
+2015.
+Engaging Pregnant Women in Kenya with a Hybrid Computer-Human SMS Communication System.
+In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15).
+Association for Computing Machinery, New York, NY, USA, 1429–1438.
+https://doi.org/10.1145/2702123.2702124
+- Pfeifer Vardoulakis et al.
+(2012) Laura Pfeifer Vardoulakis, Amy Karlson, Dan Morris, Greg Smith, Justin Gatewood, and Desney Tan.
+2012.
+Using mobile phones to present medical information to hospital patients.
+In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Austin, Texas, USA) (CHI ’12).
+Association for Computing Machinery, New York, NY, USA, 1411–1420.
+https://doi.org/10.1145/2207676.2208601
+- Prasad et al.
+(2020) Manya Prasad, Sumit Malhotra, Mani Kalaivani, Praveen Vashist, and Sanjeev K Gupta.
+2020.
+Gender differences in blindness, cataract blindness and cataract surgical coverage in India: a systematic review and meta-analysis.
+British Journal of Ophthalmology 104, 2 (2020), 220–224.
 - Radionova et al.
@@
 Journal of Medical Internet Research 25 (2023), e39219.
+- Rajashekar et al.
+(2024) Niroop Channa Rajashekar, Yeo Eun Shin, Yuan Pu, Sunny Chung, Kisung You, Mauro Giuffre, Colleen E Chan, Theo Saarinen, Allen Hsiao, Jasjeet Sekhon, Ambrose H Wong, Leigh V Evans, Rene F.
+Kizilcec, Loren Laine, Terika Mccall, and Dennis Shung.
+2024.
+Human-Algorithmic Interaction Using a Large Language Model-Augmented Artificial Intelligence Clinical Decision Support System.
+In Proceedings of the CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Honolulu¡/city¿, ¡state¿HI¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’24).
+Association for Computing Machinery, New York, NY, USA, Article 442, 20 pages.
+https://doi.org/10.1145/3613904.3642024
 - Roller et al.
@@
 arXiv preprint arXiv:2004.13637 (2020).
+- Seitz et al.
+(2022) Lennart Seitz, Sigrid Bekmeier-Feuerhahn, and Krutika Gohil.
+2022.
+Can we trust a chatbot like a physician?
+A qualitative study on understanding the emergence of trust toward diagnostic chatbots.
+International Journal of Human-Computer Studies 165 (2022), 102848.
+https://doi.org/10.1016/j.ijhcs.2022.102848
 - Singh et al. (2017) Kamaljeet Singh, Arshi Misbah, Pranav Saluja, and Arun Kumar Singh. 2017. Review of manual small-incision cataract surgery. Indian journal of ophthalmology 65, 12 (2017), 1281.
@@
 https://www.statista.com/statistics/1235799/india-most-used-social-media-and-messaging-apps/ Accessed: 2024-02-01.
+- Tan et al.
+(2024) Ting Fang Tan, Kabilan Elangovan, Liyuan Jin, Yao Jie, Li Yong, Joshua Lim, Stanley Poh, Wei Yan Ng, Daniel Lim, Yuhe Ke, et al.
+2024.
+Fine-tuning Large Language Model (LLM) Artificial Intelligence Chatbots in Ophthalmology and LLM-based evaluation using GPT-4.
+arXiv preprint arXiv:2402.10083 (2024).
 - Tang and Lansky (2005) Paul C Tang and David Lansky. 2005. The missing link: bridging the patient-provider health information gap. Health Aff. (Millwood) 24, 5 (Sept. 2005), 1290–1295.
 - Tang and Newcomb (1998) P C Tang and C Newcomb. 1998. Informing patients: a guide for providing patient health information. J. Am. Med. Inform. Assoc. 5, 6 (Nov. 1998), 563–570.
-- ter Hofstede et al.
-(1996) A.
-H.
-M.
-ter Hofstede, H.
-A.
-Proper, and Th.
-P.
-van der Weide.
-1996.
-Query Formulation as an Information Retrieval Problem.
-Comput.
-J.
-39, 4 (01 1996), 255–274.
-https://doi.org/10.1093/comjnl/39.4.255 arXiv:https://academic.oup.com/comjnl/article-pdf/39/4/255/1916704/390255.pdf
+- Thieme et al.
+(2023a) Anja Thieme, Maryann Hanratty, Maria Lyons, Jorge Palacios, Rita Faia Marques, Cecily Morrison, and Gavin Doherty.
+2023a.
+Designing Human-centered AI for Mental Health: Developing Clinically Relevant Applications for Online CBT Treatment.
+ACM Trans.
+Comput.-Hum.
+Interact.
+30, 2, Article 27 (mar 2023), 50 pages.
+https://doi.org/10.1145/3564752
+- Thieme et al.
+(2023b) Anja Thieme, Aditya Nori, Marzyeh Ghassemi, Rishi Bommasani, Tariq Osman Andersen, and Ewa Luger.
+2023b.
+Foundation Models in Healthcare: Opportunities, Risks & Strategies Forward.
+In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (, Hamburg, Germany,) (CHI EA ’23).
+Association for Computing Machinery, New York, NY, USA, Article 512, 4 pages.
+https://doi.org/10.1145/3544549.3583177
 - Tongue et al.
@@
 https://doi.org/10.1145/3610106
+- Tuli et al.
+(2018) Anupriya Tuli, Shaan Chopra, Neha Kumar, and Pushpendra Singh.
+2018.
+Learning from and with Menstrupedia: Towards Menstrual Health Education in India.
+Proc.
+ACM Hum.-Comput.
+Interact.
+2, CSCW, Article 174 (nov 2018), 20 pages.
+https://doi.org/10.1145/3274443
 - Wang et al.
@@
 https://doi.org/10.1145/3449083
+- Wei et al.
+(2024) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H.
+Chi, Quoc V.
+Le, and Denny Zhou.
+2024.
+Chain-of-thought prompting elicits reasoning in large language models.
+In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22).
+Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
 - Weisz et al.
@@
 - Williams and Moser (2019) Michael Williams and Tami Moser. 2019. The art of coding and thematic exploration in qualitative research. International Management Review 15, 1 (2019), 45–55.
+- Wutz et al.
+(2023) Maximilian Wutz, Marius Hermes, Vera Winter, and Juliane Köberlein-Neu.
+2023.
+Factors influencing the acceptability, acceptance, and adoption of conversational agents in health care: integrative review.
+Journal of Medical Internet Research 25 (2023), e46548.
 - Xiao et al.
@@
 - Yang et al.
-(2023) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Shao Zhang, Ethan Rogers, Stephen Intille, Nawar Shara, Dakuo Wang, et al.
-2023.
-Talk2Care: Facilitating Asynchronous Patient-Provider Communication with Large-Language-Model.
-arXiv preprint arXiv:2309.09357 (2023).
-arXiv:2309.09357 [cs.CL]
+(2024) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Ethan Rogers, Shao Zhang, Stephen Intille, Nawar Shara, Guodong Gordon Gao, and Dakuo Wang.
+2024.
+Talk2Care: An LLM-based Voice Assistant for Communication between Healthcare Providers and Older Adults.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 2, Article 73 (may 2024), 35 pages.
+https://doi.org/10.1145/3659625
+- Ye et al.
+(2020) Qunru Ye, Yanxian Chen, William Yan, Wei Wang, Jingxian Zhong, Cong Tang, Andreas Müller, Bo Qiu, et al.
+2020.
+Female gender remains a significant barrier to access cataract surgery in South Asia: A systematic review and meta-analysis.
+Journal of ophthalmology 2020 (2020).
+- Yeh et al.
+(2022) Su-Fang Yeh, Meng-Hsin Wu, Tze-Yu Chen, Yen-Chun Lin, XiJing Chang, You-Hsuan Chiang, and Yung-Ju Chang.
+2022.
+How to Guide Task-oriented Chatbot Users, and When: A Mixed-methods Study of Combinations of Chatbot Guidance Types and Timings.
+In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22).
+Association for Computing Machinery, New York, NY, USA, Article 488, 16 pages.
+https://doi.org/10.1145/3491102.3501941
+- Yilmaz and Dogan (2024) Ibrahim Edhem Yilmaz and Levent Dogan. 2024. Talking technology: exploring chatbots as a tool for cataract patient education. Clinical and Experimental Optometry (2024), 1–9.
 - YM et al.
@@
 arXiv:2309.09495 [cs.HC]
+- Zhang (2010) Yan Zhang.
+2010.
+Contextualizing consumer health information searching: an analysis of questions in a social Q&A community.
+In Proceedings of the 1st ACM international health informatics symposium.
+210–219.
 Appendix A Appendix
@@
 A.1.1. Cataract Bot LLM Prompts
-The CataractBot system leverages LLM (GPT-4 in our case) for these four tasks:
+The CataractBot system leverages LLM (GPT-4 in our case) for these four tasks (Table 3):
 Response Generation: For every medical and logistical question asked by the patient/attendant, the system performs a vector search on the Knowledge Base (KB) to extract the three most relevant data chunks related to the query.
@@
 Shorten Response: If the generated response exceeds WhatsApp’s message limit of 700 characters, the system prompts LLM to summarize it within the specified character limit.
-| System Prompt | Query Prompt | ||||||||||||||||||||||||||||||||||||
+| System Prompt | Query Prompt | |||||||||||||||||||||||||||||||||||||||||||
+A.2. Chatbot Usability Form
+Responses were given on a 5-point Likert scale, ranging from Strongly Disagree to Strongly Agree.
+CataractBot understands me well.
+CataractBot’s responses are easy to understand.
+CataractBot’s responses were useful, appropriate, and informative.
+CataractBot responds quickly.
+CataractBot seems to have a good grasp of medical knowledge.
+CataractBot is kind and helpful.
+CataractBot is easy to use.
+I would be willing to use CataractBot (or a similar bot before/after a major surgical treatment) in future.